### v3

# Alignment-Sensitive Minimax Rates for Spectral Algorithms with Learned Kernels

Abstract
We study spectral algorithms in the setting where kernels are learned from data. We introduce the effective span dimension (ESD), an alignment-sensitive complexity measure that depends jointly on the signal, spectrum, and noise level $\sigma^{2}$. The ESD is well-defined for arbitrary kernels and signals without requiring eigen-decay conditions or source conditions. We prove that for sequence models whose ESD is at most $K$, the minimax excess risk scales as $\sigma^{2}K$. Furthermore, we analyze over-parameterized gradient flow and prove that it can reduce the ESD. This finding establishes a connection between adaptive feature learning and provable improvements in generalization of spectral algorithms. We demonstrate the generality of the ESD framework by extending it to linear models and RKHS regression, and we support the theory with numerical experiments. This framework provides a novel perspective on generalization beyond traditional fixed-kernel theories.
1 Introduction
Many modern learning procedures are adaptive: during training they update a representation, which effectively changes the induced kernel. Such adaptivity can improve generalization when it increases alignment between the learned kernel and the target function [4, 27, 31, 10, 47, 50]. In contrast, classical theory for fixed kernel regression does not capture this effect. For instance, the Neural Tangent Kernel (NTK) theory approximates training dynamics as kernel regression [23, 2], which enables the study of generalization using the classical regression theory in Reproducing Kernel Hilbert Spaces (RKHS) [8, 48]. However, the NTK approximation is intrinsically non adaptive and therefore cannot explain why finite width networks, whose features evolve during training, often outperform traditional methods [19, 16, 25, 37, 43, 36].
Recent simplified models isolate one tractable form of adaptivity: learning the kernel eigenvalues while keeping the eigenfunctions fixed. In these settings, reshaping the spectrum can increase signal kernel alignment and lead to better performance [29, 30]. More broadly, many analyses point to the same mechanism: generalization improves when the target places more energy on the leading kernel eigenfunctions [3, 44, 26, 33]. However, classical RKHS theory is typically developed under fixed spectral assumptions, such as polynomial eigenvalue decay, together with signal regularity conditions, such as source conditions [15]. If the kernel is learned, these assumptions are difficult to justify. To explain the advantages of adaptive kernel methods, we need a refined theoretical framework that remains valid beyond fixed-kernel assumptions.
In this paper, we propose the Effective Span Dimension (ESD) to bridge fixed-kernel theory and adaptive learning. The ESD is a population complexity measure to quantify signal-kernel alignment. Unlike classical measures that ignore the signal, ESD depends on the signal, spectrum, and noise level. Importantly, it is well defined for any kernel, including adaptive kernels whose eigenvalues and eigenfunctions evolve during training. Our framework provides new theoretical insights that are absent in classical analyses. In particular, we achieve the following:
  1. 1.
    We establish sharp minimax optimal convergence rates for sequence models using ESD.
  2. 2.
    We extend our definitions and theory to linear regression and kernel regression, which demonstrates the broad applicability of our framework.
  3. 3.
    We analyze a tractable form of eigenvalue learning to obtain provable ESD reduction. We also empirically illustrate how the ESD framework applies to adaptive kernels with evolving eigenfunctions.
The ESD quantitatively explains the relationship between signal-kernel alignment and generalization. In our framework, the benefit of adaptive learning is exactly a reduction in ESD. Before adaptive training, the kernel can be badly aligned, so the ESD is large and the target lies in a minimax hard family. After training, the kernel is reshaped so that the ESD becomes smaller, which moves the same target into an easier family where efficient spectral methods can generalize well.
Notations.
Write $a\lesssim b$ if there exists a constant $C>0$ such that $a\leq Cb$, and write $a\asymp b$ if $a\lesssim b$ and $b\lesssim a$, where the dependence of the constants on other parameters is determined by the context. For $d\in\mathbb{N}_{+}$, let $d]=\{1,2,\ldots,d\}$; for $d=\infty$, let $d]=\mathbb{N}_{+}$. $\mathbf{1}_{\{\cdot\}}$ denotes an indicator function.
2 Background on Kernel Methods
We first review kernel regression to provide context. Let $(\bm{x}_{i},y_{i})_{i=1}^{n}$ be i.i.d. samples from $y=f^{*}(\bm{x})+\epsilon$, where $\bm{x}\sim\mu$ on a compact space $\mathcal{X}$, $\epsilon$ is an independent noise with $\mathbb{E}[\epsilon]=0,\mathrm{Var}(\epsilon)=\sigma_{0}^{2}$. For an estimator $\hat{f}$ of the target function $f^{*}$, the excess risk is $\mathcal{R}(\hat{f};f^{*})=\mathbb{E}_{\bm{x}\sim\mu}\bigl[(\hat{f}(\bm{x})-f^{*}(\bm{x}))^{2}\bigr]$.
A symmetric, positive-definite, and continuous kernel $\mathbf{k}(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ induces an RKHS $\mathcal{H}\subset L^{2}(\mathcal{X},\mu)$ with inner product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ and norm $\|\cdot\|_{\mathcal{H}}$ [41, 35]. Assuming $\mathbf{k}$ is bounded, Mercer’s theorem yields
$\vskip-3.99994pt\mathbf{k}\left(\bm{x},\bm{x}^{\prime}\right)=\sum_{j=1}^{\infty}\lambda_{j}\psi_{j}(\bm{x})\psi_{j}\left(\bm{x}^{\prime}\right),\quad\bm{x},\bm{x}^{\prime}\in\mathcal{X},\vskip 0.0pt$ (1)
where $\{\lambda_{j}\}_{j\geq 1}$ are eigenvalues and $\{\psi_{j}\}_{j\geq 1}\subset\mathcal{H}$ are eigenfunctions forming an orthonormal basis of $L^{2}(\mathcal{X},\mu)$. For background, see Steinwart and Christmann [38], Steinwart and Scovel [39].
Kernel regression estimates $f^{*}$ using $f=\sum_{j}\beta_{j}\psi_{j}$ and regularizes via a filter of the kernel spectrum $\{\lambda_{j}\}$ [34, 12, 18]. If $f^{*}$ satisfies the Hölder source condition $\sum_{j}\langle f^{*},\psi_{j}\rangle^{2}/\lambda_{j}^{s}\leq R_{s}$ for some positive constants $s$ and $R_{s}$ [15, 32] and the spectrum decays polynomially $\lambda_{j}\asymp j^{-\gamma}$, then the minimax rate is $n^{-s\gamma/(s\gamma+1)}$ [48, 28, 42]. The choice of kernels can significantly affect the performance [29, 50], so it is beneficial when the kernel eigenvalues align well with the expansion of the target function.
Since the kernel is usually chosen without knowing $f^{*}$, fixed-kernel methods may encounter misalignment. To address this limitation, adaptive methods have recently emerged. For instance, Li and Lin [30] propose adapting kernel eigenvalues while fixing eigenfunctions. Specifically, they consider the kernel $\mathbf{k}_{\bm{a}}\left(x,x^{\prime}\right)=\sum_{j\geq 1}a_{j}^{2}\psi_{j}(x)\psi_{j}\left(x^{\prime}\right)$ indexed by $\bm{a}=\left(a_{j}\right)_{j\geq 1}$ and the candidate $f=\sum_{j\geq 1}\beta_{j}a_{j}\psi_{j}$, where $a_{j}$’s and $\beta_{j}$’s are learned jointly via gradient flow. Such adaptation often improves performances, yet classical analyses built on fixed spectral assumptions do not explain these gains, because
  1. 1.
    adapted eigenvalues typically deviate from standard eigenvalue decay assumptions, and
  2. 2.
    it is unclear whether the classical source condition holds with respect to the adapted kernel, and if so, what the value of $s$ is.
Given the above limitation of classical analyses, we therefore seek a refined theoretical framework that explicitly captures signal-kernel alignment and explains the gains achieved by kernel adaptation.
Bridge to the sequence model.
We next connect the RKHS regression with the sequence model to motivate our analysis in the next section. For any $j\in\mathbb{N}_{+}$, define
$\theta^{*}_{j}=\langle f^{*},\psi_{j}\rangle,\quad z_{j}=n^{-1}\sum_{i}y_{i}\psi_{j}(\bm{x}_{i}),\text{ and }\xi_{j}=n^{-1}\sum_{i}\epsilon_{i}\psi_{j}(\bm{x}_{i}).$ (2)
The law of large numbers implies that for large $n$, $n^{-1}\sum_{i}\psi_{j}(\bm{x}_{i})\psi_{k}(\bm{x}_{i})\approx\mathbb{E}[\psi_{j}(\bm{x})\psi_{k}(\bm{x})]=\mathbf{1}_{\{j=k\}}$, which implies that
$z_{j}\approx\theta^{*}_{j}+\xi_{j},\text{ and }\mathbb{E}\,[\xi_{j}]=0,\mathrm{Cov}(\xi_{j},\xi_{k})\approx n^{-1}\sigma_{0}^{2}\mathbf{1}_{\{j=k\}},\quad\forall j,k\in\mathbb{N}_{+}.$ (3)
This reduction connects RKHS regression to a sequence model where the observations are $z_{j}=\theta^{*}_{j}+\xi_{j}$ and the noise terms $\{\xi_{j}\}$ are uncorrelated with variance $\sigma^{2}:=n^{-1}\sigma_{0}^{2}$. The error in the approximation due to finite $n$ will inflate the estimation variance compared to the sequence model. This approximation error can be controlled if $f^{*}$ is bounded; see Section˜9 for a rigorous treatment.
3 Effective Span Dimension and Span Profile
To bridge existing theory and adaptive kernel methods as discussed in Section 2, we propose a novel framework to characterize the alignment between spectrum and signal. To focus on the main idea, we use the reduction in Equation˜3 and first present our framework using sequence models.
Sequence models.
A sequence model assumes observations are sampled as follows:
$z_{j}=\theta_{j}^{*}+\xi_{j},\quad 1\leq j\leq d,$ (4)
where $d\in\{\infty\}\cup\mathbb{N}_{+}$, $\bm{\theta}^{*}=(\theta_{j}^{*})_{j=1}^{d}$ is a sequence of unknown parameters, $\xi_{j}$’s are uncorrelated random variables with mean zero and variance $\sigma^{2}$ (the noise level). For an estimator $\widehat{\bm{\theta}}=(\widehat{\theta}_{j})_{j=1}^{d}$, we consider the loss $\mathcal{L}(\widehat{\bm{\theta}};\bm{\theta}^{*})=\sum_{j=1}^{d}(\widehat{\theta}_{j}-\theta_{j}^{*})^{2}$ and risk $\mathcal{R}(\widehat{\bm{\theta}};\bm{\theta}^{*})=\mathbb{E}\mathcal{L}(\widehat{\bm{\theta}};\bm{\theta}^{*})$. The sequence model captures core estimation phenomena while permitting explicit analysis [11, 24]. In Section˜8, we use whitening to deal with correlated noise and analyze fixed-design linear regression. In Section˜9, we leverage the approximation in Equation˜3 to analyze RKHS regression and random-design linear regression.
Spectral estimators. Given eigenvalues $\bm{\lambda}=(\lambda_{j})_{j=1}^{d}$, spectral estimators take the form $\widehat{\theta}_{j}=\bigl(1-\psi_{\nu}(\lambda_{j})\bigr)\,z_{j}$, where $\psi_{\nu}(\lambda)$ is a filter such that larger $\nu$ induces more shrinkage. Some examples are:
$\text{Ridge (R):}\quad\psi_{\nu}^{\operatorname{R}}(\lambda)=\frac{1}{1+\lambda/\nu},\widehat{\theta}_{j}^{\operatorname{R},\nu}=\frac{\lambda_{j}}{\lambda_{j}+\nu}\,z_{j}.$ (5)
$\text{Gradient Flow (GF):}\quad\psi_{\nu}^{\operatorname{GF}}(\lambda)=e^{-\lambda/\nu},$
$\qquad\qquad\qquad\quad\qquad\widehat{\theta}_{j}^{\operatorname{GF},\nu}=\left(1-e^{-\lambda_{j}/\nu}\right)z_{j}.$ (6)
$\text{Principal Component (PC):}\quad\psi_{\nu}^{\operatorname{PC}}(\lambda)=\mathbf{1}_{\{\lambda<\nu\}},$
$\qquad\qquad\qquad\qquad\qquad\qquad\widehat{\theta}_{j}^{\operatorname{PC},\nu}=\mathbf{1}_{\{\lambda_{j}\geq\nu\}}\,z_{j}.$ (7)
For spectral estimators, the risk decomposes into squared bias $\sum_{j}\bigl(\psi_{\nu}(\lambda_{j})\bigr)^{2}\,\theta_{j}^{2}$ and variance $\sum_{j}\bigl(1-\psi_{\nu}(\lambda_{j})\bigr)^{2}\,\sigma^{2}$, where $\nu$ controls the bias-variance trade-off. Classical analyses often assume $\bm{\theta}^{*}$ lies in an ellipsoid $\Theta_{\mathbf{a}}=\left\{\theta:\sum_{j}^{\infty}a_{j}^{2}\theta_{j}^{2}\leq C^{2}\right\}$ and derive convergence rates for sequences with $a_{i}\asymp i^{\alpha}$ [24]. Our theoretical framework aims to bypass these assumptions.
3.1 Effective Span Dimension
Our goal is to quantify how estimation difficulty depends jointly on the signal $\bm{\theta}^{*}$, the spectrum $\bm{\lambda}$, and the noise variance $\sigma^{2}$. We first introduce the effective span dimension (ESD) as an intrinsic, population-level quantity determined by $(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})$ through the ordering of the spectrum.
Definition 3.1.
Suppose $\{\lambda_{j}\}_{j\in[d]}$ are distinct, with $\pi_{i}$ indexing the $i$-th largest so that $\lambda_{\pi_{1}}>\lambda_{\pi_{2}}>\dots$. We define the Effective Span Dimension (ESD) $d^{\dagger}$ of $\bm{\theta}^{*}$ w.r.t. the spectrum $\bm{\lambda}$ and variance $\sigma^{2}$ as
$d^{\dagger}(\sigma^{2};\bm{\theta}^{*},\bm{\lambda})=\min\{k\in[d]:\frac{1}{k}\sum_{i=k+1}^{d}\left(\theta_{\pi_{i}}^{*}\right)^{2}\leq\,\sigma^{2}\}.$
When $\lambda_{j}$ is decreasing, the summation in the above definition is simply the sum of the squared tail of $\bm{\theta}^{*}$. Using the indices $\pi_{i}$, the definition applies to general ordering.
Intuitively, the ESD $d^{\dagger}$ is the smallest number $k$ of leading eigencoordinates (ordered by $\bm{\lambda}$) needed so that the remaining signal energy falls below $k\sigma^{2}$. It is a complexity index of the pair (signal, spectrum) at noise level $\sigma^{2}$, and it exists regardless of which algorithm is used for estimation.
We next show that $d^{\dagger}$ describes the best achievable risk for the PC estimator, which serves as an achievability argument for the rates that will be shown minimax optimal in Theorem˜3.3.
Theorem 3.2.
Let $\widehat{\bm{\theta}}^{\operatorname{PC},\nu}$ be the PC estimator in Equation˜7 for the sequence model in Equation˜4. Denote by $\mathcal{R}_{*}^{\operatorname{PC}}$ the minimal possible risk over all choices of $\nu$. Let $d^{\dagger}=d^{\dagger}(\sigma^{2};\bm{\theta}^{*},\bm{\lambda})$ be the ESD of $\bm{\theta}^{*}$ w.r.t. the spectrum $\bm{\lambda}$ and the variance $\sigma^{2}$. It holds that
$(d^{\dagger}-1)\,\sigma^{2}\leq\mathcal{R}_{*}^{\operatorname{PC}}\;\leq\;2\,d^{\dagger}\,\sigma^{2}.$
In classical theories on sequence models, the well-tuned PC estimator is known to be minimax rate optimal under regularity assumptions that are analogous to the polynomial eigen-decay condition and the source condition in kernel regression (see Propositions 3.11 and 4.23 of Johnstone [24]). In view of Theorem˜3.2, those minimax rates should be at the order of $O(d^{\dagger}\sigma^{2})$. Indeed, we can use the ESD to characterize the intrinsic difficulty of estimation, which is the goal of the following theorem.
Theorem 3.3.
For any $K\in[d]$, spectrum $\bm{\lambda}=\{\lambda_{j}\}_{j\in[d]}$, and variance $\sigma^{2}$, define
$\mathcal{F}_{K,\bm{\lambda}}^{(n)}=\Bigl\{\bm{\theta}\in\mathbb{R}^{d}:d^{\dagger}(\sigma^{2};\bm{\theta},\bm{\lambda})\leq K\Bigr\}.$ (8)
Suppose the sample $\bm{Z}$ is drawn from the sequence model in Equation˜4 and consider any estimator $\widehat{\bm{\theta}}$ based on $\bm{Z}$. We have
$\inf_{\widehat{\bm{\theta}}}\sup_{\bm{\theta}^{*}\in\mathcal{F}_{K,\bm{\lambda}}^{(n)}}\mathcal{R}(\widehat{\bm{\theta}},\bm{\theta}^{*})\asymp K\sigma^{2}.$
The class $\mathcal{F}_{K,\bm{\lambda}}^{(n)}$ consists of distributions whose ESDs are at most $K$. We interpret $K$ as the quota for ESD: the larger $K$, the larger $\mathcal{F}_{K,\bm{\lambda}}^{(n)}$. Theorem˜3.3 states that, over $\mathcal{F}_{K,\bm{\lambda}}^{(n)}$, the minimax risk grows linearly with $K$. Theorem˜3.3 confirms that the ESD quantifies the best possible performance of any estimator over the class $\mathcal{F}_{K,\bm{\lambda}}^{(n)}$.
Example 3.4.
For any given spectrum and $\bm{\theta}^{*}$, we can find the smallest $K^{*}$ such that $\bm{\theta}^{*}\in\mathcal{F}_{K^{*},\bm{\lambda}}^{(n)}$, over which the minimax risk is $\sigma^{2}K^{*}$. Below are several examples:
  • (1)
    If $\lambda_{i}=i^{-\beta}$ and $\sum_{i=1}^{d}\lambda_{i}^{-s}\theta_{i}^{*2}\leq R$ for some $\beta,s>0$,
    $\sigma^{2}K^{*}\;\asymp\;\min\!\left\{\sigma^{\frac{2s\beta}{1+s\beta}},\,d\sigma^{2}\right\}.$
  • (2)
    If $\theta_{i}^{*}=i^{-\alpha/2}$ for some $\alpha>1$ and $\{\lambda_{i}\}_{i=1}^{d}$ is decreasing, then
    $\sigma^{2}K^{*}\;\asymp\;\min\!\left\{\sigma^{2-\frac{2}{\alpha}},\,d\sigma^{2}\right\}.$
  • (3)
    If $d<\infty$, $\theta_{i}^{*}=i^{-1/2}$, and $\{\lambda_{i}\}_{i=1}^{d}$ is decreasing, then
    $\sigma^{2}K^{*}\;\asymp\;\min\!\left\{d\sigma^{2},\,\log\!\big(d\sigma^{2}/\log(d\sigma^{2})\big)\right\}.$
  • (4)
    If $d<\infty$, $0<\alpha<1$, $\theta_{i}^{*}=i^{-\alpha/2}$, and $\{\lambda_{i}\}_{i=1}^{d}$ is decreasing, then
    $\sigma^{2}K^{*}\;\asymp\;d\min\!\left\{d^{-\alpha},\,\sigma^{2}\right\}.$
Details and proofs are deferred to Section˜A.3.
In Setting (1), we may take $\sigma^{2}=\sigma_{0}^{2}/n$ to align with the approximation developed in Equation˜3, and then the upper bound becomes $\sigma_{0}^{2}\min\left(n^{-\frac{s\beta}{1+s\beta}},d/n\right)$, which matches the well-known optimal rate under the source condition and the polynomial eigen-decay condition in the case when $d=\infty$. When $d<\infty$, there is a phase transition around $d_{0}:=n^{\frac{1}{1+s\beta}}$: if $d\lesssim d_{0}$, the upper bound is $d\sigma_{0}^{2}/n$; if $d\gtrsim d_{0}$, the upper bound is the same as if $d=\infty$.
These examples suggest that the notion of ESD allows us not only to recover classical results but also to explore new settings where the classical framework is inapplicable. ∎
ESD is a population complexity measure for theoretical analysis rather than a training input. Comparing the ESD of the same signal across kernels indicates which kernel can generalize better for that signal. Section˜10.1 illustrates a sparse signal example where the ESD provides a quantitative comparison of two different spectra while the existing measures like $d_{\text{eff }}(\nu)$ and $k^{*}$ do not.
Discussion on estimator-guided measures.
The ESD equals the truncation point for the PC estimator where the squared tail bias first becomes comparable to (or less than) the estimation variance. It may seem that one can define an alternative complexity index by calibrating the bias and variance of any particular estimator. There are two reasons why we focus on the ESD.
First, Theorem˜3.3 clarifies that any complexity measure that characterizes minimax optimal rates for the classes we consider above must agree with ESD up to constant factors. Therefore, ESD is the canonical measure for alignment-sensitive minimax rates. We use ESD because it admits a clean analytic characterization and enables the study of the dynamic signal-kernel alignment in adaptive learning.
Second, such an estimator-guided index may reflect defects of the chosen procedure rather than intrinsic difficulty. A concrete example is ridge regression: due to the saturation effect, a ridge-guided index calibrated to ridge’s bias–variance balance will fail to describe the optimal estimation risk even in classical settings with source conditions. We rigorously formalize this point in Section˜5.
Comparisons to other alignment measures.
Alternative alignment measures exist. The cosine similarity-based kernel-target alignment yields generalization bounds [13, 14], but these bounds are typically too loose to explain fast rates in adaptive kernel methods. Recently, Barzilai and Shamir [7] extended benign-overfitting analyses [5, 40] to kernel ridge regression, which may encounter saturation effects that prevent optimal rates for overly smooth target functions.
Comparisons to other effective dimensions.
We discuss the differences between ESD and some well-known measures used in the classical analysis of spectral methods.
Zhang [52] introduces the effective dimension to quantify the complexity of any regularized method. For ridge regularization in Equation˜5, the effective dimension is defined as $d_{\mathrm{eff}}(\nu)=\sum_{j}\frac{\lambda_{j}}{\lambda_{j}+\nu}$ (see Proposition A.1 in Zhang [52]). $d_{\mathrm{eff}}(\nu)$ depends only on the spectrum $\bm{\lambda}$ and the regularization parameter $\nu$, but not on the signal $\bm{\theta}^{*}$ or the noise level $\sigma^{2}$. Consequently, the effective dimension is not suitable for measuring signal-spectrum alignment. Furthermore, the effective dimension, as a function of $\nu$, does not directly connect to any minimax risk.
In linear regression, Bartlett et al. [6] analyze the minimum-norm interpolator via the effective rank $r_{k}=\frac{\sum_{i>k}\lambda_{\pi_{i}}}{\lambda_{\pi_{k+1}}}$ (using the relationship in Equation˜3). They define the index $k^{*}=\min\{k\geq 0:\,\sigma^{2}r_{k}\geq b\}$ for some constant $b$ and establish risk bounds using $\sigma^{2}k^{*}$. $k^{*}$ differs from ESD in two important aspects: (i) $k^{*}$ does not involve the signal and thus cannot measure signal-spectrum alignment; and (ii) $k^{*}$ is tailored to the minimum-norm estimator and does not characterize the minimax risk over a class.
Both $d_{\text{eff }}(\nu)$ and $k^{*}$ are signal-agnostic: they depend on the spectrum $\bm{\lambda}$ (and either $\nu$ or $\sigma^{2}$) only, and therefore remain invariant under any change in the alignment between the signal and the kernel’s eigenfunctions. For instance, if adaptive training improves alignment by reordering the eigenfunctions to better align with the signal while preserving the set of eigenvalues, then both $d_{\text{eff }}(\nu)$ and $k^{*}$ are unchanged. In contrast, the ESD $d^{\dagger}\left(\sigma^{2};\bm{\theta}^{*},\bm{\lambda}\right)$ is signal-aware, because it is defined by the bias-variance crossing for the specific $\bm{\theta}^{*}$. As signal-kernel alignment improves, the ESD decreases. Consequently, the ESD can mechanistically explain the generalization benefits of adaptive kernel learning, while $d_{\text{eff }}(\nu)$ and $k^{*}$ cannot.
3.2 Span Profile
The definition of ESD explicitly depends on the noise level $\sigma^{2}$, which distinguishes it from other complexity measures in the literature. The dependence on $\sigma^{2}$ reflects the bias-variance trade-off nature of ESD: as $\sigma^{2}$ decreases, more coordinates can be unbiasedly estimated while controlling the overall variance, thereby more bias is removed. To focus on the alignment between a given signal $\bm{\theta}^{*}$ and a spectrum $\bm{\lambda}$, we examine the ESD by varying the noise level.
Definition 3.5.
We define the span profile of $\bm{\theta}^{*}$ w.r.t. the spectrum $\bm{\lambda}$ as $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}:\tau\mapsto d^{\dagger}(\tau;\bm{\theta}^{*},\bm{\lambda})$.
The span profile $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}$ is a well-defined object that depends only on $\bm{\theta}^{*}$ and the ordering of $\bm{\lambda}$, and it summarizes how $\sigma^{2}$ affects the ESD. Theorem˜3.2 suggests that for two spectra $\bm{\lambda}^{(1)}$ and $\bm{\lambda}^{(2)}$, we can compare their alignments with the signal by the ratio of $r(\tau)={\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(1)}}(\tau)}/{\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(2)}}(\tau)}$ for small $\tau$, because, if this ratio is very small (and in particular if the limit is 0 for $\tau\to 0$), then a kernel method using $\bm{\lambda}^{(1)}$ can achieve a smaller risk than one that uses $\bm{\lambda}^{(2)}$. Such comparisons are not as convenient in classical theory. See Section˜10 for more illustrations.
A closely related object is the trade-off function of $\bm{\theta}^{*}$ relative to $\bm{\lambda}$, which is defined for any $k\in[d]$ as
$\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\;=\;\frac{1}{k}\sum_{i=k+1}^{d}\left(\theta_{\pi_{i}}^{*}\right)^{2}\;=\;\frac{1}{k}\sum_{\,i:\,\lambda_{i}<\lambda_{\pi_{k}}\,}\left(\theta_{i}^{*}\right)^{2}.$ (9)
The quantity $\sigma^{-2}\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)$ equals the bias-variance ratio of the PC estimator using the $k$ leading coordinates. Properties of span profiles and trade-off functions are summarized as follows.
Proposition 3.6.
(1) Both $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}:\tau\mapsto[d]$ and $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}:[d]\mapsto[0,\infty)$ are nonincreasing. (2) For any $\tau$, it holds that $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\tau)=\min\{k\in[d]:~~\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\leq\tau\}$. (3) For two spectra $\bm{\lambda}^{(1)}$ and $\bm{\lambda}^{(2)}$, if $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}^{(1)}}(k)\leq\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}^{(2)}}(k)$ for all $k\in[d]$, then $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(1)}}(\tau)\leq\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(2)}}(\tau),\quad\forall\tau>0$.
Property (3) in Proposition˜3.6 suggests that the faster $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(\cdot)$ decreases, the better the spectrum $\bm{\lambda}$ aligns with the signal $\bm{\theta}^{*}$. In the extreme case where the ordering of $\lambda_{i}$ matches the ordering of $|\theta_{i}^{*}|^{2}$}, the decay of $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(\cdot)$ is the fastest, which leads to the most favorable span profile.
Extensions.
To save space, we defer the extensions to linear models and kernel regression to Sections˜8 and 9, respectively. For the kernel regression model in Equations˜1 and 2, we define the ESD of $f^{*}$ w.r.t. the kernel $\mathbf{k}$ and the effective noise variance $\sigma^{2}:=(\sigma_{0}^{2}+\|f^{*}\|_{\infty}^{2})/n$ as
$d^{\dagger}(\sigma^{2};f^{*},\mathbf{k})=\min\{k\in\mathbb{N}_{+}\cup\{\infty\}:\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\leq\sigma^{2}\}.$
4 Minimax Optimal Convergence Rates
When using the span profile to characterize the signal-spectrum alignment, it is of interest to establish the optimal convergence rates. Since the setting where $d=d_{n}$ grows along with $n$ has been studied in Theorem 3.3, we focus on the case where $d=\infty$ and the spectrum $\bm{\lambda}$ is given with ordering denoted by $\{\pi_{j}\}$ such that $\lambda_{\pi_{1}}>\lambda_{\pi_{2}}>\ldots$. For the convenience of the asymptotic analysis, we examine the span profile at $\tau=\sigma_{0}^{2}/n$, where $\sigma_{0}^{2}$ is fixed and $n$ enumerates $\mathbb{N}_{+}$.
We begin by defining a class of populations whose span profile is bounded by a sequence of quotas $\bm{K}=\{K_{n}\}_{n=1}^{\infty}$. This leads to the following class of parameters:
$\mathcal{F}_{\bm{K},\bm{\lambda}}:=\Bigl\{\bm{\theta}\in\mathbb{R}^{\infty}:\mathbf{D}_{\bm{\theta},\bm{\lambda}}\!\Bigl(\tfrac{\sigma_{0}^{2}}{n}\Bigr)\leq K_{n},\quad\forall~n\geq n_{0}\text{ for some }n_{0}\Bigr\}.$ (10)
For each $\bm{\theta}\in\mathcal{F}_{\bm{K},\bm{\lambda}}$, the sequence model in Equation˜4 with $\bm{\theta}^{*}=\bm{\theta}$ and $\sigma^{2}=\sigma_{0}^{2}/n$ will have an ESD no greater than $K_{n}$. For a sample $\bm{Z}^{(n)}$ from this sequence model and any estimator $\widehat{\bm{\theta}}$ based on $\bm{Z}^{(n)}$, we aim to determine the convergence rate of the following minimax risk:
$\inf_{\widehat{\bm{\theta}}}\sup_{\bm{\theta}\in\mathcal{F}_{\bm{K},\bm{\lambda}}}\mathcal{R}(\widehat{\bm{\theta}},\bm{\theta}).$ (11)
We emphasize that $\bm{K}$ is a model-class descriptor. It is not a parameter of the distribution, but rather describes a condition on the distribution. For example, the sparsity assumption in high-dimensional regression states that $\|\bm{\beta}\|_{0}\leq s$, so $s$ describes a class of distributions; yet $s$ is not a parameter of the distribution. Our minimax result requires a regularity condition on the quota sequence $\bm{K}$. Let $\bar{K}:=\sup\{K_{n}\}\in\mathbb{N}\cup\{\infty\}$. For any $k\in[\bar{K}]$, let $M_{k}:=\max\{n:K_{n}=k\}$ (the largest $n$ such that $K_{n}=k$).
Condition 4.1.
(1) $K_{n+1}-K_{n}\leq 1$ for all $n$ sufficiently large. (2) For all $k\in[\bar{K}]$, it holds that $(k+1)/M_{k+1}\leq k/M_{k}$.
˜4.1 ensures that $K_{n}$ does not grow faster than $n$, and the ratio sequence $\{k/M_{k}\}$ is nonincreasing. ˜4.1 is easily satisfied by common growth laws.
Example 4.2.
(1) Suppose $K_{n}\asymp n^{a}$ where $0<a<1$. For any $k$, we have $M_{k}\asymp k^{1/a}$. Since $k/k^{1/a}$ is decreasing, ˜4.1 holds.
(2) Suppose $K_{n}\asymp(\log n)^{b}$ where $b>0$. For any $k$, we have $M_{k}\asymp e^{k^{1/b}}$. Since $k/e^{k^{1/b}}$ is decreasing, ˜4.1 holds.
The next theorem provides a lower bound on the minimax risk in Equation˜11.
Theorem 4.3.
Suppose ˜4.1 holds for a quota sequence $\bm{K}=\{K_{n}\}_{n=1}^{\infty}$. Let $c_{0}=1/4$. If $\bm{Z}^{(n)}$ is drawn from the sequence model with $\bm{\theta}^{*}=\bm{\theta}$ and $\sigma^{2}=\sigma_{0}^{2}/n$, it holds that
$\inf_{\widehat{\bm{\theta}}}\sup_{\bm{\theta}\in\mathcal{F}_{\bm{K},\bm{\lambda}}}\mathcal{R}(\widehat{\bm{\theta}},\bm{\theta})\geq c_{0}\sigma_{0}^{2}\frac{K_{n}}{n}.$
Theorem˜4.3 shows that given a quota sequence $\bm{K}$, no estimator can, uniformly over the class $\mathcal{F}_{\bm{K},\bm{\lambda}}$, achieve a faster convergence rate of risk than $\sigma_{0}^{2}K_{n}/n$. On the other hand, Theorem˜3.2 (using $\sigma^{2}=\sigma_{0}^{2}/n$) provides a matched upper bound on the risk of the optimal PC estimator, which is $2\sigma_{0}^{2}K_{n}/n$. We thus conclude that the minimax optimal rate over $\mathcal{F}_{\bm{K},\bm{\lambda}}$ is $\sigma_{0}^{2}K_{n}/n$.
Since our theory does not invoke any source condition or eigenvalue-decay condition, it goes beyond the classical analysis in the literature. It suggests that the ESD is an essential quantity, and the span profile provides a useful characterization of the attainable error rate for spectral methods.
To see why our framework is more general, the next example presents a case where the minimax convergence rate is slower than the rate in the fixed-dimensional setting while being faster than the standard rate in classical infinite-dimensional settings.
Example 4.4.
Let $b\geq 1$ be a constant and $K_{n}=\lceil(\log n)^{1/b}\rceil$ for $n\in\mathbb{N}_{+}$. Suppose $\{\lambda_{j}\}_{j=1}^{\infty}$ is decreasing and $\theta^{*}_{j+1}=\sqrt{\,\sigma_{0}^{2}\left[je^{-j^{b}}-(j+1)e^{-(j+1)^{b}}\right]}$ for $j\geq 1$ and $\theta^{*}_{1}=0$. Then, $\bm{\theta}^{*}\in\mathcal{F}_{\bm{K},\bm{\lambda}}$ and the optimal rate is $\sigma_{0}^{2}\,(\log n)^{1/b}\,n^{-1}$. In contrast, the traditional convergence rate based on the source condition is $\sigma_{0}^{2}\,n^{-\alpha/(1+\alpha)}$ for arbitrary $\alpha>0$, which is not sharp.
Connecting adaptivity and generalization. Our span profile framework clarifies why adaptive machine learning methods often outperform classical fixed-kernel approaches. In classical methods with a fixed kernel spectrum $\bm{\lambda}^{(0)}$, the target signal $\bm{\theta}^{*}$ might exhibit poor alignment, resulting in a large span profile $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(0)}}$. Consequently, the signal resides in a class $\mathcal{F}_{\mathbf{K}^{(0)},\bm{\lambda}^{(0)}}$ with a large quota sequence $\mathbf{K}^{(0)}$, which implies high minimax risk. By contrast, adaptive methods modify the kernel during training. Successful adaptation improves the alignment by reducing the span profile $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(a)}}$ of the same signal w.r.t. the adapted kernel spectrum $\bm{\lambda}^{(a)}$. This adaptation places the signal in a class $\mathcal{F}_{\mathbf{K}^{(a)},\bm{\lambda}^{(a)}}$ with a smaller quota sequence $\mathbf{K}^{(a)}$, which implies lower minimax risk.
5 Why ridge cannot define an intrinsic complexity measure
This section explains why a complexity measure defined by using the ridge estimator as a guideline is not suitable for characterizing intrinsic generalization difficulty. The key issue is the classical saturation effect of ridge: ridge cannot fully exploit the smoothness property of signals that are sufficiently smooth, so its optimal bias–variance trade-off becomes insensitive to the true source class.
5.1 Ridge estimator and risk decomposition
We work with the same Gaussian sequence model as in Equation˜4. The ridge estimator can be written in the generic spectral form $\widehat{\theta}^{\mathrm{R},\nu}_{j}=\bigl(1-\psi_{\nu}^{\mathrm{R}}(\lambda_{j})\bigr)z_{j}$ with $\psi_{\nu}^{\mathrm{R}}(\lambda)=\nu/(\lambda+\nu)$, namely
$\widehat{\theta}^{\mathrm{R},\nu}_{j}=\frac{\lambda_{j}}{\lambda_{j}+\nu}\,z_{j}.$
Its mean squared error admits the standard bias–variance decomposition
$\mathcal{R}_{\mathrm{R}}(\nu)=\mathbb{E}\Bigl[\|\widehat{\bm{\theta}}^{\mathrm{R},\nu}-\bm{\theta}^{*}\|_{2}^{2}\Bigr]=B_{\mathrm{R}}(\nu)+V_{\mathrm{R}}(\nu),$
where the squared bias is given by
$B_{\mathrm{R}}(\nu)=\sum_{j=1}^{d}\Bigl(\psi_{\nu}^{\mathrm{R}}(\lambda_{j})\Bigr)^{2}(\theta^{*}_{j})^{2}=\sum_{j=1}^{d}\Bigl(\frac{\nu}{\lambda_{j}+\nu}\Bigr)^{2}(\theta^{*}_{j})^{2},$
and the variance is
$V_{\mathrm{R}}(\nu)=\sigma^{2}\sum_{j=1}^{d}\Bigl(1-\psi_{\nu}^{\mathrm{R}}(\lambda_{j})\Bigr)^{2}=\sigma^{2}\sum_{j=1}^{d}\Bigl(\frac{\lambda_{j}}{\lambda_{j}+\nu}\Bigr)^{2}.$
Before moving on, we introduce the ordering and a convenient reparametrization. Given the spectrum $\bm{\lambda}=\{\lambda_{j}\}_{j=1}^{d}$, let $\pi$ be the permutation that sorts the eigenvalues in decreasing order: $\lambda_{\pi_{1}}>\lambda_{\pi_{2}}>\cdots$ (assume no ties for simplicity), and write $\lambda_{(i)}:=\lambda_{\pi_{i}}$. For any $\nu\in(0,\lambda_{(1)}]$, define the index
$k_{\bm{\lambda}}(\nu)=\#\{j:\lambda_{j}\geq\nu\}.$
In words, $k_{\bm{\lambda}}(\nu)$ is the number of eigenvalues no smaller than $\nu$. By definition, we have $\lambda_{(k_{\bm{\lambda}}(\nu))}\geq\nu>\lambda_{(k_{\bm{\lambda}}(\nu)+1)}.$
5.2 Lower bound on the variance
We introduce a convenient lower bound for the variance term $V_{\mathrm{R}}(\nu)$.
For any $k\in[d]$, define
$\widetilde{N}(k,d;\bm{\lambda})=\frac{\sum_{i=k+1}^{d}\lambda_{(i)}^{2}}{k\,\lambda_{(k)}^{2}},$
and define
$\widetilde{V}(k):=\sigma^{2}\,k\bigl(1+\widetilde{N}(k,d;\bm{\lambda})\bigr)=\sigma^{2}\,\left(k+\frac{\sum_{i=k+1}^{d}\lambda_{(i)}^{2}}{\,\lambda_{(k)}^{2}}\right).$
Proposition 5.1.
Fix $\nu\in(0,\lambda_{(1)}]$ and let $k:=k_{\bm{\lambda}}(\nu)$. It holds that
$\frac{1}{4}\widetilde{V}(k)\leq V_{\mathrm{R}}(\nu).$
5.3 A lower bound on the ridge bias at small regularization
The saturation phenomenon is driven by the fact that for small $\nu$ the ridge bias behaves as
$\nu^{2}\sum_{j}(\theta^{*}_{j}/\lambda_{j})^{2},$
which involves the rescaled coefficients $\theta^{*}_{j}/\lambda_{j}$ and therefore can stop decreasing once the signal is smoother than ridge can exploit.
To be concrete, we place the following assumption.
Assumption 5.2.
There exist an index $i_{0}\in[d]$ and a constant $c_{r}>0$ such that
$\sum_{i=1}^{i_{0}}\frac{(\theta^{*}_{\pi_{i}})^{2}}{\lambda_{(i)}^{2}}\geq 4c_{r}.$
We have the following lower bound on the squared bias term $B_{\mathrm{R}}(\nu)$
Proposition 5.3.
Under ˜5.2, it holds for any $\nu\leq\lambda_{(i_{0})}$ that
$B_{\mathrm{R}}(\nu)\geq c_{r}\,\nu^{2}.$ (12)
For any $k$, we define the proxy for the bias as
$\widetilde{B}(k):=c_{r}\,\lambda_{(k)}^{2}.$
Note that $\widetilde{B}(k)$ is comparable to the lower bound $c_{r}\nu^{2}$ if $k=k_{\bm{\lambda}}(\nu)$ and $k\geq i_{0}$.
5.4 A ridge-guided complexity.
Suppose ˜5.2 holds. Combining the bias proxy $\widetilde{B}(k)$ and the variance proxy $\widetilde{V}(k)$, we consider the proxy risk
$\widetilde{R}(k)=\widetilde{B}(k)+\widetilde{V}(k)=c_{r}\,\lambda_{(k)}^{2}+\sigma^{2}\,k\bigl(1+\widetilde{N}(k,d;\bm{\lambda})\bigr).$
This leads to the ridge-guided analogue of the trade-off sequence used in the definition of ESD. Specifically, we define the ridge trade-off sequence
$\bar{h}(k;\bm{\theta}^{*},\bm{\lambda})=\frac{\lambda_{(k)}^{2}}{k\bigl(1+\widetilde{N}(k,d;\bm{\lambda})\bigr)}.$
Definition 5.4.
Define the ridge saturating dimension by
$d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})=\max\Bigl\{k:\ c_{r}\,\bar{h}(k;\bm{\theta}^{*},\bm{\lambda})>\sigma^{2}\Bigr\}.$
By construction, at $k=d^{\Delta}$ the two proxy terms are balanced up to constants:
$c_{r}\,\lambda_{(d^{\Delta})}^{2}>\sigma^{2}\,d^{\Delta}\bigl(1+\widetilde{N}(d^{\Delta},d;\bm{\lambda})\bigr),\qquad c_{r}\,\lambda_{(d^{\Delta}+1)}^{2}\leq\sigma^{2}\,(d^{\Delta}+1)\bigl(1+\widetilde{N}(d^{\Delta}+1,d;\bm{\lambda})\bigr).$
It turns out, we can use $d^{\Delta}$ to lower bound the risk of the ridge estimator. The key point is that $d^{\Delta}$ is the quantity that ridge cannot beat (up to constants) when $\nu$ ranges over the natural regime $\nu\leq\lambda_{(i_{0})}$.
Proposition 5.5.
Suppose ˜5.2 holds. For any integer $k\geq i_{0}$, it holds that
$\widetilde{R}(k)\geq\sigma^{2}\,d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2}).$
The proof is straightforward:
  • if $k\geq d^{\Delta}$, then $\widetilde{R}(k)\geq\widetilde{V}(k)\geq\sigma^{2}k\geq\sigma^{2}d^{\Delta}$;
  • if $k\leq d^{\Delta}$, then $\widetilde{R}(k)\geq\widetilde{B}(k)\geq\widetilde{B}(d^{\Delta})>\sigma^{2}d^{\Delta}$.
Theorem 5.6 (Lower bound for Ridge risk).
Suppose ˜5.2 holds. For any $\nu\in(0,\lambda_{(i_{0})}]$, let $k:=k_{\bm{\lambda}}(\nu)$. Then
$\mathcal{R}_{\mathrm{R}}(\nu)\geq\max\Bigl\{c_{r}\nu^{2},\ \frac{1}{4}\widetilde{V}(k)\Bigr\}\geq\frac{1}{4}\sigma^{2}\,d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2}).$
Consequently,
$\inf_{\nu\in(0,\lambda_{(i_{0})}]}~\mathcal{R}_{\mathrm{R}}(\nu)\geq\frac{1}{4}\sigma^{2}\,d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2}).$
5.5 Why this ridge-guided index fails to recover source-class structure
As discussed in the main text (see Theorem˜3.3), the minimax risk over the class
$\mathcal{F}_{K,\bm{\lambda}}^{(n)}=\Bigl\{\bm{\theta}\in\mathbb{R}^{d}:d^{\dagger}(\sigma^{2};\bm{\theta},\bm{\lambda})\leq K\Bigr\}$ (13)
scales as $\sigma^{2}K$ up to constants. Therefore, $d^{\dagger}$ captures the intrinsic difficulty of estimation.
By contrast, the risk of ridge estimator is lower bounded by $d^{\Delta}$, which depends on $\bm{\lambda}$ through $\lambda_{(k)}^{2}$ and $\widetilde{N}(k,d;\bm{\lambda})$, and it becomes insensitive to additional smoothness once the signal is beyond ridge’s qualification. In particular, we can use the ratio
$\frac{d^{\dagger}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})}{d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})}$
to study whether the saturation effect appears for any given signal and spectral. In particular, if this ratio is very small, Theorem˜9.6 and Theorem˜5.6 together suggests the risk of ridge estimator is much larger than the risk of PC estimator. We demonstrate this idea using the following example.
Example 5.7.
Suppose ˜5.2 holds and consider the classical polynomial spectrum that $\lambda_{(i)}=i^{-\beta}$ where the degree $\beta>1/2$. It can be showed that $\widetilde{N}(k,d;\bm{\lambda})$ is bounded by a constant for all $k$. Therefore, we have
$\bar{h}(k;\bm{\theta}^{*},\bm{\lambda})\asymp\frac{\lambda_{(k)}^{2}}{k}=\frac{k^{-2\beta}}{k}=k^{-(1+2\beta)}.$
Given any $\sigma^{2}>0$, we have
$d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})\gtrsim(\sigma^{2})^{-1/(1+2\beta)}.$
Now consider a classical source condition: there exists $R>0$ and $s>0$ such that
$\sum_{i=1}^{d}\lambda_{i}^{-s}(\theta_{i}^{*})^{2}\leq R.$
This implies
$h(k;\bm{\theta}^{*},\bm{\lambda})=\frac{1}{k}\sum_{i=k+1}^{d}(\theta^{*}_{\pi_{i}})^{2}\leq\frac{R}{k}\,\lambda_{(k)}^{s}=R\,k^{-(1+s\beta)},$
and hence
$d^{\dagger}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})\lesssim(\sigma^{2})^{-1/(1+s\beta)}.$
If $s>2$ (i.e., the signal is smoother than ridge’s qualification), then
$\frac{d^{\dagger}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})}{d^{\Delta}(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})}\ \lesssim\ (\sigma^{2})^{\frac{1}{1+2\beta}-\frac{1}{1+s\beta}}\ \to 0\qquad\text{as }\sigma^{2}\to 0,$
so the ridge-guided index $d^{\Delta}$ is asymptotically much larger than the intrinsic complexity $d^{\dagger}$. Equivalently, ridge is forced to pay a risk on the order of $\sigma^{2}d^{\Delta}$, even though the minimax rate over the same source class continues to improve with $s$ and is on the order of $\sigma^{2}d^{\dagger}$. ∎
The theory developed here and the above example demonstrate that when the signal is sufficiently smooth, ridge cannot translate that additional smoothness into a smaller optimal bias term, and the resulting ridge-guided index complexity index $d^{\Delta}$ becomes defective to characterize the optimal risk. This is why it is unsuitable as a replacement for ESD in our rate statements.
6 Adaptive Eigenvalues via Over-parameterized Gradient Flow
This section will investigate the benefits of learning eigenvalues via over-parameterized gradient flow (OP-GF) in sequence models [29] through the lens of ESDs. For more general adaptive kernels with evolving eigenfunctions, Section˜10.2 outlines how ESD can be used as an explanatory framework and we provide supporting evidence via a linear network experiment. Analyzing eigenfunction evolution technically more difficult than eigenvalues evolution and is left for future work.
Inspired by the over-parameterized nature of deep neural networks, Li and Lin [29] parameterized $\theta_{j}=a_{j}b_{j,1}\cdots b_{j,D}\beta_{j}$, where $D$ stands for the number of layers and $(a_{j},b_{j,i},\beta_{j})$ are parameters to be learned. The gradient flow w.r.t. the empirical loss $L=\frac{1}{2}\sum_{j}(\theta_{j}-y_{j})^{2}$ is given by
$\dot{a}_{j}=-\nabla_{a_{j}}L,$ $\quad\dot{b}_{j,i}=-\nabla_{b_{j,i}}L~~(i\in[D]),\quad\dot{\beta}_{j}=-\nabla_{\beta_{j}}L,$ (14)
$a_{j}(0)=\lambda_{j}^{1/2},$ $\quad b_{j,i}(0)=b_{0}>0,~~(i\in[D]),\quad\beta_{j}(0)=0,~~j\in[d],$
where $\lambda_{j}$’s are the initial eigenvalues and $b_{0}$ is the common initialization of all $b_{j,i}$. At time $t$, the learned eigenvalues are given by $\tilde{\lambda}_{j}(t)=(a_{j}(t)b_{j,1}(t)\cdots b_{j,D}(t))^{2}$ and the OP-GF estimates are $\hat{\theta}_{j}^{OP}(t)=\tilde{\lambda}^{\frac{1}{2}}_{j}(t)\beta_{j}(t)$ for $j\in[d]$. Li and Lin [29] consider infinite-dimensional sequence models with a polynomial decay condition on the initial eigenvalues and establish upper bounds on the risk of the OP-GF estimator with proper early stopping.
Here we study the dynamics of eigenvalues in OP-GF and how it changes the ESD. At time $t$, the learned eigenvalues are $\tilde{\bm{\lambda}}(t):=(\tilde{\lambda}_{j}(t))_{j\in[d]}$, and the ESD is $d^{\dagger}(t)=d^{\dagger}(\sigma^{2};\bm{\theta}^{*},\tilde{\bm{\lambda}}(t))$. We aim to show that under some regularity conditions, OP-GF can adjust the ordering of eigenvalues $\tilde{\bm{\lambda}}(t)$ to reduce the ESD $d^{\dagger}(t)$, which leads to a better signal-spectrum alignment.
We begin with some notations for the sequence model in Equation˜4. We focus on the large-sample case where $\sigma^{2}=\frac{\sigma_{0}^{2}}{n}$ and $\sigma_{0}=1$ without loss of generality. Denote $\tilde{d}=\sum_{i=1}^{d}\lambda_{i}$ (i.e., sum of initial eigenvalues). Let $\pi^{-1}_{t}(i)$ denote the rank of $\tilde{\lambda}_{i}(t)$ at time $t$.
Assumption 6.1.
We assume (1) Each noise $\xi_{j}$ in Equation˜4 is sub-Gaussian with variance proxy bounded by $C_{\text{proxy}}\sigma^{2}$.    (2) Let $\varepsilon=2C_{\text{proxy}}^{-1/2}n^{-1/2}\sqrt{\ln n\tilde{d}\cdot\ln n}$ and $\varepsilon^{\prime}=2C_{\text{proxy}}^{-1/2}n^{-1/2}\sqrt{\ln n}$. Define $S:=\{j\in[d]:|\theta_{j}^{*}|>\varepsilon\}$. We have $|S|\leq n$.    (3) $\inf_{j\in S}\lambda_{j}>n^{-\delta}$ for some $\delta\in(0,1)$.
Theorem 6.2.
Suppose that ˜6.1 holds and the initialization in Equation˜14 is $b_{0}=c_{B}D^{\frac{D+1}{D+2}}\varepsilon^{\frac{1}{D+2}}$. Define $t_{2}=C\cdot D^{\frac{D}{D+2}}(\varepsilon)^{-\frac{2D+2}{D+2}}$. There exist some constants $c$, $C$, $C_{M}$, $C_{max}$, $C_{\eta}$, $c_{\eta}$, $c_{B}$, and $c^{\prime}$, such that with probability larger than 14/n1-4/n, we have
$d^{\dagger}(t_{2})\leq d^{\dagger}(t_{1})$
for any $t_{1}\in[0,t_{2})$ if the followings hold:
  1. 1.
    For any $j\in S$, we have $M\leq|\theta^{*}_{j}|$, where $M:=C_{M}\varepsilon$;
  2. 2.
    For any $j\in S^{c}$, we have $|\theta^{*}_{j}|\leq\tilde{\sigma}$, where $\tilde{\sigma}=c^{\prime}\varepsilon$.
  3. 3.
    For any $i,j\in S$, let $\eta_{i,j}:=|\theta_{i}^{*}|-|\theta_{j}^{*}|$. At least one of the followings hold: (a) $\eta_{i,j}\leq 0$, (b) $\eta_{i,j}\geq C_{\eta}\varepsilon$ and $|\theta_{i}^{*}|\leq C_{max}M$, or (c) $\frac{|\theta_{i}^{*}|}{|\theta_{j}^{*}|}>(1+\frac{c_{\eta}}{D})$.
  4. 4.
    At time $t_{1}$, define two subsets of $S^{c}$: $A_{1}:=\{i\in S^{c}:\pi_{t_{1}}^{-1}(i)<d^{\dagger}(t_{1}),\lambda_{i}<c\cdot D^{-\frac{D}{D+2}}\cdot M^{\frac{2}{D+2}}\}$ and $B_{1}:=\{i\in S^{c}:\pi_{t_{1}}^{-1}(i)>d^{\dagger}(t_{1})\}$, and define a subset of $S$: $B_{2}:=\{i\in S:\pi_{t_{1}}^{-1}(i)>d^{\dagger}(t_{1})\}$. It holds that $|B_{2}|+\min\left[(|A_{1}|-|B_{2}|)_{+},|B_{1}|\right]\leq|B_{2}|(C_{M}/c^{\prime})^{2}$.
Theorem˜6.2 shows that OP-GF reduces the ESD. The conditions ensure that strong signals with small initial eigenvalues have larger adapted eigenvalues than weak signals.
7 Numerical experiments
Data Generation. We utilize the misalignment setting in Li and Lin [29] to specify a $d$-dimensional sequence model. We fixed the eigen-decay rate $\gamma>0$, the signal decay rate $p>0$, and the number of nonzero signals $J$. Given any misalignment parameter $q\geq 1$, we set eigenvalues as $\lambda_{j}=j^{-\gamma},j\in[d]$, and set the true nonzero parameters as $\theta_{\ell(j)}^{*}=C\cdot j^{-\frac{p+1}{2}}$, where $\ell(j)=[j^{q}]$ and $j\leq J$. Here all other elements of $\theta_{j}^{*}$ are zero and $d\geq J^{q}$ so $\|\bm{\theta}^{*}\|_{0}=J$. The observations are sampled as $y_{i}\sim N(\theta_{i}^{*},\sigma^{2})$. This setting provides a flexible way to control the alignment between the signal structure and the spectrum. When $q=1$, the ordering of $\bm{\theta}^{*}$ aligns perfectly with the ordering of $\bm{\lambda}$. As $q$ increases, more nonzero elements of $\bm{\theta}^{*}$ are located on the tail where the eigenvalues are smaller, and more large eigenvalues are associated with zero signals, creating a worse signal-spectrum alignment.
Figure 1: Evolution of span profiles during the training of an over-parameterized gradient flow. The misalignment level $q$ varies from $1$ to $3$. Fixed parameters are $n=10000$, $\sigma_{0}=1$, $d=5000$, $J=15$, $p=2.5$, and $\gamma=1$.
Evolution of Span Profile
The first experiment visualizes the span profile of the signal w.r.t. the learned spectrum at various stages in the OP-GF process with $D=0$. Given a sample, we approximate the gradient flow in Equation˜14 by discrete-time gradient descent and obtain the solution $\{(a_{j}(t),\beta_{j}(t))_{t\geq 0}:j\in[d]\}$. The trained eigenvalue sequence $\tilde{\bm{\lambda}}(t)$ at time $t$ is given by $\tilde{\lambda}_{j}(t)=a_{j}^{2}(t)$ for $j\in[d]$. Here we focus on time points before the optimal stopping time. Figure 1 illustrates the evolution of the span profile w.r.t. the learned spectrum for different training times $t$ and various values of $q$.
When $q=1$ (Top-Left panel), the span profiles at different training times $t$ are nearly identical. This is because the initial spectrum already aligns perfectly with the signal and there is no room for improvement. For $q>1$ (Top-Right, Bottom-Left, Bottom-Right panels), we observe that as the training time $t$ increases 0 to $80$, the span profile shifts downwards. This suggests that the training process refines the alignment between the spectrum and the signal. In addition, the reduction in the span profile is more significant for $q=3$ compared to $q=1.5$, because $q=3$ corresponds to a greater initial misalignment between the signal and the spectrum, rendering the improvement from OP-GF more substantial.
Evolution of ESD and Estimation Error of PC Estimators
We next empirically investigate the evolution of the ESD $d^{\dagger}$ and the estimation error as well as the impact of layers $D$. At any time $t$, we compute the ESD $d^{\dagger}(t)$ based on the learned eigenvalue sequence $\tilde{\bm{\lambda}}(t)$ and also the PC estimate $\widehat{\bm{\theta}}(t)$ based on $\tilde{\bm{\lambda}}(t)$, with number of components determined by $d^{\dagger}(t)$. Theorems˜3.2 and 3.3 suggest that the PC estimator tuned by the ESD can achieve the minimax risk rate, so we expect $\widehat{\bm{\theta}}(t)$ to perform well.
Figure 2: Averaged squared error of the tuned PC estimator and ESD as a function of the training time. Each average is computed based on 20 replications and each error bar represents a standard deviation.
The empirical evaluation involved 20 Monte Carlo repetitions. Figure˜2 displays the averaged $d^{\dagger}(t)$ and the averaged estimation error of $\widehat{\bm{\theta}}(t)$ in Figure˜2 as a function of training time $t$. We observe that both the ESD and the squared error of the tuned PC estimator exhibit a general decay trend over training time $t$. Furthermore, for the shallow model with $D=0$ (with no $b_{i,j}$ parameters), the initial decrease in ESD and MSE occurs earlier compared to the deeper models with $D=1$ or $D=3$. However, with sufficient training iterations, the deeper models with $D=1$ or $D=3$ can achieve lower ESD values than the shallow model with $D=0$. These findings suggest that increased model depth ($D>0$) may facilitate a better adaptation of the spectrum, and thus lead to lower estimation error. This observation offers a perspective on the benefits of depth in spectral learning, but a comprehensive study for general models is left for future research.
8 Extension to Correlated Noise and Fixed-Design Linear Model
This section extends the concepts of ESD and span profile, developed in Section˜3 for the sequence model, to the setting of fixed-design linear regression. In addition, we demonstrate how the minimax optimal prediction risk in this setting can be characterized using the span profile, paralleling the analysis in Section˜4.
8.1 Strategy of reduction
Before introducing the linear model, it is helpful to outline our general transformation strategy in the context of a sequence model with correlated noise. Suppose $d\in\mathbb{N}_{+}$ and the observations are
$\mathbf{Z}=\bm{\theta}^{\star}+\bm{\xi},\quad\bm{\xi}\sim\mathcal{N}_{d}\left(0,\sigma^{2}\mathbf{\Sigma}_{\xi}\right)$ (15)
where $\mathbf{\Sigma}_{\xi}\in\mathbb{R}^{d\times d}$ is known, symmetric, and positive definite. For a correlated sequence model, it is usually of interest to measure the estimation error using the squared Mahalanobis distance defined as $L(\bm{\theta}^{\star},\widehat{\bm{\theta}}^{\star})=(\widehat{\bm{\theta}}^{\star}-\bm{\theta}^{\star})^{\top}\mathbf{\Sigma}_{\xi}^{-1}(\widehat{\bm{\theta}}^{\star}-\bm{\theta}^{\star})$.
Let $\mathbf{L}=\mathbf{\Sigma}_{\xi}^{-1/2}$ be a symmetric square root of $\mathbf{\Sigma}_{\xi}^{-1}$. Define the whitened observation and transformed parameters as
$\mathbf{\tilde{Z}}=\mathbf{L}\mathbf{Z},\quad\bm{\tilde{\theta}}^{\star}=\mathbf{L}\bm{\theta}^{\star}.$
It follows that $\mathbf{\tilde{Z}}=\bm{\tilde{\theta}}^{\star}+\bm{\tilde{\xi}}$, where $\bm{\tilde{\xi}}=\mathbf{L}\bm{\xi}\sim\mathcal{N}_{d}\left(0,\sigma^{2}\mathbf{I}_{d}\right)$. Accordingly, any estimator $\widehat{\bm{\theta}}$ for $\bm{\theta}^{\star}$ is equivalent to the estimator $\bm{\widehat{\tilde{\theta}}}:=\mathbf{L}\widehat{\bm{\theta}}$ for $\bm{\tilde{\theta}}^{\star}$, whose squared loss is $\|\bm{\widehat{\tilde{\theta}}}-\bm{\tilde{\theta}}^{\star}\|^{2}=(\widehat{\bm{\theta}}^{\star}-\bm{\theta}^{\star})^{\top}\mathbf{\Sigma}_{\xi}^{-1}(\widehat{\bm{\theta}}^{\star}-\bm{\theta}^{\star})=L(\bm{\theta}^{\star},\widehat{\bm{\theta}}^{\star})$.
Therefore, the transformed model is equivalent to the standard sequence model with uncorrelated noise in Equation˜4 and the estimation is equivalent to the estimation therein. Consequently, the ESD and span profile for the model in Equation˜15 can be naturally defined using the original definitions for the transformed model. Specifically, for the correlated-noise model, we define the ESD w.r.t. $\Sigma_{\xi}$ as
$d_{\mathbf{\Sigma}_{\xi}}^{\dagger}\left(\tau;\bm{\theta}^{\star}\right):=d^{\dagger}\left(\tau;\bm{\tilde{\theta}}^{\star}=\mathbf{L}\bm{\theta}^{\star}\right).$
Note that for the risk $\mathbb{E}L(\bm{\theta}^{\star},\widehat{\bm{\theta}}^{\star})$, our minimax risk characterization still applies, i.e., the minimax risk scales as $K\sigma^{2}$ across all distributions whose ESD $d_{\mathbf{\Sigma}_{\xi}}^{\dagger}\left(\sigma^{2};\bm{\theta}^{\star}\right)$ is bounded by $K$.
The relationship between Euclidean distance and Mahalanobis distance satisfies that
$\lambda_{\min}(\mathbf{\Sigma}_{\xi})L(\bm{\theta}^{\star},\widehat{\bm{\theta}}^{\star})\leq\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}^{\star}\|^{2}\leq\lambda_{\max}(\mathbf{\Sigma}_{\xi})L(\bm{\theta}^{\star},\widehat{\bm{\theta}}^{\star}),$
so the minimax risk in terms of $\mathbb{E}\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}^{\star}\|^{2}$ can still be characterized sharply when the condition number of $\mathbf{\Sigma}_{\xi}$ is bounded.
This strategy of reducing a complex model to a simple model will be used in our analysis of the linear model and in RKHS regression in Section˜9.
8.2 Linear model
Consider the following fixed-design linear regression model:
$\mathbf{Y}=\mathbf{X}\bm{\beta}^{*}+\bm{\epsilon},$ (16)
where $\mathbf{Y}\in\mathbb{R}^{n}$ is the vector of observations, $\mathbf{X}\in\mathbb{R}^{n\times p}$ is the fixed-design matrix of rank $r\leq\min(n,p)$, $\bm{\beta}^{*}\in\mathbb{R}^{p}$ is the unknown vector of true coefficients, and $\bm{\epsilon}\in\mathbb{R}^{n}$ is the noise vector. We assume the components of $\bm{\epsilon}$ are uncorrelated with mean zero and variance $\sigma_{0}^{2}$. For this model, we consider the loss (in-sample prediction error) $\mathcal{L}(\widehat{\bm{\beta}};\bm{\beta}^{*})=\frac{1}{n}\|\mathbf{X}\left(\widehat{\bm{\beta}}-\bm{\beta}^{*}\right)\|^{2}$ and risk $\mathcal{R}(\widehat{\bm{\beta}};\bm{\beta}^{*})=\mathbb{E}\mathcal{L}(\widehat{\bm{\beta}};\bm{\beta}^{*})$. For random design linear regression, we treat it as a special case of RKHS regression and discuss it in Section˜9.
To connect this model to the sequence model analysis presented earlier, we utilize the Singular Value Decomposition (SVD) of the design matrix $\mathbf{X}$ as follows:
$\frac{1}{\sqrt{n}}\mathbf{X}=\mathbf{U}\mathbf{S}\mathbf{V}^{\top},$ (17)
where $\mathbf{U}\in\mathbb{R}^{n\times n}$ and $\mathbf{V}\in\mathbb{R}^{p\times p}$ are orthogonal matrices, and $\mathbf{S}\in\mathbb{R}^{n\times p}$ is a rectangular diagonal matrix with non-negative singular values $s_{1}\geq s_{2}\geq\dots\geq s_{r}>0$ on its diagonal, and $s_{j}=0$ for $j>r$.
For any matrix $\mathbf{A}$ and subsets $R$ and $T$, we write $\mathbf{A}_{\cdot,R}$ for the submatrix formed by the columns of $\mathbf{A}$ with indices in $R$, and write $\mathbf{A}_{T,\cdot}$ for the submatrix formed by the rows of $\mathbf{A}$ with indices in $T$.
Multiplying the model in Equation˜16 by $\frac{1}{\sqrt{n}}\mathbf{U}_{\cdot,[r]}^{\top}$, we obtain a $r$-dimensional transformed model:
$\mathbf{Z}=\bm{\theta}^{*}+\bm{\xi},$ (18)
where we have defined $\mathbf{Z}=\frac{1}{\sqrt{n}}\mathbf{U}_{\cdot,[r]}^{\top}\mathbf{Y}$, $\bm{\theta}^{*}=\frac{1}{\sqrt{n}}\mathbf{U}_{\cdot,[r]}^{\top}\mathbf{X}\bm{\beta}^{*}=\mathbf{S}_{[r],\cdot}\mathbf{V}^{\top}\bm{\beta}^{*}$, and $\bm{\xi}=\frac{1}{\sqrt{n}}\mathbf{U}_{\cdot,[r]}^{\top}\bm{\epsilon}$. Since $\mathbf{U}$ is orthogonal, the transformed noise vector $\bm{\xi}$ still has uncorrelated components with mean zero and variance $\sigma^{2}:=\sigma_{0}^{2}/n$.
The transformed model in Equation˜18 is analogous to the sequence model in Equation˜4, where the signal is $\bm{\theta}^{*}$ and the noise variance for each component is $\sigma^{2}$. The “spectrum” relevant to this problem is derived from the singular values of $\mathbf{X}$. Specifically, we define the eigenvalues as $\lambda_{j}=s_{j}^{2}$ for $j=1,\dots,r$, and $\lambda_{j}=0$ for $j>r$. Let $\{\pi_{k}\}_{k=1}^{r}$ denote the indices corresponding to the eigenvalues sorted in descending order, $\lambda_{\pi_{1}}\geq\lambda_{\pi_{2}}\geq\dots\geq\lambda_{\pi_{r}}>0$.
For any estimator $\widehat{\bm{\beta}}$ for the linear model Equation˜16, define $\widehat{\bm{\theta}}=\mathbf{S}_{[r],\cdot}\mathbf{V}^{\top}\widehat{\bm{\beta}}$. We can then write the prediction risk as $n^{-1}\mathbb{E}\|\mathbf{X}\widehat{\bm{\beta}}-\mathbf{X}\bm{\beta}^{*}\|^{2}=\mathbb{E}\|\mathbf{U}\mathbf{S}\mathbf{V}^{\top}\widehat{\bm{\beta}}-\mathbf{U}\mathbf{S}\mathbf{V}^{\top}\bm{\beta}^{*}\|^{2}=\mathbb{E}\|\widehat{\bm{\theta}}-\bm{\theta}^{*}\|^{2}$. Conversely, given an estimator $\widehat{\bm{\theta}}$ for the sequence model Equation˜18, we can define $\widetilde{\bm{\beta}}=\mathbf{V}\mathbf{S}^{\dagger}\widehat{\bm{\theta}}$ where $\mathbf{S}^{\dagger}\in\mathbb{R}^{p\times r}$ is a diagonal matrix whose diagonal elements are $\{1/s_{j}\}_{j\in[r]}$. It is easy to check that $\mathbf{S}_{[r],\cdot}\mathbf{V}^{\top}\widetilde{\bm{\beta}}=\widehat{\bm{\theta}}$. Therefore, we establish an equivalence between the model Equation˜18 and the model Equation˜16.
The usual ridge regression estimator for the linear model Equation˜16 is given by
$\widehat{\bm{\beta}}_{\nu}=\bigl(\mathbf{X}^{\top}\mathbf{X}+\nu\,\mathbf{I}_{p}\bigr)^{-1}\mathbf{X}^{\top}\mathbf{Y},$
which transforms into
$\widehat{\bm{\theta}}_{\nu}=\mathbf{S}_{[r],\cdot}\mathbf{V}^{\top}\widehat{\bm{\beta}}_{\nu}=\bigl(\mathbf{I}_{r}-\psi_{\nu}(\operatorname{Diag}\left(\lambda_{1},\ldots,\lambda_{r}\right))\bigr)\mathbf{Z},\qquad\text{ where }\psi_{\nu}(\lambda)=\frac{1}{\lambda/\nu+1}.$
In the above expression, we have used the identity that $\mathbf{S}_{[r],\cdot}\mathbf{S}_{[r],\cdot}^{\top}=\operatorname{Diag}\left(s_{1}^{2},\ldots,s_{r}^{2}\right)$ and $\psi_{\nu}(\cdot)$ is applied element-wise. If we replace $\psi_{\nu}(\lambda)$ by other functions as discussed in Section˜3, we recover other spectral methods.
8.3 ESD for linear models
We can now adapt the definitions from Section˜3 to linear models.
Definition 8.1 (ESD for Linear Regression).
Suppose the SVD of the design matrix $\mathbf{X}$ is given in Equation˜17. The Effective Span Dimension (ESD) of $\bm{\beta}^{*}$ with respect to the design $\mathbf{X}$ and the per component variance $\sigma_{0}^{2}/n$ is defined as
$d^{\dagger}=d^{\dagger}(\sigma_{0}^{2}/n;\bm{\beta}^{*},\mathbf{X})=\min\{k\in[r]:\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\leq\sigma_{0}^{2}/n\},$
where $\bm{\theta}^{*}=\mathbf{S}_{\cdot,[r]}\mathbf{V}^{\top}\bm{\beta}^{*}$ and $\lambda_{j}=s_{j}^{2}$.
The Principal Component Regression (PCR) estimator for $\bm{\beta}^{*}$ corresponds to the Principal Component (PC) estimator in the transformed space Equation˜18. Specifically, for any $k\in[r]$, define $\widehat{\bm{\beta}}^{\operatorname{PC},k}=\frac{1}{\sqrt{n}}\mathbf{V}\mathbf{S}_{k}^{\dagger}\mathbf{U}_{\cdot,[r]}^{\top}\mathbf{Y}$, where $\mathbf{S}_{k}^{\dagger}\in\mathbb{R}^{p\times r}$ is a diagonal matrix whose diagonal elements are $\{\frac{1}{s_{j}}\mathbf{1}_{\{s_{j}\geq s_{\pi_{k}}\}}\}$. In the $\mathbf{Z}$ space, this means
$\hat{\theta}_{j}^{\operatorname{PC},k}=\mathbf{1}\{s_{j}\geq s_{\pi_{k}}\}Z_{j},~~j\in[r].$ (19)
Analogous to Theorem˜3.2, the minimal prediction risk achievable by PCR over $k$ is characterized by the ESD.
Proposition 8.2 (Optimal PCR Prediction Risk).
Let $\widehat{\bm{\beta}}^{\operatorname{PC},k}$ be the PCR estimator using the first $k$ principal components. Let $\mathcal{R}_{*}^{\operatorname{PC}}$ be the minimal possible prediction risk over $k\in[r]$, i.e., $\mathcal{R}_{*}^{\operatorname{PC}}=\min_{k\in[r]}\mathcal{R}(\widehat{\bm{\beta}}^{\operatorname{PC},k};\bm{\beta}^{*})$. It holds that
$(d^{\dagger}-1)\sigma_{0}^{2}/n\leq\mathcal{R}_{*}^{\operatorname{PC}}\leq 2d^{\dagger}\sigma_{0}^{2}/n,$
where $d^{\dagger}=d^{\dagger}(\sigma_{0}^{2}/n;\bm{\beta}^{*},\mathbf{X})$ is the ESD defined in Definition˜8.1.
Proposition˜8.2 directly follows from Theorem˜3.2 and its proof is omitted. This result shows that the optimal prediction risk for PCR is determined by the ESD $d^{\dagger}$, which measures the effective number of principal components needed to balance the bias-variance trade-off.
We can further extend the minimax analysis from Theorem˜3.3. Let $K$ be a quota on ESD. Define a class of coefficient vectors based on this quota:
$\mathcal{B}_{K}^{(n)}=\left\{\bm{\beta}^{*}\in\mathbb{R}^{p}:d^{\dagger}(\sigma_{0}^{2}/n;\bm{\beta}^{*},\mathbf{X})\leq K\right\},$ (20)
This class contains signals whose ESD relative to the design $\mathbf{X}$ is controlled by $K$. We can establish the minimax optimal rate for prediction over this class.
Theorem 8.3 (Minimax Prediction Risk for Linear Regression).
Suppose $K\leq r$. For the linear model Equation˜16 with noise variance $\sigma_{0}^{2}$, the minimax prediction risk over the class $\mathcal{B}_{K}^{(n)}$ defined in Equation˜20 satisfies:
$\inf_{\widehat{\bm{\beta}}}\sup_{\bm{\beta}^{*}\in\mathcal{B}_{K}^{(n)}}\mathcal{R}(\widehat{\bm{\beta}};\bm{\beta}^{*})\asymp\sigma_{0}^{2}\frac{K}{n}.$
The proof of Theorem˜8.3 is essentially the same as that of Theorem˜3.3 and is omitted.
Through this extension, the span profile framework connects the optimal prediction performance in fixed-design linear regression to the alignment between the signal structure (transformed via the design matrix) and the spectrum derived from the design matrix’s singular values.
8.4 Numerical illustration
This section illustrates the ESD in fixed-design linear models in two examples. Throughout, we fix the noise variance at $\sigma_{0}^{2}=1$, the sample size at $n=300$, and the dimension at $p=400$.
Experimental setup
The baseline design matrix $\mathbf{X}_{0}$ is randomly generated with covariance matrix $\Sigma=\operatorname{Diag}\{\lambda_{j}\}_{j\in[p]}$ and then held fixed. We consider two cases:
  1. 1.
    Geometric decay spectrum and polynomial decay signal: $\lambda_{j}\propto 0.95^{\,j}$ and $\beta^{*}_{j}=j^{-0.2}$;
  2. 2.
    Logarithmic decay spectrum and signal: $\lambda_{j}=1/\log(j+1)$ with $\beta^{*}_{j}=1/\log(j+1)$.
The response will be generated from $\bm{Y}=\mathbf{X}_{0}\bm{\beta}^{*}+\bm{\epsilon}$ with random noise $\bm{\epsilon}$.
We are interested in the ESD and the minimum risk for different transformations of the design matrix. For this purpose, we introduce a class of non-orthogonal column transformations indexed by $\alpha>0$ as follows:
$\mathbf{A}(\alpha)=\operatorname{diag}\bigl(\exp\{\alpha\,t_{j}\}\bigr),\quad t_{j}=(j-1)/(p-1)-1/2,\quad j\in[p].$
The transformed design $\mathbf{X}(\alpha)=\mathbf{X}_{0}\mathbf{A}(\alpha)$ and the correspondingly transformed coefficient vector is $\bm{\beta}(\alpha)=\mathbf{A}(\alpha)^{-1}\bm{\beta}^{*}$. These transformations will change the ordering of the spectrum, from well-aligned to misaligned. We are interested in the following at each $\alpha$:
  • Effective Span Dimension: $d^{\dagger}(\alpha)=d^{\dagger}\bigl(\sigma_{0}^{2}/n;\,\bm{\beta}(\alpha),\mathbf{X}(\alpha)\bigr)$;
  • Minimal PCR risk: $\mathcal{R}_{*}(\alpha)=\min_{k}\mathbb{E}\bigl[n^{-1}\|\mathbf{X}(\alpha)\widehat{\bm{\beta}}_{k}-\mathbf{X}(\alpha)\bm{\beta}(\alpha)\|^{2}\bigr]$, where $\widehat{\bm{\beta}}_{k}$ is the $k$-component principal-component estimator based on $(\bm{Y},\mathbf{X}(\alpha))$.
Figure 3: Oracle PCR risk versus Effective Span Dimension for (a) geometric eigen-decay and (b) logarithmic eigen-decay. The dashed line plots $\text{Risk}\times n/\sigma_{0}^{2}$; the solid line is $d^{\dagger}(\alpha)$. The risk is computed based on 20 replications and the error bar represents the standard deviation.
Figure 3 plots $d^{\dagger}(\alpha)$ (solid) and the rescaled oracle risk defined as $n\mathcal{R}_{*}(\alpha)/\sigma_{0}^{2}$ (dashed) against $\alpha$. The two curves coincide over the entire path, which empirically verifies the bound in Proposition˜8.2 that $\mathcal{R}_{*}(\alpha)\asymp\frac{\sigma_{0}^{2}}{n}\,d^{\dagger}(\alpha)$. As $\alpha$ grows, the diagonal stretch $\mathbf{A}(\alpha)$ shifts signal energy towards directions that carry smaller singular values. This raises $d^{\dagger}$, and consequently, the achievable risk.
This experiment illustrates that ESD, rather than raw spectral decay, is the pivotal measure governing learnability.
9 Extension to RKHS Regression
This section extends the concepts of Effective Span Dimension (ESD) and span profile, developed in Section˜3, to the setting of RKHS regression. Since our goal is to develop a population-level complexity measure, we focus on the simple case where the eigenfunctions of the kernel are fully known and computable, and leave a thorough analysis for future studies.
9.1 RKHS regression
We recall the standard random-design kernel regression model from Section˜2:
$y_{i}=f^{*}(x_{i})+\epsilon_{i},\quad\epsilon_{i}\text{ are i.i.d. with }\mathbb{E}[\epsilon_{i}]=0,\ \mathrm{Var}(\epsilon_{i})=\sigma_{0}^{2},\quad i=1,\dots,n,$ (21)
where $x_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mu$, and $f^{*}\in L^{2}(\mathcal{X},\mu)$ is the target function. We use a kernel $\mathbf{k}(\cdot,\cdot)$ with Mercer decomposition $\mathbf{k}(x,x^{\prime})=\sum_{j=1}^{\infty}\lambda_{j}\psi_{j}(x)\psi_{j}(x^{\prime})$, where $\{\psi_{j}\}_{j=1}^{\infty}$ form an orthonormal basis for $L^{2}(\mathcal{X},\mu)$ and $\bm{\lambda}=\{\lambda_{j}\}_{j\geq 1}$ is the sequence of eigenvalues not necessarily sorted. For simplicity, we assume there are no ties among the eigenvalues and let $\pi$ be the permutation that sorts them in descending order, so that $\lambda_{\pi_{1}}>\lambda_{\pi_{2}}>\dots>\lambda_{\pi_{k}}>\ldots$. The coefficients of $f^{*}$ in this basis are $\theta_{j}^{*}=\langle f^{*},\psi_{j}\rangle_{L^{2}(\mu)}$. An estimator $\hat{f}$ has risk $\mathcal{R}(\hat{f};f^{*})=\mathbb{E}\|\hat{f}-f^{*}\|_{L^{2}(\mu)}^{2}$. If we define $\hat{\theta}_{j}=\langle\hat{f},\psi_{j}\rangle_{L^{2}(\mu)}$, we can also write the risk as
$\mathcal{R}(\hat{f};f^{*})=\mathbb{E}\|\hat{f}-f^{*}\|_{L^{2}(\mu)}^{2}=\mathbb{E}\bigl[\sum_{j=1}^{\infty}\left(\hat{\theta}_{j}-\theta_{j}^{*}\right)^{2}\bigr]=\sum_{j=1}^{\infty}\mathbb{E}\left(\hat{\theta}_{j}-\theta_{j}^{*}\right)^{2}.$ (22)
The expression in Equation˜22 suggests that we can equivalently estimate each $\theta_{j}^{*}$ separately using the transformed observation $z_{j}=n^{-1}\sum_{i}y_{i}\psi_{j}(x_{i})$ for any $j\geq 1$ as introduced in Equation˜2.
Here we kindly remind the reader that in RKHS, we use subscript $i$ for samples and subscript $j$ for eigen-coordinates. The subscript $j$ aligns with the use of notation in the sequence model, where we use indices $j$ to denote coordinates.
Inflated variance in Equation˜3.
Before we introduce the definition of ESD, we demonstrate that the approximation in Equation˜3 can be made exact by increasing the variance to absorb the approximation error.
For each $j\geq 1$, we can write the transformed observation as
$z_{j}$ $=\frac{1}{n}\sum_{i=1}^{n}y_{i}\,\psi_{j}(x_{i})$
$=\frac{1}{n}\sum_{i=1}^{n}\big(f^{*}(x_{i})+\epsilon_{i}\big)\,\psi_{j}(x_{i})$
$=\frac{1}{n}\sum_{i=1}^{n}\left(\sum_{k\geq 1}\theta_{k}^{*}\psi_{k}(x_{i})\right)\psi_{j}(x_{i})\;+\;\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\,\psi_{j}(x_{i})$
$=\sum_{k\geq 1}\theta_{k}^{*}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}\psi_{k}(x_{i})\psi_{j}(x_{i})\right)}_{:=G_{kj}}\;+\;\underbrace{\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\,\psi_{j}(x_{i})}_{=\xi_{j}}$
$=\sum_{k\geq 1}G_{kj}\,\theta_{k}^{*}\;+\;\xi_{j},$
where $G_{kj}:=\frac{1}{n}\sum_{i=1}^{n}\psi_{k}(x_{i})\psi_{j}(x_{i})$ are entries of the empirical Gram matrix. For $\xi_{j}$, we have $\mathbb{E}(\xi_{j}\mid\{x_{i}\}_{i\in[n]})=0$ and $\mathbb{E}(\xi_{j}\xi_{k}\mid\{x_{i}\}_{i\in[n]})=n^{-1}\sigma_{0}^{2}G_{jk}$.
Since $x_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mu$ and $\{\psi_{j}\}$ are orthonormal in $L^{2}(\mathcal{X},\mu)$, we have $\mathbb{E}[G_{kj}]=\mathbf{1}_{\{k=j\}}$. Hence $\mathbb{E}[z_{j}]=\theta_{j}^{*}$.
We may further decompose $z_{j}$ as follows
$z_{j}-\theta_{j}^{*}=(G_{jj}-1)\theta_{j}^{*}+\left(\sum_{k\geq 1,k\neq j}G_{kj}\theta_{k}^{*}\right)+\xi_{j}=\Delta_{j}+\xi_{j},$
where we have defined $\Delta_{j}=(G_{jj}-1)\theta_{j}^{*}+\left(\sum_{k\geq 1,k\neq j}G_{kj}\theta_{k}^{*}\right)$. This term does not appear in the sequence model; its randomness arises solely from the random covariates $x_{i}$. As $n\to\infty$, this term vanishes because $G_{jj}=n^{-1}\sum_{i}\psi_{j}\left(x_{i}\right)^{2}\to 1$ and $G_{kj}=n^{-1}\sum_{i=1}^{n}\psi_{j}\left(x_{i}\right)\psi_{j^{\prime}}\left(x_{i}\right)\to 0$ for $k\neq j$. Furthermore, since $\mathbb{E}(\xi_{j}\mid\{x_{i}\}_{i\in[n]})=0$, we have
$\mathbb{E}(\Delta_{j}\xi_{j})=\mathbb{E}(\Delta_{j}\mathbb{E}(\xi_{j}\mid\{x_{i}\}_{i\in[n]}))=0.$
The presence of $\Delta_{j}$ effectively inflates the variance in $z_{j}$ to $\sigma_{0}^{2}/n+\mathrm{Var}(\Delta_{j})$. One can show that $\mathrm{Var}(\Delta_{j})=\mathrm{Var}(f^{*}(x)\psi_{j}(x))$, which is bounded by $\|f^{*}\|_{\infty}^{2}$. This is how we will control the impact of $\Delta_{j}$ in the following development.
9.2 ESD for RKHS Regression
We start by analyzing the counterpart of the PC estimator, the Kernel Principal Component Projection Estimator (KPCPE), defined as
$\hat{f}_{k}^{\operatorname{PC}}(x):=\sum_{j:\lambda_{j}\geq\lambda_{\pi_{k}}}z_{j}\psi_{j}(x),$ (23)
where $k$ is the number of leading eigenfunctions to be included.
The risk $\mathcal{R}_{k}:=\mathbb{E}\|\hat{f}_{k}^{\operatorname{PC}}-f^{*}\|_{L^{2}(\mu)}^{2}$ decomposes into squared bias $B(k)$ and variance $V(k)$. Since $\mathbb{E}[\psi_{j^{\prime}}(x_{i})\psi_{j}(x_{i})]=\mathbf{1}_{\{j=j^{\prime}\}}$, we have $\mathbb{E}[z_{j}]=\theta_{j}^{*}$. Therefore, the bias is due to truncation as
$B(k)=\sum_{j=k+1}^{\infty}(\theta_{j}^{*})^{2}.$ (24)
The integrated variance is $V(k)=\sum_{j:\lambda_{j}\geq\lambda_{\pi_{k}}}\mathrm{Var}(z_{j})$. Using the law of total variance, we have
$\mathrm{Var}(z_{j})=\frac{1}{n}\mathrm{Var}(y_{i}\psi_{j}(x_{i}))=\frac{1}{n}(\sigma_{0}^{2}+\tau_{j}^{2}),\quad\text{where}\quad\tau_{j}^{2}:=\mathrm{Var}(f^{*}(x)\psi_{j}(x)).$ (25)
The term $\tau_{j}^{2}$ arises from the randomness of the design $x$. To ensure $V(k)$ grows at the rate of $k/n$, we need to uniformly bound the design-induced variance $\tau_{j}^{2}$. To illustrate the idea, we assume $f^{*}$ is bounded in the sense that $|f^{*}(X)|\leq\|f^{*}\|_{\infty}$, $\mu$-almost surely. Here, $\|f^{*}\|_{\infty}$ denotes the essential supremum of $|f^{*}|$ w.r.t. the measure $\mu$.
Assumption 9.1 (Bounded target).
$f^{*}\in L^{\infty}(\mathcal{X},\mu)$ and $\|f^{*}\|_{\infty}=\operatorname{ess}\sup|f^{*}|$.
˜9.1 is very mild: for compact $\mathcal{X}$, if $f$ is continuous, then $f$ is bounded.
Under Assumption 9.1, $\tau_{j}^{2}\leq\mathbb{E}[f^{*}(x)^{2}\psi_{j}(x)^{2}]\leq\|f^{*}\|_{\infty}^{2}$. Subsequently, the variance is bounded by $V(k)\leq\frac{k}{n}(\sigma_{0}^{2}+\|f^{*}\|_{\infty}^{2})$. This motivates us to define the effective noise variance per component as
$\sigma^{2}:=\frac{\sigma_{0}^{2}+\|f^{*}\|_{\infty}^{2}}{n}.$ (26)
The effective noise variance $\sigma^{2}$ includes the term $\|f^{*}\|_{\infty}^{2}/n$, which inflates the noise compared to an idealized sequence model.
We can now adapt the definitions from Section˜3.1 using the effective noise variance $\sigma^{2}$.
Definition 9.2 (ESD for RKHS Regression).
The Effective Span Dimension (ESD) of $f^{*}$ with respect to the kernel $\mathbf{k}$ and the effective noise variance $\sigma^{2}=(\sigma_{0}^{2}+\|f^{*}\|_{\infty}^{2})/n$ is defined as
$d^{\dagger}=d^{\dagger}(\sigma^{2};f^{*},\mathbf{k})=\min\{k\in\mathbb{N}_{+}\cup\{\infty\}:\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\leq\sigma^{2}\},$ (27)
where $\bm{\theta}^{*}=\{\theta_{j}^{*}\}_{j\geq 1}$ and $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)$ is defined as in Equation˜9.
The risk of the KPCPE estimator can be bounded using this ESD.
Proposition 9.3 (Optimal KPCPE Risk Bound).
Let $\hat{f}_{k}^{\operatorname{PC}}$ be the KPCPE estimator defined in Equation˜23. Let $\mathcal{R}_{*}^{\operatorname{PC}}=\min_{k\geq 1}\mathcal{R}(\hat{f}_{k}^{\operatorname{PC}};f^{*})$. Under Assumption 9.1, it holds that:
$(d^{\dagger}-1)\frac{\sigma_{0}^{2}}{n}\leq\mathcal{R}_{*}^{\operatorname{PC}}\leq 2d^{\dagger}\sigma^{2}=2d^{\dagger}\frac{\sigma_{0}^{2}+\|f^{*}\|_{\infty}^{2}}{n},$ (28)
where $d^{\dagger}=d^{\dagger}(\sigma^{2};f^{*},\mathbf{k})$ is the ESD from Definition 9.2. In particular, if $\|f^{*}\|_{\infty}^{2}\lesssim\sigma_{0}^{2}$, we can conclude that $\mathcal{R}_{*}^{\operatorname{PC}}\asymp d^{\dagger}\sigma_{0}^{2}/n$.
We comment that Zhang et al. [51] have established the minimax optimality of the well-tuned PC estimator. Proposition˜9.3 suggests that the risk of the well-tuned PC estimator scales as $d^{\dagger}/n$. Therefore, we essentially express the minimax rate therein using the ESD without reliance on the classical eigen-decay conditions or source conditions.
We can also extend the minimax framework in Section˜4 to RKHS regression. Let $K$ be a quota on the ESD, and let $C_{0}$ be a constant. Define the class based on the span profile as follows:
$\mathcal{F}_{K,\mathbf{k}}^{(n)}=\{f^{*}\in L^{2}(\mathcal{X},\mu)\cap L^{\infty}(\mathcal{X},\mu):\|f^{*}\|_{\infty}\leq\sigma_{0}C_{0},\quad d^{\dagger}(\bar{\sigma}^{2}/n;f^{*},\mathbf{k})\leq K\},$ (29)
where $\bar{\sigma}^{2}=\sigma_{0}^{2}(1+C_{0}^{2})$. We further impose the following assumption on the spectrum.
Assumption 9.4.
The kernel $\mathbf{k}$ is said to be $(K,n)$-regular if there are some constants $c_{1}\in(0,1)$ and $C_{1}$ such that $\sum_{i\leq c_{1}K}\lambda_{\pi_{i}}^{-1}\leq C_{1}n$.
Theorem 9.5 (Minimax Risk over Span Profile Classes).
If $\mathbf{k}$ is $(K,n)$-regular, then the minimax risk over $\mathcal{F}_{K,\mathbf{k}}^{(n)}$ satisfies:
$\inf_{\hat{f}}\sup_{f^{*}\in\mathcal{F}_{K,\mathbf{k}}^{(n)}}\mathcal{R}(\hat{f};f^{*})\asymp\frac{\sigma_{0}^{2}K}{n},$ (30)
where the infimum is over all estimators $\hat{f}$.
Combining Theorem˜9.5 and Proposition˜9.3, the optimally tuned KPCPE estimator is minimax rate optimal over $\mathcal{F}_{K,\mathbf{k}}^{(n)}$, with rate $\sigma_{0}^{2}K/n$.
The KPCPE serves as a simple benchmark for spectral methods. This analysis, via the ESD, characterizes the performance of the optimally tuned KPCPE based directly on the properties of the specific signal $f^{*}$ (via $\theta^{*}$) and kernel spectrum $\bm{\lambda}$, without requiring standard assumptions like source conditions or polynomial eigenvalue decay. Therefore, we consider the ESD evaluated at the design-adjusted noise level $\sigma^{2}$ as a key measure of statistical complexity in RKHS regression. In summary, the span profile framework provides a unified perspective on the generalization performance of spectral methods across a variety of models.
Minimax convergence rates.
Following the framework in Section˜4, we can quantify a class of populations using a quota sequence $\bm{K}=\{K_{n}\}_{n=1}^{\infty}$. For some $n_{0}\in\mathbb{N}_{+}$, define
$\mathcal{F}_{\bm{K},\mathbf{k}}=\{f^{*}\in L^{2}(\mathcal{X},\mu)\cap L^{\infty}(\mathcal{X},\mu):\|f^{*}\|_{\infty}\leq\sigma_{0}C_{0},\quad d^{\dagger}(\bar{\sigma}^{2}/n;f^{*},\mathbf{k})\leq K_{n},\forall n\geq n_{0}\},$ (31)
where $\bar{\sigma}^{2}=\sigma_{0}^{2}(1+C_{0}^{2})$. For a sample $\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from the model in Equation˜21 and any estimator $\hat{f}$, we aim to determine the optimal convergence rate of the following minimax risk:
$\inf_{\hat{f}}\sup_{f^{*}\in\mathcal{F}_{\bm{K},\mathbf{k}}}\mathcal{R}(\hat{f},f^{*}).$ (32)
We have the following result.
Theorem 9.6.
Suppose ˜4.1 holds for a quota sequence $\bm{K}=\{K_{n}\}_{n=1}^{\infty}$. Furthermore, suppose $\mathbf{k}$ is $(K_{n},n)$-regular for all $n\geq n_{0}$. If $\{(x_{i},y_{i})\}_{i=1}^{n}$ is drawn from the model in Equation˜21, it holds that
$\inf_{\hat{f}}\sup_{f^{*}\in\mathcal{F}_{\bm{K},\mathbf{k}}}\mathcal{R}(\hat{f},f^{*})\asymp\bar{\sigma}^{2}\frac{K_{n}}{n}.$
˜9.4 is a mild condition. The following is an example where we use Theorem˜9.6 to recover the minimax convergence rate derived under the classical polynomial eigen-decay conditions and source conditions.
Example 9.7.
Suppose $\mathbf{k}$ admits spectrum such that $\lambda_{\pi_{i}}\asymp i^{-\beta}$ with $\beta>0$. Let $K_{n}=\lfloor n^{\frac{1}{1+s\beta}}\rfloor$ for any $s\geq 1$. It is easy to see that
$\sum_{i\leq c_{1}K_{n}}\lambda_{\pi_{i}}^{-1}\asymp(c_{1}K_{n})^{\beta+1}\asymp n^{\frac{\beta+1}{s\beta+1}}\lesssim n.$
Therefore, the minimax optimal rate for the class $\mathcal{F}_{\bm{K},\mathbf{k}}$ is $\frac{\bar{\sigma}^{2}K_{n}}{n}\asymp\sigma_{0}^{2}n^{-\frac{s\beta}{1+s\beta}}$, which is the same as the optimal rate given by the source condition with smoothness parameter $s$.
Remark 9.8.
In Section˜2, we simplify our discussion by assuming that $\mathbf{k}$ is positive definite. In practice, positive semi-definite (PSD) kernels may also be used. In the case where $\mathbf{k}$ has rank $d<\infty$, spectral algorithms inevitably induce a systematic (squared) bias
$\Delta_{f^{*},\mathbf{k}}=\|f^{*}\|_{L^{2}(\mu)}^{2}-\sum_{j\in[d]}\left(\theta_{\pi_{i}}^{*}\right)^{2},$
regardless of the regularization parameter. In that case, we may modify the definition of ESD by adding the systematic bias into the summation to the sum of squared tails.
Specifically, for a PSD kernel $\mathbf{k}$ of rank $d<\infty$, we modify the definition of ESD in Definition˜9.2 as
$\bar{d}^{\dagger}=\bar{d}^{\dagger}(\sigma^{2};f^{*},\mathbf{k})=\min\left\{j\in[d]:\frac{1}{j}\left[\Delta_{f^{*},\mathbf{k}}+\sum_{i=j+1}^{d}\left(\theta_{\pi_{i}}^{*}\right)^{2}\right]\leq\,\sigma^{2}\right\}.$
Again, $\sigma^{2}\bar{d}^{\dagger}$ characterizes the risk of the well-tuned PC estimator as in Proposition˜9.3.
9.3 Connection to Random Design Linear Regression
The analysis in this section also covers an important case: random-design linear regression (as opposed to fixed-design linear regression). This is because linear regression can be viewed as an RKHS regression w.r.t. any positive-definite linear kernels $\mathbf{k}(x,x^{\prime})=x^{\top}\mathbf{K}x^{\prime}$ for $x,x^{\prime}\in\mathbb{R}^{p}$, where $\mathbf{K}\in\mathbb{R}^{p\times p}$ is positive definite.
Let the support $\mathcal{X}$ be a compact subset of $\mathbb{R}^{p}$. Suppose $\bm{\Sigma}_{x}=\mathbb{E}_{x\sim\mu}(xx^{\top})$ is positive definite and $\mathbf{L}$ is a symmetric square root of $\bm{\Sigma}_{x}$. Let the eigen-decomposition of $\mathbf{L}\mathbf{K}\mathbf{L}$ be
$\mathbf{L}\mathbf{K}\mathbf{L}=\sum_{j=1}^{p}\lambda_{j}v_{j}v_{j}^{\top}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\top},$
where $\mathbf{V}$ is the matrix with columns formed by $v_{j}$ and $\mathbf{\Lambda}=\operatorname{Diag}(\lambda_{1},\dots,\lambda_{p})$.
Define $\mathbf{\Psi}=\mathbf{L}^{-1}\mathbf{V}$. Then $\mathbf{\Psi}^{\top}\bm{\Sigma}_{x}\mathbf{\Psi}=\mathbf{V}^{\top}\mathbf{V}=\mathbf{I}_{p}$. Furthermore, we have
$\mathbf{K}=\mathbf{L}^{-1}\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\top}\mathbf{L}^{-1}=\mathbf{\Psi}\mathbf{\Lambda}\mathbf{\Psi}^{\top},$
Suppose the columns of $\mathbf{\Psi}$ are $\psi_{j}$. We can then write
$\mathbb{E}_{x\sim\mu}\left[\langle\psi_{j},x\rangle\langle\psi_{j},x\rangle\right]=\left(\mathbf{\Psi}^{\top}\bm{\Sigma}_{x}\mathbf{\Psi}\right)_{jk}=\mathbf{1}_{\{j=k\}},~~~\forall j,k\in[p],$
so $\{\psi_{j}\}$ is an orthonormal system in $L^{2}(\mathcal{X},\mu)$. Furthermore, the kernel can be expressed as
$\mathbf{k}(x,x^{\prime})=x^{\top}\mathbf{K}x^{\prime}=\sum_{j=1}^{p}\lambda_{j}\langle\psi_{j},x\rangle\langle\psi_{j},x^{\prime}\rangle.$
Hence, $(\{\lambda_{i}\},\{\psi_{j}\})$ is the eigenpair for $\mathbf{k}$.
For linear regression where $y=f^{*}(x)+\epsilon$ and $f^{*}(x)=\langle\beta^{*},x\rangle$. Define $\theta^{*}=\mathbf{\Psi}^{-1}\beta^{*}=\mathbf{V}^{\top}\mathbf{L}\beta^{*}$. We can write
$f^{*}(x)=\langle\beta^{*},x\rangle=\langle\mathbf{\Psi}^{-1}\beta^{*},\mathbf{\Psi}^{\top}x\rangle=\sum_{j=1}^{p}\theta_{j}^{*}\langle\psi_{j},x\rangle.$
It is also clear that $\|f^{*}\|_{\infty}\leq\sup_{x\in\mathcal{X}}\langle\beta^{*},x\rangle\leq\|\beta^{*}\|_{2}C_{\mathcal{X}}<\infty$, where $C_{\mathcal{X}}$ is finite and depends on $\mathcal{X}$. We can define the effective noise level $\sigma^{2}=n^{-1}(\sigma_{0}^{2}+\|f^{*}\|_{\infty}^{2})$.
Therefore, with respect to the kernel $\mathbf{k}$ and the basis $\{\psi_{j}\}$, we define the ESD exactly as in Definition˜9.2 using the coefficients $\{\theta^{*}_{j}\}$ and eigenvalues $\{\lambda_{j}\}$.
9.4 Numerical Illustration
This section provides numerical validation of the relationship between the ESD and the optimally tuned KPCPE risk, mirroring the setup for linear models in Section˜8.4. We use the cosine basis eigenfunctions $\psi_{j}(x)=\sqrt{2}\cos(2\pi jx)$ on the domain $[0,1]$ with inputs sampled as $x_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Unif}[0,1]$. The sample size is fixed at $n=400$, and for numerical purpose, we consider the first $J=800$ eigenfunctions. The noise variance is set as $\sigma_{0}^{2}=1$.
Experimental Setup: We set the baseline kernel eigenvalue spectrum as $\lambda_{j,0}=j^{-1.1}$ and the fixed signal coefficients as $\theta_{j}^{*}=j^{-4}$. To study the impact of misalignment between the kernel spectrum and the signal, we introduce a severity parameter $\alpha\geq 0$ and define the modified eigenvalue spectrum as
$\lambda_{j}(\alpha)=\lambda_{j,0}\exp\left(\alpha t_{j}\right),\quad t_{j}=\frac{j-1}{D-1}\text{ for }j\leq D,\text{ and }t_{j}=0\text{ otherwise},$
with $D=80$. As $\alpha$ increases, the leading $D$ eigenvalues become progressively magnified, with the largest index having the most significant increase. Consequently, the modified kernel places more emphasis on directions that receive less signal energy, so the optimal KPCPE selects more principal components.
As the severity parameter $\alpha$ grows, only the first $D$ eigenvalues are changed while the rest of the spectrum is untouched. Among the changed ones, the leading eigenvalues are magnified by a smaller constant, so that the resulting kernel has its leading subspaces being on the directions in which the signal has less of its energy and thus increases the misalignment.
For each $\alpha$ in a specified grid, we compute two quantities:
  • The Effective Span Dimension
    $d_{\mathrm{eff}}^{\dagger}(\alpha)=d^{\dagger}\left(\sigma_{\mathrm{eff}}^{2};f^{*},\lambda(\alpha)\right),\quad\sigma_{\mathrm{eff}}^{2}:=\frac{\sigma_{0}^{2}+\sigma_{f,4}^{2}}{n},$
    where the design-induced variance $\sigma_{f,4}^{2}=\max_{j\leq J}\operatorname{Var}\left\{f^{*}(X)\psi_{j}(X)\right\}$ is computed numerically based on the fixed sample of $x_{i}$ and the true $f^{*}$.
  • The optimally tuned KPCPE risk
    $\mathcal{R}_{*}(\alpha)=\min_{k}\mathbb{E}\|\hat{f}_{k}^{\operatorname{PC}}(\alpha)-f^{*}\|_{L^{2}(\mu)}^{2},$
    where the estimator $\hat{f}_{k}^{\operatorname{PC}}(\alpha)$ is computed using the spectrum $\lambda(\alpha)$ and the expectation is estimated by averaging prediction error over $B=10$ Monte Carlo replications.
Figure 4: Effective Span Dimension and Optimal KPCPE risk. The dashed line plots Risk $\times n/\sigma_{0}^{2}$; the solid line is $d^{\dagger}(\alpha)$. The risk is computed based on 20 replications and the error bar represents the standard deviation.
Figure 4 plots the empirically computed optimal KPCPE risk (orange solid line) alongside the theoretical lower bound $(d^{\dagger}-1)\sigma_{0}^{2}/n$ (blue dashed line) and upper bound $2d^{\dagger}\sigma_{\mathrm{eff}}^{2}$ (green dotted line). The empirical risk consistently lies between the two theoretical curves, confirming the validity of the bounds derived in Proposition 9.3.
As the severity parameter $\alpha$ increases, the resulting spectral perturbation shifts energy into higher-index eigenfunctions. This inflates the ESD and consequently the minimal achievable risk.
Overall, this experiment demonstrates that our span profile framework provides an accurate and robust characterization of generalization performance in RKHS regression, consistent with earlier observations made for the sequence and linear regression models.
10 Measuring Alignment via ESD
We illustrate how the notion of ESD can be used to measure the alignment between the signal and the kernel.
10.1 An example comparing signal-spectrum alignment
The following simple example illustrates how to compare signal-spectrum alignment across the spectra discussed in Section˜3.2.
Suppose $\bm{\theta}^{*}$ is $s$-sparse with support $S\subset[d]$ and $s=|S|\ll d$. Consider the following two spectra with the same set of eigenvalues but different allocations:
(1) The $k$ largest eigenvalues of $\bm{\lambda}^{(1)}$ are located on $S$;
(2) The $k$ largest eigenvalues of $\bm{\lambda}^{(2)}$ are located on $S^{c}=[d]\setminus S$.
Intuitively, $\bm{\lambda}^{(1)}$ aligns better with $\bm{\theta}^{*}$ than $\bm{\lambda}^{(2)}$. However, a quantitative analysis is not straightforward without using the notion of ESD.
First, we note that the effective dimensions [52] is the same for both spectra because they share the same sets of eigenvalues. Similarly, the covariance-splitting index $k^{*}$ [6] is the same for both spectra. Thus, these signal-agnostic complexity measures do not distinguish signal-spectrum alignment between the two spectra.
Next, we consider the ESD and the span profile. Rigorously, we can show for any $\tau$,
$\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(1)}}(\tau)\leq s,\text{ and }\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(2)}}(\tau)\geq\min\left(d-s,\|\bm{\theta}^{*}\|^{2}/\tau\right).$
Hence, for sufficiently small $\tau$, their ratio
$r(\tau)={\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(1)}}(\tau)}/{\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}^{(2)}}(\tau)}\leq s/(d-s)\ll 1.$
In view of Theorem˜3.3, this suggests that the minimax risk under $\bm{\lambda}^{(1)}$ is substantially lower than under $\bm{\lambda}^{(2)}$ when the noise level is small. Therefore, spectral estimators using $\bm{\lambda}^{(1)}$ are preferred.
10.2 Pathwise ESD for Learned Kernels
Section˜6 analyzes eigenvalue learning because OP-GF admits tractable dynamics under a fixed eigenbasis. This is a limitation of that specific analysis, not of the ESD concept. In fact, ESD applies to general representation learning. We illustrate how decreases in ESD explain minimax risk reduction for learned kernels, whether adaptation acts through eigenvalues, eigenfunctions, or both.
Let $\mathbf{k}_{t}$ be the kernel learned at training time $t$, with eigenvalues $\{\lambda_{j}(t)\}$ (sorted in decreasing order) and eigenfunctions $\{\psi_{j}^{(t)}\}$ that are orthonormal in $L^{2}(\mathcal{X},\mu)$. To understand how the signal-kernel alignment evolves, we define the pathwise ESD as
$d^{\dagger}(t):=d^{\dagger}(\sigma^{2};f^{*},\mathbf{k}_{t}),t\geq 0,$
where we have followed Definition˜9.2 to define $d^{\dagger}(\sigma^{2};f^{*},\mathbf{k}_{t})$ as the ESD of $f^{*}$ w.r.t. the kernel $\mathbf{k}_{t}$ using $\theta_{j}^{*,(t)}=\langle f^{*},\psi_{j}^{(t)}\rangle$ and $\sigma^{2}=n^{-1}(\sigma_{0}^{2}+\|f^{*}\|_{\infty}^{2})$.
Let $\mathbf{H}_{t}(k):=\frac{1}{k}\sum_{i>k}\left[\theta^{*,(t)}_{i}\right]^{2}$. If training aligns the leading eigenfunctions $\psi_{j}^{(t)}$ better with $f^{*}$, then $\{\theta_{j}^{*,(t)}\}$ concentrate more on the leading indices, and thus $\mathbf{H}_{t}(k)$ decreases for all $k$, which implies a decrease in $d^{\dagger}(t)$.
Experiment on Deep Linear Networks.
To demonstrate this pathwise perspective, we simulate a random-design linear regression.
Each covariate coordinate is drawn independently from $\{\pm 1\}$, so $\bm{\Sigma}_{x}=\mathbb{E}(XX^{\top})=\mathbf{I}_{p}$ and $\|X\|_{\infty}=1$. We set $p=900$, and specify the true parameters as follows: $\beta^{*}$ follows a power-law decay with $\beta^{*}_{j}=j^{-1.1}$ for $1\leq j\leq 200$ and $\beta^{*}_{j}=0$ for $j>200$. The response is $Y=\langle\beta^{*},X\rangle+\varepsilon$ with $\varepsilon\sim N(0,\sigma_{0}^{2})$ and $\sigma_{0}=0.1$.
We draw $n=1000$ samples and train a deep linear network with $D=4$ hidden affine layers without bias using full-batch Adam with learning rate $10^{-4}$. The hidden weight matrices of the network are $\mathbf{W}_{\ell}(t)\in\mathbb{R}^{p\times p}$ for $\ell=1,\ldots,D$ (using a near-identity initialization), and the weight of the final linear layer is $w(t)\in\mathbb{R}^{p}$.
The estimated function at time $t$ is given by $f_{t}(x)=w(t)^{\top}\mathbf{A}(t)x$, where $\mathbf{A}(t):=\mathbf{W}_{D}(t)\cdots\mathbf{W}_{1}(t)$. We form the learned kernel $\mathbf{k}_{t}(x,x^{\prime})=\langle\mathbf{A}(t)x,\mathbf{A}(t)x^{\prime}\rangle=x^{\top}\mathbf{G}_{t}x^{\prime}$, where $G_{t}:=\mathbf{A}(t)^{\top}\mathbf{A}(t)$. We then follow the derivation in Section˜9.3 and define the ESD $d^{\dagger}(t)$ of $f^{*}$ w.r.t. the kernel $\mathbf{k}_{t}$. Since $\|X\|_{\infty}=1$ $\mu$-a.s., we have $\|f^{*}\|_{\infty}^{2}=\|\beta^{*}\|_{1}^{2}$; this is used in computing the effective noise level.
Figure 5: Pathwise ESD and risk under a learned kernel using a 4-layer linear network.
Figure˜5 shows that adaptive representation learning progressively reduces the ESD $d^{\dagger}(t)$ over training time, along with the true risk. This confirms that ESD captures the evolving alignment between signal and kernel.
11 Discussion
This paper introduces the effective span dimension (ESD) and span profile to analyze the interplay between the signal structure and the kernel spectrum. Our framework moves beyond classical static assumptions relative to a fixed kernel (e.g., source conditions and polynomial eigenvalue decay) and offers a dynamic, noise-dependent perspective on signal complexity. Unlike traditional source conditions, the ESD is more flexible and remains applicable when the spectrum itself is learned from data.
Like the sparsity level in high dimensional statistics, ESD is a population quantity for theory rather than a training input. It is useful as a quantitative diagnostic: comparing ESD across kernels clarifies which representation supports better generalization for a given signal.
In summary, the ESD framework provides a novel view of generalization that connects classical kernel methods with modern adaptive learning. We expect to relate this framework to learned representations in neural networks to explain their superior generalization performance.
Appendix A Proof
A.1 Proofs of Results on ESD of Sequence Models
Proof of Theorem˜3.2.
For any $\nu>0$, define
$k_{\Lambda}(\nu)\;=\;\#\{\,j:\lambda_{j}\geq\nu\},$
which counts how many eigenvalues exceed the threshold $\nu$. The KPCR estimator sets
$\hat{\theta}_{i}^{\nu}\;=\;\mathbf{1}_{\{\lambda_{i}\geq\nu\}}\;z_{i},\;i\in[d].$
Its squared bias and variance are given by
$B^{\operatorname{PC}}(v)\;=\;\sum_{i:\lambda_{i}<v}\left(\theta_{i}^{*}\right)^{2},\;V^{\operatorname{PC}}(v)\;=\;\sum_{i:\lambda_{i}\geq v}\sigma^{2}\;=\;k_{\Lambda}(v)\,\sigma^{2}.$
For any threshold $v$, we can reparameterize the bias and variance using $k\;=\;k_{\Lambda}(v)$ as
$B^{\operatorname{PC}}(k)\;=\;\sum_{i=k+1}^{d}\left(\theta_{\pi_{i}}^{*}\right)^{2},\quad\text{and}\quad V^{\operatorname{PC}}(k)\;=\;k\,\sigma^{2},k=0,1,\ldots,d.$
The function $B^{\operatorname{PC}}(k)$ decreases in $k$, while $V^{\operatorname{PC}}(k)$ increases in $k$. The risk function is given by $\mathcal{R}^{\operatorname{PC}}(k)=B^{\operatorname{PC}}(k)+V^{\operatorname{PC}}(k)$.
For any integer $k\geq 1$, we have
$\mathcal{R}^{\operatorname{PC}}(k)\;=\;k\;\Bigl(\sigma^{2}\;+\;\frac{1}{k}\sum_{i:\lambda_{i}<\lambda_{\pi_{k}}}\left(\theta_{i}^{*}\right)^{2}\Bigr).$
Upper bound
For $k=d^{\dagger}$, we have $\frac{1}{k}\sum_{i:\lambda_{i}<\lambda_{\pi_{k}}}\left(\theta_{i}^{*}\right)^{2}\leq\sigma^{2}$. By definition of the optimal risk, we have
$\mathcal{R}_{*}^{\operatorname{PC}}\;\leq\;\mathcal{R}^{\operatorname{PC}}(d^{\dagger})\;\leq\;2\,d^{\dagger}\,\sigma^{2}.$
Therefore, the upper bound is proved.
Lower bound
Without loss of generality, assume $d^{\dagger}\geq 2$. For any $k\leq d^{\dagger}-1$, we have
$\mathcal{R}^{\operatorname{PC}}(k)\geq B^{\operatorname{PC}}(k)=\sum_{i=k+1}^{d}\left(\theta_{\pi_{i}}^{*}\right)^{2}\geq\sum_{i=d^{\dagger}}^{d}\left(\theta_{\pi_{i}}^{*}\right)^{2}>(d^{\dagger}-1)\sigma^{2},$
where the last inequality comes from the definition of $d^{\dagger}$. For any $k\in[d^{\dagger},d]$, we have
$\mathcal{R}^{\operatorname{PC}}(k)\geq k\sigma^{2}\geq d^{\dagger}\sigma^{2}.$
Therefore, the lower bound is proved. ∎
A.2 Proof on Minimax Results
Proof of Theorem˜4.3.
Throughout the proof, the quota sequence is fixed. Recall the definition of $M_{k}$ in ˜4.1. Define $\psi(k)=\sigma_{0}^{2}\frac{k}{M_{k}}$ for any $k\in\bar{K}$. Also define $\mathbf{S}_{\bm{\theta},\bm{\lambda}}(k)=k\mathbf{H}_{\bm{\theta},\bm{\lambda}}(k)$.
We can express $\mathcal{F}_{\bm{K},\bm{\lambda}}$ as follows.
Lemma A.1.
Under ˜4.1, we have
$\mathcal{F}_{\bm{K},\bm{\lambda}}\;=\;\Bigl\{\bm{\theta}\in\mathbb{R}^{\infty}:\mathbf{S}_{\bm{\theta},\bm{\lambda}}(k)\leq\psi(k)\text{ for all }k\in[\bar{K}]\Bigr\}.$
Proof of Lemma˜A.1.
Observe the relation that
$\bm{\theta}\in\mathcal{F}_{\bm{K},\bm{\lambda}}\quad\Longleftrightarrow\quad\mathbf{D}_{\bm{\theta},\bm{\lambda}}\!\Bigl(\tfrac{\sigma_{0}^{2}}{n}\Bigr)\;\leq\;K_{n},~~\forall n\geq 1\quad\Longleftrightarrow\quad\mathbf{S}_{\bm{\theta},\bm{\lambda}}(K_{n})\leq\sigma_{0}^{2}\frac{K_{n}}{n},~~\forall n\geq 1.$
By (1) of ˜4.1, we have
$\mathbf{S}_{\bm{\theta},\bm{\lambda}}(K_{n})\leq\sigma_{0}^{2}\frac{K_{n}}{n},~~\forall n\geq 1\quad\Longleftrightarrow\quad\mathbf{S}_{\bm{\theta},\bm{\lambda}}(k)\leq\psi(k)\text{ for all }k\in[\bar{K}].$
Therefore, we can rewrite
$\mathcal{F}_{\bm{K},\bm{\lambda}}\;=\;\Bigl\{\bm{\theta}\in\mathbb{R}^{\infty}:\mathbf{S}_{\bm{\theta},\bm{\lambda}}(k)\leq\psi(k)\text{ for all }k\in[\bar{K}]\Bigr\}.$
Fix any $n$. Define $\delta=\sqrt{c\sigma_{0}^{2}/n}$ with the constant $c=\tfrac{1}{4}\wedge\tau$, where $\tau$ comes from ˜4.1. Consider assigning nonzero signals on the block $B_{n}=\{\pi_{1},\dots,\pi_{K_{n}}\}$ to construct a subset of populations.
Specifically, we define the collection of hypercubes vertices $\mathcal{V}=\{-1,1\}^{K_{n}}$. For every vertex $v\in\mathcal{V}$, define a parameter vector $\bm{\theta}^{(v)}=(\theta_{j}^{(v)})_{j=1}^{d}$ as follows:
$\theta_{\pi_{i}}^{(v)}=\delta\,v_{i},\text{ for }i=1,\ldots,K_{n},\qquad\text{ and }\theta_{j}^{(v)}=0\text{ for }j\notin B_{n}.$ (33)
There are $2^{K_{n}}$ such vectors $\{\bm{\theta}^{(v)}\}$, and they satisfy the following property.
Lemma A.2.
For any $v\in\mathcal{V}$, the parameter vector $\bm{\theta}^{(v)}$ constructed in Equation˜33 lies in $\mathcal{F}_{\bm{K},\bm{\lambda}}$.
Proof of Lemma˜A.2.
For any $k\in[\bar{K}]$, if $k\geq K_{n}$, then $\mathbf{S}_{\bm{\theta}^{(v)},\bm{\lambda}}(k)$ is 0.
If $1\leq k\leq K_{n}-1$, then $\mathbf{S}_{\bm{\theta}^{(v)},\bm{\lambda}}(k)\leq K_{n}\delta^{2}=c\sigma_{0}^{2}K_{n}/n$. Denote $k_{0}=K_{n}-1$ and $L=1+M_{k_{0}}$. By definition of $M_{k_{0}}$, we have $n\geq L$. Since $k\leq k_{0}$, we have $L\geq 1+M_{k}$. We have
$\frac{K_{n}}{n}$ $\leq\frac{K_{n}}{L}$ (34)
$=\frac{k_{0}+1}{1+M_{k_{0}}}$
$\leq 2\frac{k_{0}}{M_{k_{0}}}$
$\leq 2\frac{k}{M_{k}},$
where the second last inequality is because $(1+k_{0})/(1+m)\leq 2k_{0}/m\Leftrightarrow m+k_{0}m\leq 2k_{0}+2k_{0}m$ and the last inequality is due to (2) of ˜4.1. Since $2c<1$, we see that $\sigma_{0}^{2}cK_{n}/n\leq\psi(k)$ for all $k<K_{n}$.
In either case, we have $\mathbf{S}_{\bm{\theta}^{(v)},\bm{\lambda}}(k)\leq\psi(k)$ for all $k\in[\bar{K}]$, and thus $\bm{\theta}^{(v)}\in\mathcal{F}_{\bm{K},\bm{\lambda}}$. ∎
For each $v\in\mathcal{V}$, let $P_{v}$ be the sampling distribution of the sequence model in Equation˜4 with $\bm{\theta}^{*}=\bm{\theta}^{(v)}$, $\sigma^{2}=\sigma_{0}^{2}/n$, and $\{\xi_{j}\}_{j\in[d]}$ being i.i.d. from $N(0,\sigma^{2})$. Let $\rho$ be Hamming distance on $\mathcal{V}$. If $v$ and $w\in\mathcal{V}$ differ in exactly one coordinate (i.e., $\rho(v,w)=1$), then
  • $\|\bm{\theta}^{(v)}-\bm{\theta}^{(w)}\|^{2}\geq(2\delta)^{2}$, and
  • the Kullback-Leibler divergence between $P_{v}$ and $P_{w}$ satisfies $\operatorname{KL}\!\bigl(P_{v}\,\|\,P_{w}\bigr)=\frac{1}{2\sigma^{2}}(2\delta)^{2}=2c\leq\tfrac{1}{2}$, and by the Pinsker’s inequality, $\|P_{v}\wedge P_{w}\|=1-\operatorname{TV}(P_{v},P_{w})\geq 1-\sqrt{\operatorname{KL}\!\bigl(P_{v}\,\|\,P_{w}\bigr)/2}\geq 1/2$.
By Assouad’s Lemma (Lemma 2 in Yu [49]), for any estimator $\widehat{\bm{\theta}}$ based on a sample $Y^{(n)}$ drawn from $P_{v}$, we have
$\sup_{v\in\mathcal{V}}\mathbb{E}_{v}\bigl\lVert\widehat{\bm{\theta}}-\bm{\theta}^{(v)}\bigr\rVert^{2}\geq K_{n}\frac{(2\delta)^{2}}{4}=c\sigma_{0}^{2}\frac{K_{n}}{n}.$
Proof of Theorem˜3.3.
The upper bound is given by Theorem˜3.2, so we only need to prove the lower bound. The main idea is the same as the proof for Theorem˜4.3.
Let $\delta=\sqrt{c\sigma_{0}^{2}/n}$ with the constant $c=\tfrac{1}{4}$. Let $B_{n}=\{\pi_{1},\dots,\pi_{K_{n}}\}$. We define the collection of hypercubes vertices $\mathcal{V}=\{-1,1\}^{K}$. For every vertex $v\in\mathcal{V}$, define a parameter vector $\bm{\theta}^{(v)}=(\theta_{j}^{(v)})_{j=1}^{d}$ as in Equation˜33. There are $2^{K}$ such vectors $\{\bm{\theta}^{(v)}\}$. For each $v\in\mathcal{V}$, let $P_{v}$ be the sampling distribution of the sequence model in Equation˜4 with $\bm{\theta}^{*}=\bm{\theta}^{(v)}$, $\sigma^{2}=\sigma_{0}^{2}/n$, and $\{\xi_{j}\}_{j\in[d]}$ being i.i.d. normal. Let $\rho$ be Hamming distance on $\mathcal{V}$. The rest of the proof is identical to that of Theorem˜4.3 and is omitted. ∎
A.3 Details of Example˜3.4
We provide the details of Example˜3.4 for illustration of the concepts of ESD and span profile through several examples.
Example A.3 (Polynomial spectrum with source condition).
Assume $\lambda_{i}=i^{-\beta}$ for some $\beta>0$ and the source condition $\sum_{i=1}^{d}\lambda_{i}^{-s}\,\theta_{i}^{*2}\leq R$ with $s>0$. The trade-off function satisfies
$\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\;=\;\frac{1}{k}\sum_{i>k}\theta_{i}^{*2}\;\leq\;\frac{\lambda_{k}^{s}}{k}\sum_{i>k}\lambda_{i}^{-s}\theta_{i}^{*2}\;\leq\;R\,k^{-(1+s\beta)},$
which follows that $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma^{2})\;\lesssim\;\,[\sigma^{2}]^{-\frac{1}{1+s\beta}}$. Since $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma^{2})\leq d$, the optimal risk of PC estimator satisfies
$\mathcal{R}_{*}^{\operatorname{PC}}\;\lesssim\;\min\left(\,[\sigma^{2}]^{\frac{s\beta}{1+s\beta}},d\sigma^{2}\right).$
In Example˜A.3, we note that for $\sigma^{2}=\sigma_{0}^{2}/n$, the upper bound becomes $\sigma_{0}^{2}\min\left(n^{-\frac{s\beta}{1+s\beta}},d/n\right)$. When $d=\infty$, this upper bound matches the well-known optimal rate under the source condition and the polynomial eigen-decay condition. When $d<\infty$, there is a phase transition around $d_{0}\asymp n^{\frac{1}{1+s\beta}}$: if $d\lesssim d_{0}$, the upper bound is $d\sigma_{0}^{2}/n$; if $d\gtrsim d_{0}$, the upper bound is the same as if $d=\infty$. Using the span profile, we can extend classical results to finite-dimensional models and reveal new phenomena.
Example A.4 (Polynomial signals ($\alpha>1$)).
Suppose $\theta_{i}^{*}=i^{-\alpha/2}$ for some constant $\alpha>1$, and $\{\lambda_{i}\}_{1}^{d}$ are decreasing. By an integral approximation, we can get $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\leq\frac{1}{\alpha-1}\,k^{-\alpha}$. Therefore, we have $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma^{2})\;\lesssim\;\,[\sigma^{2}]^{-\frac{1}{\alpha}}$. The optimal risk of PC estimator satisfies
$\mathcal{R}_{*}^{\operatorname{PC}}\;\leq\;2\sigma^{2}\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma^{2})\;\lesssim\;\min\left(\,[\sigma^{2}]^{\,1-\frac{1}{\alpha}},d\sigma^{2}\right).$
Example A.5 (Polynomial signals ($\alpha=1$)).
Suppose $d<\infty$, $\theta_{i}^{*}=i^{-1/2}$, and $\{\lambda_{i}\}_{1}^{d}$ are decreasing. We show in the supplementary material that for some constant $C$, if $d\sigma^{2}\leq e$, then $\mathcal{R}_{*}^{\operatorname{PC}}\leq Cd\sigma^{2}$, and if $d\sigma^{2}>e$, then $\mathcal{R}_{*}^{\operatorname{PC}}\leq C\log(d\sigma^{2}/\log(d\sigma^{2}))$.
Example A.6 (Polynomial signals ($\alpha<1$)).
Suppose $d<\infty$, $\theta_{i}^{*}=i^{-1/2}$, and $\{\lambda_{i}\}$ is decreasing. We show in the supplementary material that $\mathcal{R}_{*}^{\operatorname{PC}}\;\lesssim\;d\min\left(d^{-\alpha},\sigma^{2}\right)$.
These examples suggest that using our framework of span profile, we are able not only to recover classical results but also to extend it to various settings where the classical framework is inapplicable.
Details of Example˜A.5.
We have $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\leq k^{-1}\int_{k}^{d}\frac{1}{x}dx=k^{-1}\bigl(\log d-\log k\bigr)$.
By dropping the term $\log k$ in the numerator, it is easy to see that a sufficient condition for $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\leq\sigma^{2}$ is given by $k\geq\sigma^{-2}\log(d)$. Therefore, we have $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma^{2})\;\leq\;\lceil\sigma^{-2}\log(d)\rceil$.
The upper bound can be improved. Suppose $A>1$ satisfies $d\sigma^{2}\leq A\log A$. If $k\geq\sigma^{-2}\log A$, then
$\frac{k}{d}\geq\frac{\log A}{d\sigma^{2}}\geq\frac{d\sigma^{2}/A}{d\sigma^{2}}=\frac{1}{A},$
which follows that $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\leq k^{-1}\log A\leq\sigma^{2}$. Therefore,
$\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma^{2})\leq\min\left(d,\;\lceil\sigma^{-2}\log A\rceil\right).$
By elementary calculus, if $y>e$, the solution to $x\log x=y$ satisfies that $x\in(e,y)$, and thus $\log x\in(1,\log\left(y\right))$, which implies $x>y/\log\left(y\right)$ and thus $x<y/\log\left(y/\log\left(y\right)\right)=y/\left(\log y-\log\log y\right)<2y/\log(y)$.
If $d\sigma^{2}\leq e$, we can take $A=e$ and conclude
$\mathcal{R}_{*}^{\operatorname{PC}}\;\leq\;2\sigma^{2}\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma^{2})\;\lesssim\;d\sigma^{2}.$
If $d\sigma^{2}>e$, then $\log(d\sigma^{2})>1$ and we can take $A=2d\sigma^{2}/\log\left(d\sigma^{2}\right)$, which implies that
$\mathcal{R}_{*}^{\operatorname{PC}}\;\leq\;2\sigma^{2}\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma^{2})\;\lesssim\;\log\left(d\sigma^{2}\right)-\log\left(\log\left(d\sigma^{2}\right)\right).$
Detail of Example˜A.6.
By an integral approximation, we see that
$\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\;\asymp\;k^{-1}\bigl(d^{\,1-\alpha}-k^{\,1-\alpha}\bigr).$
Case 1:
$\sigma^{2}d^{\,\alpha}<2$. We have the default bound $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma^{2})\leq d$.
Case 2:
$\sigma^{2}d^{\,\alpha}\geq 2$. If $k\geq d^{1-\alpha}/\sigma^{2}$, then $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)\leq\sigma^{2}$. Therefore, we have $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma^{2})\leq\lceil d^{1-\alpha}/\sigma^{2}\rceil$, which is not larger than $\lceil d/2\rceil$.
Combining both cases, we have $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma^{2})\lesssim d\min(1/(d^{\alpha}\sigma^{2}),1)$. Multiplying by $2\sigma^{2}$ on both sides, we have
$\mathcal{R}_{*}^{\operatorname{PC}}\;\leq 2\sigma^{2}\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma^{2})\;\lesssim\;d\min\left(d^{-\alpha},\sigma^{2}\right).$
A.4 Detail of Example˜4.4
Let $f(x)=\sigma_{0}^{2}xe^{-x^{b}}$. Then $(\theta^{*}_{j+1})^{2}=f(j)-f(j+1)$ for $j\geq 1$. Since $\theta_{1}^{*}=0$, for any $k\geq 1$, the tail sum is
$\sum_{j=k+1}^{\infty}(\theta^{*}_{j})^{2}=\sum_{j=k}^{\infty}(f(j)-f(j+1))=f(k)=\sigma_{0}^{2}ke^{-k^{b}},$
since $f(N)\to 0$. As $\{\lambda_{j}\}$ is assumed to be decreasing, the trade-off function is $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)=\frac{1}{k}\sum_{j=k+1}^{\infty}(\theta^{*}_{j})^{2}=\sigma_{0}^{2}e^{-k^{b}}$.
For any $n\geq 3$, let $k=K_{n}$. By definition of the ceiling function, $k\geq(\log n)^{1/b}$, which implies $k^{b}\geq\log n$, and thus $e^{k^{b}}\geq n$. Then, $\mathbf{H}_{\bm{\theta}^{*},\bm{\lambda}}(k)=\sigma_{0}^{2}e^{-k^{b}}\leq\sigma_{0}^{2}/n$. By Proposition˜3.6, we have $\mathbf{D}_{\bm{\theta}^{*},\bm{\lambda}}(\sigma_{0}^{2}/n)\leq k=K_{n}$.
Since this holds for all sufficiently large $n$, we conclude that $\bm{\theta}^{*}\in\mathcal{F}_{\bm{K}}$. Theorem˜4.3 guarantee the optimal convergence rate is $\Theta(\sigma_{0}^{2}K_{n}/n)=\Theta(\sigma_{0}^{2}(\log n)^{1/b}/n)$.
Lastly, we consider the standard source condition that for some $s>0$, there is some constant $R_{s}$ such that
$\sum_{j=1}^{\infty}\lambda_{j}^{-s}(\theta_{j}^{*})^{2}\leq R_{s}.$ (35)
Let’s assume a polynomial eigenvalue decay $\lambda_{j}\asymp j^{-\gamma}$ for some $\gamma>0$. Let $S$ be the left hand side of Equation˜35. Since $\theta_{1}^{*}=0$, we have
$S=\sum_{j=2}^{\infty}(j^{-\gamma})^{-s}(\theta_{j}^{*})^{2}$ $=\sum_{j=2}^{\infty}j^{s\gamma}(\theta_{j}^{*})^{2}$
$=\sum_{k=1}^{\infty}(k+1)^{s\gamma}(\theta_{k+1}^{*})^{2}.$
Using $(\theta_{k+1}^{*})^{2}=f(k)-f(k+1)$ with $f(x)=\sigma_{0}^{2}xe^{-x^{b}}$:
$S=\sum_{k=1}^{\infty}(k+1)^{s\gamma}(f(k)-f(k+1)).$
Using summation by part, we have
$S=(1+1)^{s\gamma}f(1)-\lim_{N\to\infty}(N+1)^{s\gamma}f(N+1)+\sum_{k=1}^{\infty}((k+2)^{s\gamma}-(k+1)^{s\gamma})f(k+1).$
Since $\lim_{N\to\infty}(N+1)^{s\gamma}Ne^{-N^{b}}=0$ for $b\geq 1$, the limit term vanishes. $f(1)=\sigma_{0}^{2}e^{-1}$. The difference term $(k+2)^{s\gamma}-(k+1)^{s\gamma}>0$. $f(k+1)=\sigma_{0}^{2}(k+1)e^{-(k+1)^{b}}>0$. The sum $\sum_{k=1}^{\infty}((k+2)^{s\gamma}-(k+1)^{s\gamma})f(k+1)$ converges because $f(k+1)$ decays faster than any polynomial grows. Specifically, $(k+2)^{s\gamma}-(k+1)^{s\gamma}\approx s\gamma k^{s\gamma-1}$, and the sum $\sum k^{s\gamma-1}(k+1)e^{-(k+1)^{b}}$ converges. Therefore, $S$ converges for any $s>0$ and any $\gamma>0$.
The classical theory predicts a rate of $n^{-\frac{s\gamma}{s\gamma+1}}$. Since the source condition holds for arbitrarily large $s$, the classical rate can be made arbitrarily close to $n^{-1}$. However, this $n^{-1}$ rate ignores the logarithmic factor $(\log n)^{1/b}$ present in the true optimal rate $\Theta(\sigma_{0}^{2}(\log n)^{1/b}/n)$. Thus, the traditional convergence analysis based on the source condition is not sharp for this signal.
Appendix B Proofs for results in Section˜9
Proof of Proposition˜9.3.
Upper bound: Take $k=d^{\dagger}$, and we have $B(k)=k\mathbf{H}_{f^{*},\lambda}(k)\leq k\sigma^{2}$. The variance $V(k)=\sum_{j=1}^{k}(\sigma_{0}^{2}+\tau_{j}^{2})/n\leq k(\sigma_{0}^{2}+\sigma_{f,4}^{2})/n=k\sigma^{2}$. Thus $\mathcal{R}_{*}^{\operatorname{PC}}\leq\mathcal{R}_{k}=B(k)+V(k)\leq 2k\sigma^{2}=2d^{\dagger}\sigma^{2}$.
Lower bound: Let $k^{*}$ be the optimal tuning parameter. If $k^{*}\geq d^{\dagger}$, then $\mathcal{R}_{*}\geq d^{\dagger}\sigma_{0}^{2}/n$. If $k^{*}\leq d^{\dagger}-1$, by definition of ESD, we have $\mathcal{R}_{*}\geq B(k^{*})\geq B(d^{\dagger}-1)\geq(d^{\dagger}-1)\sigma^{2}\geq(d^{\dagger}-1)\sigma_{0}^{2}/n$.
Proof of Theorem˜9.5.
Upper bound: The upper bound follows the proof of the upper bound in Proposition˜9.3. To see this, we note that since $f^{*}\in\mathcal{F}_{K,\bm{\lambda},n}$, we have $\|f\|_{\infty}^{2}\leq\sigma_{0}^{2}C_{0}^{2}$. Therefore, $\sigma^{2}\leq\bar{\sigma}^{2}/n$. We can then apply the argument in Proposition˜9.3 with $\sigma^{2}$ replaced by $\bar{\sigma}^{2}/n$.
Lower bound: We establish the lower bound using Assouad’s method.
Let $m=\lfloor c_{1}K\rfloor$. Consider the first $m$ eigenfunctions $\{\psi_{\pi_{j}}\}_{j\leq m}$ corresponding to the largest eigenvalues $\{\lambda_{\pi_{j}}\}_{j\leq m}$. Define the collection of hypercubes vertices $\mathcal{V}=\{-1,1\}^{m}$. For every vertex $v\in\mathcal{V}$, define a function
$f^{(v)}(x)=\gamma\sum_{j=1}^{m}v_{j}\psi_{\pi_{j}}(x),$ (36)
where the amplitude $\gamma$ is to be chosen. Since $\mathbf{k}$ is $(K,n)$-regular, we have
$f^{(v)}(x)^{2}\leq\gamma^{2}\sum_{j\leq m}\lambda_{\pi_{j}}^{-1}\sum_{j\leq m}\lambda_{j}\psi^{2}_{\pi_{j}}(x)\leq\gamma^{2}C_{1}n\kappa^{2},$ (37)
where $\kappa^{2}=\sup_{x}\mathbf{k}(x,x)<\infty$ by assumption.
We choose
$\gamma^{2}=n^{-1}\min\left(\frac{\bar{\sigma}^{2}}{4(1+C_{0}^{2})},\frac{\sigma_{0}^{2}C_{0}^{2}}{C_{1}\kappa^{2}}\right).$
It then follows that $\|f^{(v)}\|_{\infty}^{2}\leq\sigma_{0}^{2}C_{0}^{2}$.
For each $v\in\mathcal{V}$, let $P_{v}$ be the sampling distribution of $\{z_{i}=(x_{i},y_{i})\}_{i\leq n}$ from the regression model Equation˜21 with $f^{*}=f^{(v)}$. Let $\rho$ be the Hamming distance on $\mathcal{V}$. If $v$ and $w\in\mathcal{V}$ differ in exactly one coordinate (i.e., $\rho(v,w)=1$), then
  • $\|f^{(v)}-f^{(w)}\|_{L^{2}(\mu)}^{2}\geq(2\gamma)^{2}$, and
  • the Kullback-Leibler divergence between $P_{v}$ and $P_{w}$ satisfies $\operatorname{KL}\bigl(P_{v}\,\|\,P_{w}\bigr)=\frac{n}{2\sigma_{0}^{2}}(2\gamma)^{2}\leq\tfrac{1}{2}$, where the last equation is due to the definition of the constant $c$. By the Pinsker’s inequality, $\|P_{v}\wedge P_{w}\|=1-\operatorname{TV}(P_{v},P_{w})\geq 1-\sqrt{\operatorname{KL}\!\bigl(P_{v}\,\|\,P_{w}\bigr)/2}=1/2$.
By Assouad’s Lemma (Lemma 2 in Yu [49]), for any estimator $\widehat{f}$ based on a sample $\{z_{i}=(x_{i},y_{i})\}_{i\leq n}$ drawn from $P_{v}$, we have
$\sup_{v\in\mathcal{V}}\mathbb{E}_{v}\bigl\lVert\widehat{\bm{\theta}}-\bm{\theta}^{(v)}\bigr\rVert^{2}\geq m\frac{(2\gamma)^{2}}{4}=c\frac{\sigma_{0}^{2}K}{n},$
where $c$ is a constant that depends on $C_{0}$, $\kappa$, $c_{1}$, $C_{1}$.
Proof of Theorem˜9.6.
Since
$\mathcal{F}_{\mathbf{K},\mathbf{k}}=\bigcap_{n\geq n_{0}}\mathcal{F}_{K_{n},\mathbf{k}}^{(n)},$
the upper bound $\bar{\sigma}^{2}K_{n}/n$ is immediately implied by Theorem˜9.5.
The lower bound follows the same argument as in the proof of Theorem˜4.3, but replace the construction of parameter vectors in Equation˜33 by the construction of functions in Equation˜36. Following the proof for Theorem˜4.3, we use ˜4.1 to ensure the constructed functions all belong to $\mathcal{F}_{\mathbf{K},\mathbf{k}}$. Then the lower bound is given using Assouad’s Lemma as in the proof of Theorem˜9.5. Below, we provide the details for completeness.
Mercer’s theorem yields
$\mathbf{k}\left(\bm{x},\bm{x}^{\prime}\right)=\sum_{j=1}^{\infty}\lambda_{j}\psi_{j}(\bm{x})\psi_{j}\left(\bm{x}^{\prime}\right),\quad\bm{x},\bm{x}^{\prime}\in\mathcal{X},\vskip 0.0pt$ (38)
where $\{\psi_{j}\}_{j\geq 1}$ is a $L^{2}(\mathcal{X},\mu)$-orthonormal eigenbasis. Without loss of generality, assume $\lambda_{j}$ is sorted decreasingly.
Fix $n$ and set $m:=\lfloor c_{1}K_{n}\rfloor$ where $c_{1}$ comes from ˜9.4.
For a sign vector $v=(v_{j})_{j\leq m}\in\{-1,+1\}^{m}$, define the sequence of coefficients as
$\theta_{j}^{(v)}\ :=\ \begin{cases}\gamma\,v_{j},&j\leq m,\\ 0,&j>m,\end{cases}\qquad f_{v}(x)\ :=\ \sum_{j\geq 1}\theta_{j}^{(v)}\,\psi_{j}(x)\ =\gamma\sum_{j\leq m}v_{j}\,\psi_{j}(x).$
Since $\mathbf{k}$ is $(K_{n},n)$-regular, Equation˜37 holds and reads as
$f^{(v)}(x)^{2}\leq\gamma^{2}C_{1}n\kappa^{2}.$
If $\gamma_{2}C_{1}n\kappa^{2}\leq\sigma_{0}^{2}C_{0}^{2}$, then $\|f^{(v)}\|_{\infty}^{2}\leq\sigma_{0}^{2}C_{0}^{2}$. Furthermore, if $m\gamma^{2}\leq(2n)^{-1}\sigma_{0}^{2}K_{n}$, we can the same argument in Lemma˜A.2 (in particular, using ˜4.1 to dervie Equation˜34) to show that $f^{(v)}\in\mathcal{F}_{\mathbf{K},\mathbf{k}}$.
We choose
$\gamma^{2}=n^{-1}\min\left(\sigma_{0}^{2},\frac{\bar{\sigma}^{2}}{4(1+C_{0}^{2})},\frac{\sigma_{0}^{2}C_{0}^{2}}{C_{1}\kappa^{2}}\right),$
which implies $f^{(v)}\in\mathcal{F}_{\mathbf{K},\mathbf{k}}$.
We then follows the same argument in the proof of lower bound in Theorem˜9.5 to obtain
$\sup_{v\in\mathcal{V}}\mathbb{E}_{v}\bigl\lVert\widehat{\bm{\theta}}-\bm{\theta}^{(v)}\bigr\rVert^{2}\geq m\frac{(2\gamma)^{2}}{4}=c\frac{\sigma_{0}^{2}K}{n},$
where $c$ is a constant that depends on $C_{0}$, $\kappa$, $c_{1}$, $C_{1}$.
B.1 Proofs for results in Section˜5
Proof of Proposition˜5.1.
We note that for the top $k$ terms, $\lambda_{(i)}\geq\nu$, so $\frac{\lambda_{(i)}}{\lambda_{(i)}+\nu}\geq\frac{1}{2}$ and hence
$\sum_{i=1}^{k}\Bigl(\frac{\lambda_{(i)}}{\lambda_{(i)}+\nu}\Bigr)^{2}\geq\frac{k}{4}.$
For the tail terms, $\lambda_{(i)}<\nu$ for $i\geq k+1$, so $\frac{\lambda_{(i)}}{\lambda_{(i)}+\nu}\geq\frac{\lambda_{(i)}}{2\nu}$ and thus
$\sum_{i=k+1}^{d}\Bigl(\frac{\lambda_{(i)}}{\lambda_{(i)}+\nu}\Bigr)^{2}\geq\frac{1}{4\nu^{2}}\sum_{i=k+1}^{d}\lambda_{(i)}^{2}.$
Combining the two cases, we have
$V_{\mathrm{R}}(\nu)$ $=\sigma^{2}\sum_{i=1}^{d}\Bigl(\frac{\lambda_{(i)}}{\lambda_{(i)}+\nu}\Bigr)^{2}$
$\geq\sigma^{2}\Biggl(\frac{k}{4}+\frac{1}{4v^{2}}\sum_{i=k+1}^{d}\lambda_{(i)}^{2}\Biggr)$
$\geq\frac{1}{4}\sigma^{2}\Biggl(k+\sum_{i=k+1}^{d}\Bigl(\frac{\lambda_{(i)}}{\lambda_{(k)}}\Bigr)^{2}\Biggr).$
Proof of Proposition˜5.3.
For any $\nu\leq\lambda_{(i_{0})}$, we have $\lambda_{(i)}+\nu\leq 2\lambda_{(i)}$ for all $i\leq i_{0}$. We then have
$B_{\mathrm{R}}(\nu)\geq\sum_{i=1}^{i_{0}}\Bigl(\frac{\nu}{\lambda_{(i)}+\nu}\Bigr)^{2}(\theta^{*}_{\pi_{i}})^{2}\geq\frac{\nu^{2}}{4}\sum_{i=1}^{i_{0}}\frac{(\theta^{*}_{\pi_{i}})^{2}}{\lambda_{(i)}^{2}}\geq c_{r}\,\nu^{2},$
where the last inequality is due to ˜5.2. ∎
Proof of Theorem˜5.6.
Fix $\nu\in(0,\lambda_{(i_{0})}]$ and let $k:=k_{\bm{\lambda}}(\nu)$. By definition of $k_{\bm{\lambda}}(\nu)$, we have $\lambda_{(k)}\geq\nu>\lambda_{(k+1)}$ and, since $\nu\leq\lambda_{(i_{0})}$, we have $k\geq i_{0}$.
Under ˜5.2, for any $\nu\leq\lambda_{(i_{0})}$ we already showed in Equation˜12 that
$B_{\mathrm{R}}(\nu)\geq c_{r}\nu^{2}.$
Proposition˜5.1 implies that
$V_{\mathrm{R}}(\nu)\geq\frac{1}{4}\widetilde{V}(k).$
Combining the above two lower bounds, we have
$\mathcal{R}_{\mathrm{R}}(\nu)=B_{\mathrm{R}}(\nu)+V_{\mathrm{R}}(\nu)\geq\max\Bigl\{B_{\mathrm{R}}(\nu),V_{\mathrm{R}}(\nu)\Bigr\}\geq\max\Bigl\{c_{r}\nu^{2},\frac{1}{4}\widetilde{V}(k)\Bigr\}.$
Recall
$d^{\Delta}=\max\Bigl\{m:\ c_{r}\,\lambda_{(m)}^{2}>\sigma^{2}\,m\bigl(1+\widetilde{N}(m,d;\bm{\lambda})\bigr)\Bigr\}=\max\Bigl\{m:\ c_{r}\,\lambda_{(m)}^{2}>\widetilde{V}(m)\Bigr\}.$
We consider two cases.
Case 1: $k\geq d^{\Delta}$. Then $\widetilde{V}(k)\geq\sigma^{2}k\geq\sigma^{2}d^{\Delta}$, so
$\mathcal{R}_{\mathrm{R}}(\nu)\geq\frac{1}{4}\widetilde{V}(k)\geq\frac{1}{4}\sigma^{2}d^{\Delta}.$
Case 2: $k\leq d^{\Delta}-1$. Since $\nu>\lambda_{(k+1)}$ and $k+1\leq d^{\Delta}$, monotonicity of the spectrum gives $\lambda_{(k+1)}\geq\lambda_{(d^{\Delta})}$, hence $\nu\geq\lambda_{(d^{\Delta})}$ and
$\mathcal{R}_{\mathrm{R}}(\nu)\geq c_{r}\nu^{2}\geq c_{r}\lambda_{(d^{\Delta})}^{2}.$
By definition of $d^{\Delta}$, we have $c_{r}\lambda_{(d^{\Delta})}^{2}>\widetilde{V}(d^{\Delta})\geq\sigma^{2}d^{\Delta}$, so
$\mathcal{R}_{\mathrm{R}}(\nu)\geq c_{r}\lambda_{(d^{\Delta})}^{2}>\sigma^{2}d^{\Delta}\geq\frac{1}{4}\sigma^{2}d^{\Delta}.$
In both cases, $\mathcal{R}_{\mathrm{R}}(\nu)\geq\frac{1}{4}\sigma^{2}d^{\Delta}$. Taking the infimum over $\nu\in(0,\lambda_{(i_{0})}]$ yields the second display. ∎
Appendix C Proofs for results on overparameterized gradient flow
In this section, we prove Theorem 6.2. The high-level idea is as follows: To show the ESD decreases, it is enough to show that the squared signal tail sorted by the learned eigenvalues at the new time is smaller than that at the old time. The key idea is to study how the gradient flow changes the eigenvalues depending on the signal’s strength. Our analysis reveals that eigenvalues associated with the strong signal coordinates will often grow much faster than those associated with weak ones. Consequently, more of the largest learned eigenvalues correspond to the strong signals. This implies that the signal energy is concentrated in the top principal components of the learned kernel, which reduces the signal tail and thus reduces the ESD.
We first remark that for any $j\in[d]$, due to the same initialization $b_{j,k}=b_{0}$ for all $k$, one can prove that throughout the time $b_{j,k}$ (for all $k$) have the same value $b_{j}$. Therefore, we can rewrite the over-parameterization as $\theta_{j}=a_{j}b_{j}^{D}\beta_{j}$, and consider the following gradient flow
$\dot{a}_{j}=-\nabla_{a_{j}}L_{j}=b_{j}^{D}\beta_{j}(z_{j}-\theta_{j}),$ (39)
$\dot{b}_{j}=-\nabla_{b_{j}}L_{j}=Da_{j}b_{j}^{D-1}\beta_{j}(z_{j}-\theta_{j}),$
$\dot{\beta}_{j}=-\nabla_{\beta_{j}}L_{j}=a_{j}b_{j}^{D}(z_{j}-\theta_{j}),$
$a_{j}(0)=\lambda_{j}^{\frac{1}{2}}>0,\quad\bm{b}(0)=b_{0}>0,\quad\bm{\beta}(0)=0,$
where $L_{j}=\frac{1}{2}(z_{j}-\theta_{j})^{2}$.
C.1 Proof of Theorem 6.2
Recall that we define these sets as follows:
$A_{1}:$ $=\{i:\pi_{t_{1}}^{-1}(i)<d^{\dagger}(t_{1}),\lambda_{i}<c\cdot D^{-\frac{D}{D+2}}\cdot M^{\frac{2}{D+2}},|\theta_{i}^{*}|<\tilde{\sigma}\}$;
$A_{2}:$ $=\{i:\pi_{t_{1}}^{-1}(i)<d^{\dagger}(t_{1}),|\theta_{i}^{*}|>M\};$
$A_{3}:$ $=\{i:\pi_{t_{1}}^{-1}(i)<d^{\dagger}(t_{1}),|\theta_{i}^{*}|<\tilde{\sigma}\}\setminus A_{1}.$
and
$B_{1}:=\{i:\pi_{t_{1}}^{-1}(i)>d^{\dagger}(t_{1}),|\theta_{i}^{*}|<\tilde{\sigma}\};~~~B_{2}:=\{i:\pi_{t_{1}}^{-1}(i)>d^{\dagger}(t_{1}),|\theta_{i}^{*}|>M\}.$
where $\tilde{\sigma}^{2}\leq\min\{\frac{|B_{1}|}{C_{B_{1}}}\varepsilon^{2},c^{\prime}\varepsilon^{2}\}$, $C_{B_{1}}=\min\{1\vee(|A_{1}|-|B_{1}|),|B_{2}|\}$ and $c^{\prime}$ is a constant $\leq 1$. Also recall from the assumption of Theorem˜6.2 that suppose $|\theta^{*}_{i}|>|\theta^{*}_{j}|>M$, then if $C_{max}M>|\theta^{*}_{i}|$, $\eta_{i,j}=|\theta_{i}^{*}|-|\theta_{j}^{*}|>C_{\eta}\varepsilon^{\prime}$; otherwise, $\frac{|\theta_{i}^{*}|}{|\theta_{j}^{*}|}>(1+\frac{c_{\eta}}{D})$.
[…]
Throughout the proof, we assume all the events $\{E_{k}\}$ in Lemma C.9 hold. We divide the proof into several parts.
Part 1: Very small eigenvalues can be ignored. From ˜6.1, we have $\inf_{j\in S}\lambda_{j}>n^{-\delta}$, where $\delta$ is a constant. For $i\in R$, if $\lambda_{i}<n^{-(2.1\delta\vee 5)}$, Proposition˜C.2 implies that at $t_{2}$, we have
$\tilde{\lambda}_{i}(t_{2})<n^{1.1}\cdot\lambda_{i}^{0.99}<n^{-\delta}.$
If $B_{2}$ is empty, then all the signals in $R$ is 0 by the definition of $\tilde{\sigma}$. Otherwise, for $\tilde{\lambda}_{i}<n^{-\delta}$, and $\lambda_{\pi_{t_{1}}(d^{\dagger}(t_{1}))}>n^{-\delta}$. According to the monotonicity of eigenvalues in Lemma˜C.4, any the index $i$ such that $\lambda_{i}<n^{-(2.1\delta\vee 5)}$ can not rank among the first $d^{\dagger}(t_{1})$ at time $t_{2}$, i.e., it makes no difference to the variation of $d^{\dagger}$ from $t_{1}$ to $t_{2}$.
Part 2: Analysis for $B_{2}$
If $j\in B_{2}$, by Proposition C.1, we have $\left|\theta_{j}(t_{2})-\theta_{j}^{*}\right|<\varepsilon^{\prime}\leq\frac{1}{C_{M}}M$. We apply Equation˜51 to get $\beta^{2}=a^{2}-\lambda<a^{2}$ and $\beta^{2}=D^{-1}(b^{2}-b_{0}^{2})<D^{-1}b^{2}$, and thus
$|\theta_{j}(t)|=a_{j}(t)b_{j}^{D}(t)|\beta_{j}(t)|\leq a_{j}(t)b_{j}^{D}(t)\cdot a_{j}^{\frac{1}{D+1}}(t)\cdot D^{-\frac{D}{2(D+1)}}b_{j}^{\frac{D}{D+1}}(t)=D^{-\frac{D}{2(D+1)}}\left[a_{j}(t)b_{j}^{D}(t)\right]^{\frac{D+2}{D+1}}.$
Therefore, at time $t=t_{2}$, we have for some constants $c$ and $C$ that
$D^{-\frac{D}{2(D+1)}}\left[a_{j}(t)b_{j}^{D}(t)\right]^{\frac{D+2}{D+1}}\geq\left|\theta_{j}(t)\right|\geq(1-c)M,\quad\Longrightarrow\quad a_{j}(t)b_{j}^{D}(t)\geq C\cdot D^{\frac{D}{2(D+2)}}M^{\frac{D+1}{D+2}}.$
It follows that $\tilde{\lambda}_{j}(t_{2})\geq cD^{\frac{D}{(D+2)}}M^{\frac{2(D+1)}{D+2}}$. Moreover, for $\lambda_{j}$ and $D^{-1}b_{0}^{2}$ that are much smaller than $c\cdot D^{-\frac{D}{D+2}}\cdot M^{\frac{2}{D+2}}$, we use Equation˜51 to obtain
$\beta_{j}^{2}\asymp a_{j}^{2}\asymp D^{-1}b_{j}^{2}\asymp D^{-\frac{D}{D+2}}M^{\frac{2}{D+2}}.$ (40)
Part 3: $A_{1}$ and $B_{2}$ will exchange
In the following, suppose $i\in A_{1}$ and $j\in B_{2}$. We will prove $\tilde{\lambda}_{i}(t_{2})<\tilde{\lambda}_{j}(t_{2})$ by contradiction.
If $\tilde{\lambda}_{i}(t_{2})\geq\tilde{\lambda}_{j}(t_{2})$ , then by Proposition C.1, we have
$|\theta_{j}(t_{2})-\theta_{j}^{*}|<2\varepsilon^{\prime}.$
By Lemma C.4, we have
$|\theta_{i}(t_{2})|<|\theta_{i}^{*}|+\kappa_{i}.$
We have $|\theta_{j}|>C_{D}|\theta_{i}|$ where $C_{D}>1+\frac{c}{D}$ for $M=C_{M}\varepsilon>C_{M}|\kappa_{i}|$. At $t_{2}$, the following holds:
$|\beta_{j}(t_{2})|>C_{D}|\beta_{i}(t_{2})|.$
It follows that $a_{i}^{2}(t_{2})b_{i}^{2D}(t_{2})>a_{j}^{2}(t_{2})b_{j}^{2D}(t_{2})$. Combined with Equation˜51, we have
$\frac{\beta_{i}^{2}(t_{2})+\lambda_{i}}{\beta_{j}^{2}(t_{2})+\lambda_{j}}>\left(\frac{D\beta_{j}^{2}(t_{2})+b_{0}^{2}}{D\beta_{i}^{2}(t_{2})+b_{0}^{2}}\right)^{D}>\left(\frac{C_{D}}{1+\delta}\right)^{D}=(1+c(C_{D}-1))^{D}.$ (41)
Recall that $|\beta_{i}(t_{2})|<\frac{1}{C_{D}}|\beta_{j}(t_{2})|$ and $\beta_{j}^{2}(t_{2})>CD^{-\frac{2}{D+2}}M^{\frac{2}{D+2}}$. If we choose the constant $C_{D}$ such that $(1+c(C_{D}-1))^{D}$ is large enough, the inequality Equation˜41 will implies $\lambda_{i}$ larger than its upper bound in the definition of set $A_{1}$. (We can let $C_{D}=1+c\cdot\frac{1}{D}$.)
The contraction shows that $\tilde{\lambda}_{i}(t_{2})<\tilde{\lambda}_{j}(t_{2})$ for any $i\in A_{1}$, $j\in B_{2}$. If the sets $A_{1}$ and $B_{2}$ are not empty when $t=t_{1}$, then from $t_{1}$ to $t_{2}$, the elements of set $B_{2}$ will be arranged before those of set $A_{1}$ according to the eigenvalue index. We only need $\lambda_{i}<c\cdot D^{-\frac{D}{D+2}}M^{\frac{2}{D+2}}$.
For the same reason, the elements of set $A_{2}$ will be arranged before those of set $A_{1}$ at $t_{2}$.
Part 4: $A_{2}$ and $B_{2}$ will be monotonously nonincreasing
In the following, W.L.O.G we assume $\theta^{*},\theta^{*}>0$. We prove that given $i\in A_{2}$, $j\in B_{2}$, if $\theta_{i}^{*}>\theta_{j}^{*}$, we have $\tilde{\lambda}_{i}(t_{2})>\tilde{\lambda}_{j}(t_{2})$.
If $\theta_{i}^{*}-\theta_{j}^{*}>C_{\eta}\varepsilon$, we have $z_{i}>z_{j}$. Then by Proposition C.1, we have at $t_{2}$
$|\theta_{i}(t_{2})-\theta_{i}^{*}|<2\varepsilon^{\prime},~~~|\theta_{j}(t_{2})-\theta_{j}^{*}|<2\varepsilon^{\prime}.$
If $\lambda_{i}\geq\lambda_{j}$, by Equation˜39 and monotonicity, we always have $\tilde{\lambda}_{i}(t)>\tilde{\lambda}_{j}(t)$.
Now, consider the case where $\lambda_{i}<\lambda_{j}$. By the definition of $A_{1}$ and $B_{2}$, we have $\lambda_{j}<c\cdot D^{-\frac{D}{D+2}}\cdot M^{\frac{2}{D+2}}$. Next, we use proof by contradiction. Note that
$\frac{\theta_{i}(t_{2})}{\theta_{j}(t_{2})}=\frac{\tilde{\lambda}_{i}^{\frac{1}{2}}(t_{2})\beta_{i}(t_{2})}{\tilde{\lambda}_{j}^{\frac{1}{2}}(t_{2})\beta_{j}(t_{2})}\geq C_{D}.$
If $\tilde{\lambda}_{i}(t)\leq\tilde{\lambda}_{j}(t)$, then $\beta_{i}(t_{2})>C_{D}\beta_{j}(t_{2})$. By Equation˜40, both $\beta_{i}(t_{2})$ and $\beta_{j}(t_{2})$ are much larger then $D^{-\frac{1}{2}}b_{0}$. It then follows that
$\left(\frac{D\beta_{i}^{2}(t_{2})+b_{0}^{2}}{D\beta_{j}^{2}(t_{2})+b_{0}^{2}}\right)^{D}>C^{*}$
where $C^{*}$ is a constant that can be made large enough by choosing $C_{D}$. Therefore, using the same reason as in Equation˜41, we can get
$\frac{\beta_{i}^{2}(t_{2})+\lambda_{i}}{\beta_{j}^{2}(t_{2})+\lambda_{j}}<\frac{1}{C^{*}},$
which is impossible because $\beta_{i}(t_{2})>C_{D}\beta_{j}(t_{2})$ and $\beta_{i}^{2}(t_{2})>C\lambda_{j}$.
Part 5: $B_{1}$ will stay behind $A_{2}$, $A_{3}$ and $B_{2}$
In the following, let $j\in B_{1}$. By $y=\theta^{*}+\xi$, we have $|z_{j}|<\tilde{\sigma}+\varepsilon^{\prime}$. By Equation˜54 and $|\theta_{j}|<|z_{j}|$, we have
$|\beta_{j}(t_{2})|\leq\left(D^{-D/2}|z_{j}|\right)^{1/(D+2)}.$
Then by $\tilde{\lambda}_{j}(t_{2})=(\beta_{j}^{2}(t_{2})+\lambda_{j})(D\beta_{j}^{2}(t_{2})+b_{0}^{2})^{D}<(1+\frac{1}{c\cdot D})^{D}b_{0}^{2D}(\beta_{j}^{2}(t_{2})+\lambda_{j})$. Then if $\lambda_{j}<c\cdot D^{-D/(D+2)}\varepsilon^{2/(D+2)}$, given any $i\in A_{3}\cup A_{2}\cup B_{2}$, $\tilde{\lambda}_{i}(t_{2})>\tilde{\lambda}_{j}(t_{2})$. Otherwise, we have $\lambda_{j}>C\cdot D^{-D/(D+2)}\varepsilon^{2/(D+2)}$, then we have $\tilde{\lambda}_{j}(t_{2})<(1+\delta)\tilde{\lambda}_{j}(t_{1})$, then by the definition of $d^{\dagger}$, at least $d^{\dagger}$ eigenvalues will larger than $\tilde{\lambda}_{j}(t_{2})$.
Part 6: $A_{3}$ will be ahead of $A_{1}$
by the same reason between $B_{1}$ and $A_{1}$, for given $j\in A_{3}$, $\tilde{\lambda}_{j}(t_{2})=(\beta_{j}^{2}(t_{2})+\lambda_{j})(D\beta_{j}^{2}(t_{2})+b_{0}^{2})^{D}<(1+\frac{1}{c\cdot D})^{D}b_{0}^{2D}(\beta_{j}^{2}(t_{2})+\lambda_{j})$, and $i\in A_{1}$, $\lambda_{i}<c\cdot D^{-D/(D+2)}\varepsilon^{2/(D+2)}$, combined with
$|\beta_{i}(t_{2})|\leq\left(D^{-D/2}|z_{i}|\right)^{1/(D+2)}.$
then we have $\tilde{\lambda}_{j}(t_{2})>\tilde{\lambda}_{i}(t_{2})$ for $c$ is small and $C$ is large enough.
Part 7: Ordering of the spectrum at $t_{2}$ and $d^{\dagger}(t_{1})\geq d^{\dagger}(t_{2})$
To show $d^{\dagger}(t_{2})\leq d^{\dagger}(t_{1})$, it suffices to show $\mathbf{H}_{\bm{\theta}^{*},\tilde{\bm{\lambda}}(t_{2})}(d^{\dagger}(t_{1}))<\mathbf{H}_{\bm{\theta}^{*},\tilde{\bm{\lambda}}(t_{2})}(d^{\dagger}(t_{1}))$, which is equivalent to prove the following difference
$\sum_{i:\pi_{t_{1}}^{-1}(i)>d^{\dagger}(t_{1})}|\theta^{*}_{i}|^{2}-\sum_{i:\pi_{t_{2}}^{-1}(i)>d^{\dagger}(t_{1})}|\theta^{*}_{i}|^{2}$ (42)
is nonnegative.
We will make use of $|A_{1}|+|A_{2}|+|A_{3}|=d^{\dagger}(t_{1})$ and consider two possible cases.
  • Case 1: $|B_{2}|\leq|A_{1}|$. Since $B_{2}\cup A_{2}$ is ahead of $A_{1}\cup B_{1}$, we can see that the eigenvalue of the last element of $B_{2}\cup A_{2}$ is among the top $d^{\dagger}$ ones. Because $A_{3}$ is ahead of $A_{1}\cup B_{1}$, so only some of $A_{1}$ is swapped to the later part. Also some of $B_{1}$ may arise ahead some of $A_{1}$ . Therefore, to analyze the ordering of eigenvalues $\tilde{\bm{\lambda}}(t_{2})$, we define
    $B_{11}$ $=\{i\in B_{1}:\pi^{-1}_{t_{2}}(i)\leq d^{\dagger}(t_{1})\}.$ (43)
    $A_{11}$ $=\{i\in A_{1}:\pi^{-1}_{t_{2}}(i)>d^{\dagger}(t_{1})\}.$
    Here $A_{11}$ contains all the elements that move from the top $d^{\dagger}(t_{1})$ part to the later part, while $B_{2}$ and $B_{11}$ are the elements that move from the later part to the top $d^{\dagger}(t_{1})$ part. Therefore, we have $|A_{11}|=|B_{2}|+|B_{11}|$. Let $C_{B_{1}}:=\min\{(|A_{1}|-|B_{2}|)_{+},|B_{1}|\}$. We have $|B_{11}|\leq C_{B_{1}}$.
    W.L.O.G., we can write divide $A_{11}$ into two subsets such that $|A_{111}|=|B_{2}|$ and $|A_{112}|=|B_{11}|$. We can then write Equation˜42 as
    $\|\theta^{*}_{B_{2}}\|_{2}^{2}+\|\theta^{*}_{B_{1}1}\|_{2}^{2}-\|\theta^{*}_{A_{111}}\|_{2}^{2}-\|\theta^{*}_{A_{112}}\|_{2}^{2}.$ (44)
    The exchange between $A_{1}$ and $B_{2}$ yields
    $\|\theta^{*}_{B_{2}}\|_{2}^{2}-\|\theta^{*}_{A_{111}}\|_{2}^{2}\geq|B_{2}|(M^{2}-\tilde{\sigma}^{2}),$ (45)
    and
    $\|\theta^{*}_{B_{11}}\|_{2}^{2}-\|\theta^{*}_{A_{112}}\|_{2}^{2}\geq-|B_{11}|\tilde{\sigma}^{2}\geq-C_{B_{1}}\tilde{\sigma}^{2}.$
    Note the assumption of Theorem˜6.2 that $(|B_{2}|+C_{B_{1}})\tilde{\sigma}^{2}\leq|B_{2}|M^{2}$. We add the last two inequalities Equations˜44 and 44 together to get
    $\sum_{i:\pi_{t_{1}}(i)>d^{\dagger}(t_{1})}|\theta^{*}_{i}|^{2}-\sum_{i:\pi_{t_{2}}(i)>d^{\dagger}(t_{1})}|\theta^{*}_{i}|^{2}\geq 0,$
    where $\geq$ becomes >> if $B_{2}\neq\emptyset$. By the definition of $d^{\dagger}$, $d^{\dagger}(t_{2})\leq d^{\dagger}(t_{1})$.
  • Case 2: $|B_{2}|>|A_{1}|$. If the eigenvalue of the last element of $B_{2}\cup A_{2}$ is among the top $d^{\dagger}$, we follow the exact same proof in Case 1.
    Now suppose that the eigenvalue of the last element of $B_{2}\cup A_{2}$ is in the later part. In this case, all elements $B_{1}$ and $A_{1}$ are in the later part. We first identify all the elements that fall in the later part at time $t_{2}$: in addition to all elements of $B_{1}$, the following
    $B_{21}$ $=\{i\in B_{2}:\pi^{-1}_{t_{2}}(i)>d^{\dagger}\},$
    $A_{11}$ $=\{i\in A_{1}:\pi^{-1}_{t_{2}}(i)>d^{\dagger}\},$
    $A_{21}$ $=\{i\in A_{2}:\pi^{-1}_{t_{2}}(i)>d^{\dagger}\},$
    $A_{31}$ $=\{i\in A_{3}:\pi^{-1}_{t_{2}}(i)>d^{\dagger}\}.$
    Note that at time $t_{1}$, the elements in the later part are in $B_{1}$, $B_{21}$, and $B_{22}:=B_{2}\setminus B_{21}$. Therefore, Equation˜42 can be written as
    $\|\theta_{B_{22}}\|^{2}-\|\theta_{A_{11}}\|^{2}-\|\theta_{A_{21}}\|^{2}-\|\theta_{A_{31}}\|^{2}.$ (46)
    By definition of $B_{2}$, $A_{1}$, and $A_{3}$, each squared element in $B_{2}$ is larger than that of both $A_{1}$ and $A_{3}$. In addition, since $A_{2}$ and $B_{2}$ will be monotonously nonincreasing, for any element in $B_{22}$, its squared signal will be no less than that of any element in $\theta_{A_{21}}$. Therefore, we conclude that Equation˜46 is nonnegative and will be positive if $B_{22}$ is not empty.
C.2 Generalized signal results by dynamic equation analysis
Proposition C.1 (Shrinkage monotonicity and shrinkage time).
Suppose all the events $\{E_{k}\}$ in Lemma C.9 hold. Let $\varepsilon=2(C_{\text{proxy}})^{-1/2}\sqrt{\frac{\ln n\tilde{d}}{n}\cdot\ln n},\varepsilon^{\prime}=2(C_{\text{proxy}})^{-1/2}\sqrt{\frac{\ln n}{n}}$.
For any $j\in S$ (as defined in ˜6.1), we have
$|\theta_{j}^{*}-\theta_{j}(t)|<2\varepsilon^{\prime}.\quad\forall t\geq t(\varepsilon)$ (47)
where $t(\varepsilon)=C~b_{0}^{-D}~\varepsilon^{-1}\ln n$ for some absolute constant $C$.
Proof.
When all the events $E_{k}$ in Lemma C.9 hold, we have
$\left\|\xi_{S}\right\|_{\infty}\leq\varepsilon^{\prime}.$ (48)
Consider $j\in S$. We have $\theta^{*}_{j}\geq 8\varepsilon^{\prime}$(We let $C_{M}\geq 8$). By taking $\delta=\varepsilon^{\prime}$ and also $\kappa=\varepsilon^{\prime}$ in Lemma C.7, we have
$|\theta_{j}^{*}-\theta_{j}(t)|\leq 2\varepsilon^{\prime},\quad\forall t\geq\bar{T}^{app}(\delta),$
with
$\bar{T}^{app}(\delta)\leq T^{\textrm{sig}}+C_{2}D^{\frac{D}{D+2}}(\theta_{j}^{*})^{-\frac{2D+2}{D+2}}\ln^{+}\frac{2\theta_{j}^{*}}{\delta},$ (49)
and
$T^{\mathrm{\textrm{sig}}}\leq\begin{cases}C_{1}(\theta_{j}^{*})^{-1}b_{0}^{-D}\ln(\frac{eb_{0}}{a_{0j}\sqrt{D}})&a_{0j}<b_{0}/\sqrt{D};\\ C_{1}(\theta_{j}^{*})^{-1}a_{0j}^{-1}\ln(\frac{ea_{0j}\sqrt{D}}{b_{0}})&a_{0j}>b_{0}/\sqrt{D},~~\text{and}~~D=1;\\ C_{1}(\theta_{j}^{*})^{-1}D^{-\frac{1}{2}}a_{0j}^{-1}b_{0}^{-D+1}&a_{0j}>b_{0}/\sqrt{D},~~\text{and}~~D>1,\end{cases}$ (50)
where both $C_{1}$ and $C_{2}$ are absolute constants.
For the choice $b_{0}=cD^{\frac{D+1}{D+2}}\varepsilon^{\frac{1}{D+2}}$, the second term on the right-hand side of Equation˜49 is dominated by the right-hand side of Equation˜50, and we can choose $c$ small enough so that the summation of the two terms is bounded by $b_{0}^{-D}\varepsilon^{-1}\ln n$. This justifies our choice of $t(\varepsilon)$.
Proposition C.2.
We consider the set $R$. We let $j\in R$, and suppose all the events $\{E_{k}\}$ in Lemma C.9 hold. Given $b_{0}$ and $t(\varepsilon)$ defined in Proposition C.1. For any positive constant $\delta^{\prime}$, if the eigenvalue $\lambda_{j}<n^{-5}$, and $n$ is large enough, then we have
$\tilde{\lambda}_{j}(t)<2\lambda_{j}^{1-2\delta^{\prime}}b_{0}^{D}\cdot n^{1+\delta^{\prime}}.$
Proof.
Since the events in Lemma C.9 hold, we have $|\xi_{j}|\leq 2(C_{\text{proxy}})^{-1/2}\sqrt{\frac{\ln\tilde{j}+\ln n}{n}}$, where $\tilde{j}=\frac{\tilde{d}}{\lambda_{j}}$. Since $j\in R$, we have $|\theta_{j}^{*}|\leq\tilde{\sigma}\leq\sqrt{c^{\prime}}\varepsilon$. Since $b_{0}^{D}t(\varepsilon)=\varepsilon^{-1}\ln n$, we can check that $t$ is no more than the hitting time $T_{2}$ defined by Equation˜68 in Lemma˜C.8 as follows.
Note that
$\varepsilon^{-1}\ln n(|\theta_{j}^{*}|+|\xi_{j}|)$ $\leq\sqrt{c^{\prime}}\ln n+\sqrt{\ln n}\sqrt{\frac{n}{\ln n\tilde{d}}}\cdot\sqrt{\frac{\ln n\tilde{d}+\ln\frac{1}{\lambda_{j}}}{n}}$
$<\ln n+\sqrt{\ln n+\ln\frac{1}{\lambda_{j}}}.$
Since $b_{0}=c\cdot D^{\frac{D+1}{D+2}}\varepsilon^{\frac{1}{D+2}}$, for $n$ large enough, we have $\lambda_{j}<n^{-5}<\frac{b_{0}^{2}}{D^{2}}$ and also
$\ln\frac{b_{0}/D}{\lambda_{j}^{\frac{1}{2}}}=\ln c+\frac{1}{2(D+2)}(\ln n+\ln\ln n\tilde{d})-\frac{1}{D+2}\ln D+\frac{1}{2}\ln\frac{1}{\lambda_{j}}>\ln n+\sqrt{\ln n+\ln\frac{1}{\lambda_{j}}}.$
It then follows that
$\beta_{j}(t)$ $<\lambda_{j}^{\frac{1}{2}}\exp\left(b_{0}^{D}t(|\theta_{j}^{*}|+|\xi_{j}|)\right)$
$\leq\lambda^{\frac{1}{2}}_{j}\exp\left(\sqrt{\ln n}\sqrt{\frac{n}{\ln n\tilde{d}}}\cdot\sqrt{\frac{\ln n\tilde{d}+\ln\frac{1}{\lambda_{j}}}{n}}\right)\cdot n$
$<\lambda^{\frac{1}{2}}_{j}\exp\left(\sqrt{\ln\frac{1}{\lambda_{j}}}\right)\cdot\exp\left(\sqrt{\ln n}\right)\cdot n$
$<\lambda_{j}^{\frac{1}{2}-\delta^{\prime}}n^{1+\delta^{\prime}}.$
Using $\tilde{\lambda}_{j}(t)=(\beta_{j}^{2}(t)+\lambda_{j})(b_{0}^{2}+D\beta_{j}^{2}(t))^{D}$, we obtain the desired result for sufficiently large $n$. ∎
Remark C.3.
The above proposition provides a very weak upper bound on $\tilde{\lambda}$, but it is sufficient to show that any eigenvalue $\lambda_{j}$, such that if given any constant $C$, $\lambda_{j}<n^{-C}$, then $\tilde{\lambda}_{j}(t)$ is also less than any polynomial of $n^{-1}$. Therefore, when considering the eigenvalue ordering problem, such signals can be ignored.
C.3 Conservation quantity
We omit the subscript $j$ in the following two sections C.3 and C.4 because all the proofs are similar for $j=1,2,\cdots,d$. By Equation˜39, it is easy to see that
$\frac{d}{dt}a^{2}=\frac{1}{D}\frac{d}{dt}b^{2}=\frac{d}{dt}\beta^{2}=2ab^{D}\beta(\theta^{*}-\theta+\xi).$
Consequently, we have
$a^{2}(t)-\beta^{2}(t)\equiv a_{0}^{2},~~~b^{2}(t)-D\beta^{2}(t)\equiv b_{0}^{2}.$ (51)
Using this, we see that
$a(t)=\left(\beta^{2}(t)+a_{0}^{2}\right)^{1/2},~~~~~b(t)=\left(D\beta^{2}(t)+b_{0}^{2}\right)^{1/2}>0.$
Using these conservation quantities, we can prove the following estimations in terms of $\beta$:
$\max\left(a_{0},|\beta|\right)\leq a\leq\sqrt{2}\max\left(a_{0},|\beta|\right)$ (52)
$\max\left(b_{0},\sqrt{D}|\beta|\right)\leq b\leq\sqrt{2}\max\left(b_{0},\sqrt{D}|\beta|\right)$
which also implies that $|\theta|=\left|ab^{D}\beta\right|\geq D^{D/2}|\beta|^{D+2}$. The evolution of $\theta$. It is direct to compute that
$\dot{\theta}$ $=\dot{a}b^{D}\beta+aDb^{D-1}\dot{b}\beta+ab^{D}\dot{\beta}$ (53)
$=\left[\left(b^{D}\beta\right)^{2}+\left(Dab^{D-1}\beta\right)^{2}+\left(ab^{D}\right)^{2}\right](\theta^{*}-\theta+\xi)$
$=\theta^{2}\left(a^{-2}+D^{2}b^{-2}+\beta^{-2}\right)(\theta^{*}-\theta+\xi).$
And we also have
$|\theta|=\left|ab^{D}\beta\right|\geq D^{D/2}|\beta|^{D+2}\quad\Longrightarrow|\beta|\leq\left(D^{-D/2}|\theta|\right)^{1/(D+2)}.$ (54)
Therefore,
$\theta^{2}\left(a^{-2}+D^{2}b^{-2}+\beta^{-2}\right)\geq\theta^{2}\beta^{-2}\geq D^{-\frac{D}{D+2}}|\theta|^{\frac{2D+2}{D+2}}.$ (55)
C.4 Multi-layer dynamic
We study the dynamic of the ODE for any given $j$. Before the analysis, we streamline some notations.
Assume for some $\kappa_{j}>0$, it holds that $|\xi_{j}|\leq\kappa_{j}$. (Note that this $\kappa_{j}$ can be the high probability upper bound derived using Lemma˜C.9.) Since $j$ is given, we drop the the subscript $j$ to simplify the exposition throughout this subsection; for example, we write $\lambda$ for $\lambda_{j}$ and $\theta^{*}$ for $\theta_{j}^{*}$. We write $\ln^{+}(x)=\max(1,\ln(x))$ for any $x>0$.
Lemma C.4 (Monotonicity from equation).
Consider the equation Equation˜39. Suppose $y>0$.
  1. 1.
    $a(t),\beta(t)$, and $\theta(t)$ are all non-negative and increasing.
  2. 2.
    We have
    $y\geq\theta(t)\geq 0\quad\forall t\geq 0.$
  3. 3.
    Since $y=\theta^{*}+\xi$ and $|\xi|\leq\kappa$, we have
    $|\theta^{*}-\theta(t)|\leq|\theta^{*}|+\kappa,\quad\forall t\geq 0.$
  4. 4.
    $|\theta^{*}-\theta(t)|$ is decreasing provided that $|\theta^{*}-\theta(t)|>\kappa$.
  5. 5.
    If $\left|\theta^{*}-\theta\left(t_{1}\right)\right|\leq\kappa$ for some $t_{1}$, we have
    $|\theta^{*}-\theta(t)|\leq\kappa\text{ for all }t\geq t_{1}.$
If $y<0$, Items 3, 4, and 5 still hold, while Items 1 and 2 can be modified by symmetry.
Proof.
Items 1 and 2 are directly implied from Equation˜39. Item 3 is implied by Item 2.
To prove Item 4, consider Equation˜53, from which we have
$\dot{\theta}=\theta^{2}\left(a^{-2}+D^{2}b^{-2}+\beta^{-2}\right)(y-\theta),$
which implies $\dot{\theta}\geq 0$.
Since $|\theta^{*}-\theta(t)|>\kappa$, we have either $\theta(t)>\theta^{*}+\kappa$ or $\theta(t)<\theta^{*}-\kappa$.
The first case is not possible; otherwise, we have $0<y=\xi+\theta^{*}\leq\kappa+\theta^{*}<\theta\leq y$, which is a contradiction. In the second case, we have $|\theta^{*}-\theta(t)|=\theta^{*}-\theta(t)$, which is decreasing because $\dot{\theta}\geq 0$.
Item 5 is implied by Item 4.
Lemma C.5 (Approaching from below).
Consider the equation Equation˜39. Suppose $\theta^{*}\geq 8\kappa$ (similar results hold for $\theta^{*}\leq-8\kappa$ by symmetry). Suppose $t_{0}\geq 0$ such that $0\leq\theta\left(t_{0}\right)<\frac{1}{4}\theta^{*}$. Define
$T^{\operatorname{\textrm{sig}}}=\inf\left\{s\geq 0:\theta\left(t_{0}+s\right)\geq\theta^{*}/4\right\}.$
This is the extra time needed from $t_{0}$ for $\theta$ to reach $\theta^{*}/4$. We have
$T^{\mathrm{\textrm{sig}}}\leq\begin{cases}4(\theta^{*})^{-1}b_{0}^{-D}\ln(\frac{b_{0}}{a_{0}\sqrt{D}})&a_{0}<b_{0}/\sqrt{D};\\ 4(\theta^{*})^{-1}a_{0}^{-1}\ln(\frac{a_{0}\sqrt{D}}{b_{0}})&a_{0}>b_{0}/\sqrt{D},~~\text{and}~~D=1;\\ 4(\theta^{*})^{-1}D^{-\frac{1}{2}}a_{0}^{-1}b_{0}^{-D+1}&a_{0}>b_{0}/\sqrt{D},~~\text{and}~~D>1.\end{cases}$ (56)
Proof.
[…]
$y-\theta(t)=\theta^{*}-\theta(t)+\xi\geq\frac{3}{4}\theta^{*}-\kappa\geq\frac{1}{2}\theta^{*}.$
Let $r=\min\left(a_{0},b_{0}/D^{\frac{1}{2}}\right)$ and $R=\max\left(a_{0},b_{0}/D^{\frac{1}{2}}\right)$. Define the following time point if it exists:
$T^{\textrm{pos},1}=\inf\left\{s\geq 0:\beta\left(t_{0}+s\right)\geq r\right\};~~T^{\textrm{pos},2}=\inf\left\{s\geq 0:\beta\left(t_{0}+s\right)\geq R\right\}$
$T^{\textrm{sig}}=\inf\left\{s\geq 0:\left|\theta^{*}-\theta\left(t_{0}+s\right)\right|\leq\frac{3}{4}\theta^{*}\right\}.$
We will first bound both $T^{\textrm{pos},1}$ and $T^{\textrm{pos},2}$.
From Equation˜39, we have
$\dot{\beta}(t)=a(t)b^{D}(t)[\theta^{*}+\xi-\theta(t)]\geq\frac{1}{4}\theta^{*}a(t)b^{D}(t),\quad\text{ for }\quad t\leq t_{0}+T^{\textrm{sig}}.$ (57)
Stage 1: $0\leq s\leq T^{\textrm{pos},1}$. Note that $\sqrt{2}a_{0}>a(t)>a_{0}$, and $e\cdot b_{0}^{D}>b(t)^{D}=(D\beta(t)^{2}+b_{0}^{2})^{\frac{D}{2}}>b_{0}^{D}$. We have
$\dot{\beta}\left(t_{0}+s\right)\geq\frac{1}{4}\theta^{*}a_{0}b_{0}^{D}\geq\frac{1}{4}\theta^{*}a_{0}b_{0}^{D},$
which suggests $\beta$ increases at least linearly. Therefore, we have
$T^{\textrm{pos},1}\leq 8r\left(\theta^{*}a_{0}b_{0}^{D}\right)^{-1}.$ (58)
Stage 2: $T^{\textrm{pos},1}\leq s\leq T^{\textrm{pos},2}$. Consider two cases.
Case 1: If $a_{0}<b_{0}/\sqrt{D}$, $r=a_{0}$ and $R=b_{0}/\sqrt{D}$. Note $a\geq\beta$ in Equation˜52. We use Equation˜57 to get
$\dot{\beta}\left(t_{0}+s\right)\geq\frac{1}{4}\theta^{*}b_{0}^{D}\left|\beta\left(t_{0}+s\right)\right|,$
By Grönwall’s inequality, we have
$T^{\textrm{pos},2}-T^{\textrm{pos},1}\leq 4\left(\theta^{*}b_{0}^{D}\right)^{-1}\ln\frac{b_{0}}{a_{0}\sqrt{D}}.$ (59)
Case 2: If $a_{0}>b_{0}/\sqrt{D}$, $R=a_{0}$ and $r=b_{0}/\sqrt{D}$. We use $b\geq\sqrt{D}\beta$ in Equation˜52 together with $a>a_{0}$, $b>\sqrt{D}|\beta|$ and Equation˜57 to get that
$\dot{\beta}\left(t_{0}+s\right)\geq\frac{1}{4}\theta^{*}a_{0}D^{\frac{D}{2}}\left|\beta\left(t_{0}+s\right)\right|^{D}.$
By comparison theorem, we have
$T^{\textrm{pos},2}-T^{\textrm{pos},1}\leq\left\{\begin{array}[]{l}4\left(\theta^{*}a_{0}\right)^{-1}\ln\frac{a_{0}\sqrt{D}}{b_{0}},\quad\text{ if }\quad D=1;\\ {4\left((D-1)\theta^{*}a_{0}D^{D/2}\right)^{-1}\left[\left(b_{0}/\sqrt{D}\right)^{-(D-1)}-a_{0}^{-(D-1)}\right],\quad\text{ if }\quad D\geq 2}.\end{array}\right$. (60)
Stage 3: If $T^{\textrm{sig}}\leq T^{\textrm{pos},2}$, then we can use the for $T^{\textrm{sig}}$ in Stage 2 as a bound for $T^{\textrm{sig}}$. Now, we consider the case $T^{\textrm{pos},2}<T^{\textrm{sig}}$. We combine Equation˜52 with $a>|\beta|$, $b>\sqrt{D}|\beta|$, and Equation˜57 to get
$\dot{\beta}\left(t_{0}+T^{\textrm{pos},2}+s\right)\geq\frac{1}{4}\theta^{*}D^{D/2}|\beta\left(T^{\textrm{pos},2}+s\right)|^{D+1},\quad\text{ for }\quad s\in\left[0,T^{\textrm{sig}}-T^{\textrm{pos},2}\right].$
Beside, we have $\beta\left(t_{0}+T^{\textrm{pos},2}\right)=R>0$. By Lemma C.10, we have
$T^{\textrm{sig}}-T^{\textrm{pos},2}\leq 4D^{-\frac{D+2}{2}}(\theta^{*})^{-1}R^{-D}.$ (61)
We now bound $T^{\mathrm{\textrm{sig}}}$ using the summation of Equation˜58, Equation˜61, and Equation˜59 if $a_{0}<b_{0}/\sqrt{D}$, or the summation of Equation˜58, Equation˜61, and Equation˜60 if $a_{0}>b_{0}/\sqrt{D}$.
If $a_{0}<b_{0}/\sqrt{D}$, we can bound the right hand sides of Equation˜58 and Equation˜61 by $8(\theta^{*})^{-1}b_{0}^{-D}$ and $4(\theta^{*})^{-1}b_{0}^{-D}$ respectively.
If $a_{0}>b_{0}/\sqrt{D}$, we can bound the right hand sides of Equation˜58 and Equation˜61 by $8(\theta^{*})^{-1}D^{-\frac{1}{2}}a_{0}^{-1}b_{0}^{-D+1}$ and $4(\theta^{*})^{-1}D^{-\frac{1}{2}}a_{0}^{-1}b_{0}^{-D+1}$ respectively. Furthermore, if $D>1$, we can bound Equation˜60 by $4(\theta^{*})^{-1}(D-1)^{-1}D^{-\frac{1}{2}}a_{0}^{-1}b_{0}^{-D+1}$.
This leads to
$T^{\mathrm{\textrm{sig}}}\leq\begin{cases}4(\theta^{*})^{-1}b_{0}^{-D}\left(3+\ln(\frac{b_{0}}{a_{0}\sqrt{D}})\right)&a_{0}<b_{0}/\sqrt{D};\\ 4(\theta^{*})^{-1}a_{0}^{-1}\left(3+\ln(\frac{a_{0}\sqrt{D}}{b_{0}})\right)&a_{0}>b_{0}/\sqrt{D},~~\text{and}~~D=1;\\ 16(\theta^{*})^{-1}D^{-\frac{1}{2}}a_{0}^{-1}b_{0}^{-D+1}&a_{0}>b_{0}/\sqrt{D},~~\text{and}~~D>1.\end{cases}$ (62)
Lemma C.6 (Approximation time near $\theta^{*}$).
Consider the equation Equation˜39 with $\theta^{*}\geq 0$ ( $a$ similar result holds for $\theta^{*}\leq 0$ ). Suppose $\theta^{*}>8\kappa$. Suppose for some $t_{0}\geq 0$ such that
$\frac{1}{4}\theta^{*}\leq\theta\left(t_{0}\right)\leq\theta^{*}-\kappa.$
Then, for any $\delta>0$, we have
$|\theta^{*}-\theta(t)|\leq\kappa+\delta,\quad\forall t\geq t_{0}+4^{\frac{2D+2}{D+2}}D^{\frac{D}{D+2}}(\theta^{*})^{-\frac{2D+2}{D+2}}\ln^{+}\frac{\left|\theta^{*}-\theta\left(t_{0}\right)\right|-\kappa}{\delta}.$
Proof.
Given any $\delta>0$, if $\theta(t_{0})\geq\theta^{*}-\kappa-\delta$, we have $|\theta^{*}-\theta(t)|\leq\kappa+\delta$ for all $t\geq t_{0}$ by Lemma˜C.4 (Item 4) and the desired result is proved.
Next, suppose $\theta(t_{0})<\theta^{*}-\kappa-\delta$. Define
$T^{\textrm{app}}=\inf\left\{s\geq 0:\left|\theta^{*}-\theta\left(t_{0}+s\right)\right|\leq\kappa+\delta\right\}.$
By Lemma˜C.4 (Item 4) again, it suffices to provide an upper bound on $T^{\text{app }}$.
For all $t\geq t_{0}$, we have $\frac{1}{4}\theta^{*}\leq$ $\theta(t)$ by Lemma˜C.4 (Item 1). Consequently, Equation˜55 implies that
$\theta^{2}\left(a^{-2}+D^{2}b^{-2}+\beta^{-2}\right)\geq D^{-\frac{D}{D+2}}|\theta|^{\frac{2D+2}{D+2}}\geq 4^{-\frac{2D+2}{D+2}}D^{-\frac{D}{D+2}}(\theta^{*})^{\frac{2D+2}{D+2}}:=c_{0}.$
Furthermore, by Equation˜53, we have
$\dot{\theta}=\theta^{2}\left(a^{-2}+D^{2}b^{-2}+\beta^{-2}\right)(\theta^{*}-\theta+\xi)\geq c_{0}(\theta^{*}-\kappa-\theta).$
Let $x(s):=\theta^{*}-\kappa-\theta(t_{0}+s)$ with $x(0)=\theta^{*}-\kappa-\theta(t_{0})$. Note that $T^{app}$ is the hitting time of $x(s)$ to $\delta$. Applying Lemma˜C.11 to $x(s)$, we have
$T^{\text{app }}\leq c_{0}^{-1}\ln\frac{\left|\theta^{*}-\theta\left(t_{0}\right)\right|-\kappa}{\delta}.$
Lemma C.7.
Consider the equation Equation˜39 with $\theta^{*}\geq 0$ (a similar result holds for $\theta^{*}\leq 0$ ).
Suppose $\theta^{*}\geq 8\kappa$. For two absolute constants $C_{1},C_{2}$, we have
$|\theta^{*}-\theta(t)|\leq\kappa+\delta,\quad\forall t\geq\bar{T}^{app}(\delta),$
where
$\bar{T}^{app}(\delta):=\bar{T}^{\textrm{sig}}+C_{2}D^{\frac{D}{D+2}}(\theta^{*})^{-\frac{2D+2}{D+2}}\ln^{+}\frac{\theta^{*}}{\delta},$ (63)
and
$\bar{T}^{\mathrm{\textrm{sig}}}:=\begin{cases}C_{1}(\theta^{*})^{-1}b_{0}^{-D}\ln(\frac{b_{0}}{a_{0}\sqrt{D}})&a_{0}<b_{0}/\sqrt{D};\\ C_{1}(\theta^{*})^{-1}a_{0}^{-1}\ln(\frac{a_{0}\sqrt{D}}{b_{0}})&a_{0}>b_{0}/\sqrt{D},~~\text{and}~~D=1;\\ C_{1}(\theta^{*})^{-1}D^{-\frac{1}{2}}a_{0}^{-1}b_{0}^{-D+1}&a_{0}>b_{0}/\sqrt{D},~~\text{and}~~D>1.\end{cases}$ (64)
Proof.
We will repeatedly apply the monotonicity of Lemma˜C.4.
Recall $T^{\mathrm{\textrm{sig}}}$ defined in Lemma˜C.5 with $t_{0}=0$ and let $t_{1}$ be the upper bound on $T^{\mathrm{\textrm{sig}}}$ we found therein. Then $\theta(t_{1})\geq\frac{\theta^{*}}{4}$.
We then apply Lemma˜C.6 with $t_{0}=t_{1}$, and conclude that $|\theta^{*}-\theta(t)|\leq\kappa+\delta$ for all $t\geq t_{1}+t_{2}$, where $t_{2}=4^{\frac{2D+2}{D+2}}D^{\frac{D}{D+2}}(\theta^{*})^{-\frac{2D+2}{D+2}}\ln^{+}\frac{\left|\theta^{*}-\theta\left(t_{1}\right)\right|-\kappa}{\delta}$. Note that $\left|\theta^{*}-\theta\left(t_{1}\right)\right|-\kappa\leq\theta^{*}$. We complete the proof by defining $\bar{T}^{\mathrm{\textrm{sig}}}=t_{1}$ and $\bar{T}^{app}(\delta)=t_{1}+4^{\frac{2D+2}{D+2}}D^{\frac{D}{D+2}}(\theta^{*})^{-\frac{2D+2}{D+2}}\ln^{+}\frac{\theta^{*}}{\delta}$.
Lemma C.8.
Consider the equation Equation˜39. Denote $r^{\prime}=\min\{a_{0},b_{0}/D\},R^{\prime}=\max\{a_{0},b_{0}/D\}$. W.L.O.G., We assume that $\theta^{*}\geq 0$. Define $T_{1}=\inf\{t:|\beta(t)|>r^{\prime}\}$, and $T_{2}=\inf\{t:|\beta(t)|>R^{\prime}\}$. If $D\geq 1$, and $t$ satisfies the following:
$\sqrt{2e}a_{0}b_{0}^{D}\int_{0}^{t}(|\theta^{*}|+|\xi|)\mathrm{d}s\leq\min\left(a_{0},b_{0}/D\right),$
then we have
$|\theta(t)|\leq 2e\cdot a_{0}^{2}b_{0}^{2D}\int_{0}^{t}(|\theta^{*}|+|\xi|)\mathrm{d}s.$ (65)
Moreover, if $a_{0}\leq b_{0}/D$ and $0\leq t\leq T_{2}-T_{1}$ satisfies the following:
$\sqrt{2e}b_{0}^{D}\int_{0}^{t}(|\theta^{*}|+|\xi|)\mathrm{d}s\leq\ln\frac{b_{0}/D}{a_{0}},$
then we have
$|\beta(t)|\leq a_{0}\exp\left(b_{0}^{D}\int_{0}^{t}(|\theta^{*}|+|\xi|)\mathrm{d}s\right);$
$|\theta(t)|\leq\sqrt{2e}a_{0}^{2}b_{0}^{D}\exp\left(2\sqrt{e}b_{0}^{D}\int_{0}^{t}(|\theta^{*}|+|\xi|)\mathrm{d}s\right).$
Proof.
From Equation˜39, we have
$|\beta(t)|\leq\int_{0}^{t}a(s)b^{D}(s)(|\theta^{*}|+|\xi|)ds.$
Consider $t\leq T_{1}$. We use Equation˜52 to get $a(t)\leq\sqrt{2}a_{0}$ and by Equation˜51, $b^{2}(t)-D\beta^{2}(t)\equiv b_{0}^{2}$. Consequently, we have $b^{2}(t)\leq(1+\frac{1}{D})b_{0}^{2}$, and thus
$|\beta(t)|\leq\sqrt{2e}a_{0}b_{0}^{D}\int_{0}^{t}(|\theta^{*}|+|\xi|)dt,$ (66)
which implies Equation˜65 by using the fact that $|\theta|=|ab^{D}\beta|$. Furthermore, Equation˜66 implies that
$T_{1}\geq\inf\left\{t\geq 0:\sqrt{2e}a_{0}b_{0}^{D}\int_{0}^{t}(|\theta^{*}|+|\xi|)ds\geq r^{\prime}\right\}.$
Then when $t>T_{1}$, in the following, suppose $a_{0}\leq b_{0}/D$. We have $r^{\prime}=a_{0}$ and $R^{\prime}=b_{0}/D$.
Consider $t\in(T_{1},T_{2})$. We have $a(t)\leq\sqrt{2}|\beta(t)|$ and $b(t)^{2}\leq(1+\frac{1}{D})b_{0}^{2}$. Consequently, Equation˜39 implies that
$|\beta(t)|\leq a_{0}+\sqrt{2e}b_{0}^{D}\int_{T_{1}}^{t}|\beta(s)|(|\theta^{*}|+|\xi|)\mathrm{d}s,$ (67)
for any $t\in(T_{1},T_{2})$. By Grönwall inequality, we have
$\left|\beta\left(t\right)\right|\leq a_{0}\exp\left(\sqrt{2e}b_{0}^{D}\int_{T_{1}}^{t}(|\theta^{*}|+|\xi|)\mathrm{d}s\right),\quad t\in(T_{1},T_{2}).$
By definition of $T_{2}$, we have
$T_{2}\geq\inf\left\{t\geq T_{1}:\sqrt{2e}b_{0}^{D}\int_{T_{1}}^{t}(|\theta^{*}|+|\xi|)\mathrm{d}s=\ln\frac{b_{0}/D}{a_{0}}\right\}.$ (68)
The bound for $\theta(t)$ now follows from using the bounds $a(t)\leq\sqrt{2}|\beta|,b^{2}(t)\leq(1+\frac{1}{D})b_{0}^{2}$ to get
$\left|\theta\left(t\right)\right|=\left|ab^{D}\beta\right|\leq\sqrt{2e}b_{0}^{D}|\beta|^{2}\leq\sqrt{2e}a_{0}^{2}b_{0}^{D}\exp\left(2\sqrt{e}b_{0}^{D}\int_{T_{1}}^{t}(|\theta^{*}|+|\xi|)\mathrm{d}s\right),\forall t\in(T_{1},T_{2}).$
C.5 Auxiliary Lemma
The following lemma provides a choice of $\kappa_{j}\geq|\xi_{j}|$ with high probability.
Lemma C.9.
Recall $S$ defined by Assumption 6.1 and let $C=2(C_{\text{proxy}})^{-1/2}$. For $k\in S$, we introduce the events $\{E_{k}\}$ as follows:
$E_{k}:=\{\left|\xi_{k}\right|\leq Cn^{-1/2}\sqrt{\ln n}\}.$ (69)
For $k\in S^{C}$, we introduce the events $\{E_{k}\}$ as follows:
$E_{k}:=\left\{\left|\xi_{k}\right|\leq Cn^{-1/2}\sqrt{\ln(n\tilde{k})}\right\}$. (70)
where $\tilde{k}=\sum_{j}\lambda_{j}/\lambda_{k}$.
Then, with probability at least $1-\frac{4}{n}$, all events $E_{k},k\in[d]$ hold simultaneously.
Proof.
By ˜6.1, the noise $\xi_{k}$ is sub-Gaussian with variance proxy $C_{\text{proxy}}/n$. Therefore, $\mathbf{P}\left(\left|\xi_{k}\right|\geq s\right)\leq 2\exp(-(2C_{\text{proxy}})^{-1}ns^{2})$.
If $k\in S$, we have
$\mathbf{P}\left\{\left|\xi_{k}\right|\geq 2C_{\text{proxy}}^{-1/2}\sqrt{\frac{\ln n}{n}}\right\}\leq 2\exp\left(-2(\ln n)\right).$
By the union bound, we have
$\mathbf{P}\left\{\cap_{k\in S}E_{k}\right\}$ $\geq 1-\sum_{k\in S}\mathbf{P}\left\{\left|\xi_{k}\right|\geq 2C_{\text{proxy}}^{-1/2}\sqrt{\frac{\ln n}{n}}\right\}$ (71)
$\geq 1-|S|2\exp\left(-2(\ln n)\right)$
$\geq 1-\frac{2}{n},$
where the last inequality is because $|S|2\exp\left(-2(\ln n)\right)\leq 2n^{-1}$.
If $k\in S^{C}$, we have
$\mathbf{P}\left\{\left|\xi_{k}\right|\geq 2(C_{\text{proxy}})^{-1/2}\sqrt{\frac{\ln n\tilde{k}}{n}}\right\}\leq 2\exp\left(-(\ln n+\ln\frac{\sum_{j}\lambda_{j}}{\lambda_{k}})\right)\leq\frac{2}{n}\cdot\frac{\lambda_{k}}{\sum_{j}\lambda_{j}},$ (72)
where we recall that $\tilde{k}=\frac{\sum_{j}\lambda_{j}}{\lambda_{k}}$.
By the union bound, we have
$\mathbf{P}\left\{\cap_{k\in S^{C}}E_{k}\right\}$ $\geq 1-\sum_{k\in S^{C}}\mathbf{P}\left\{\left|\xi_{k}\right|\geq 2(C_{\text{proxy}})^{-1/2}\sqrt{\frac{\ln n\tilde{k}}{n}}\right\}$ (73)
$\geq 1-\sum_{k\in S^{C}}\frac{2}{n}\cdot\frac{\lambda_{k}}{\sum_{j}\lambda_{j}}$
$\geq 1-\frac{2}{n}.$
We combined the Equation˜71 and Equation˜73, and we derive the results.
The following two lemmas provide convenient upper bounds on hitting times of ODE solutions.
Lemma C.10.
Let $k>0$ and $p>1$.
  • Consider the ODE
    $\dot{x}\geq kx^{p},\quad x(0)=x_{0}>0$
    Then we have
    $x(t)\geq\left(x_{0}^{-(p-1)}-(p-1)kt\right)^{-\frac{1}{p-1}}$
    and thus for any $M\geq 0$,
    $\inf\{t\geq 0:x(t)\geq M\}\leq\left[(p-1)kx_{0}^{p-1}\right]^{-1}.$ (74)
  • Consider the ODE
    $\dot{x}\leq-kx^{p},\quad x(0)=x_{0}>0.$
    Then we have
    $x(t)\leq\left(x_{0}^{-(p-1)}+(p-1)kt\right)^{-\frac{1}{p-1}},$
    and thus for any $M>0$,
    $\inf\{t\geq 0:x(t)\leq M\}\leq\left[(p-1)kM^{p-1}\right]^{-1}.$ (75)
Lemma C.11.
Let $k>0$ and $x_{0}>0$.
  1. 1.
    If
    $\dot{x}(t)\;\geq\;k\,x(t),\qquad x(0)=x_{0},$
    then for all $t\geq 0$, it holds that
    $x(t)\;\geq\;x_{0}\,e^{kt},$
    and for every $M\geq x_{0}$, we have
    $\inf\{t\geq 0:\,x(t)\geq M\}\;\leq\;\frac{1}{k}\log\!\Bigl(\frac{M}{x_{0}}\Bigr).$
  2. 2.
    If
    $\dot{x}(t)\;\leq\;-k\,x(t),\qquad x(0)=x_{0},$
    then for all $t\geq 0$, it holds that
    $x(t)\;\leq\;x_{0}\,e^{-kt},$
    and for every $0<M\leq x_{0}$, we have
    $\inf\{t\geq 0:\,x(t)\leq M\}\;\leq\;\frac{1}{k}\log\!\Bigl(\frac{x_{0}}{M}\Bigr).$
Appendix D Related Work on Principal Component Regression
As discussed in Section˜3.1, the PC estimator serves as a motivating example for the concepts of ESD and span profile due to its clear illustration of bias-variance trade-offs. However, the ESD and span profile are designed to characterize the intrinsic difficulty of generalization arising from signal-kernel alignment, and their definitions do not rely on the PC estimator. Nonetheless, the analysis of PC estimators, particularly in high-dimensional linear regression, been an active area of recent research. Below, we briefly summarize some relevant contributions to provide context.
D.1 Proportional Asymptotics
Several studies analyze Principal Component Regression (PCR) in the proportional asymptotic setting where the dimension $p$ and sample size $n$ grow with $p/n\!\to\!\gamma$. In this regime, Xu and Hsu [46] study the limiting risk of PCR with Gaussian designs with diagonal covariance. They assume polynomially decaying eigenvalues or a convergent empirical spectrum, together with an isotropic prior, and they reveal a “double-descent” risk curve. In a related vein, Wu and Xu [45] extend the analysis to a general covariance matrix $\Sigma_{x}$ and an anisotropic prior satisfying $\mathbb{E}\beta_{*}\beta_{*}^{\!\top}\!=\!\Sigma_{\beta}$. They also derive an exact risk expression and demonstrate how “misalignment” between $\Sigma_{x}$ and $\Sigma_{\beta}$ affects risk; here “alignment” refers to concordance between the orderings of their eigenvalues. Both studies assume knowledge of the eigenvectors of the population covariance matrix $\Sigma_{x}$ to construct the oracle PCR. Gedon et al. [17] analyze the limiting risk of PCR under a latent factor model and explore the effect of distribution shift. Green and Romanov [20] derive the exact limits of estimation risk, in-sample prediction risk, and out-of-sample prediction risk of PCR under the assumption that both the empirical spectrum distribution and the distribution of the true signal’s mass over the eigenspaces of $\Sigma_{x}$ converges weakly.
D.2 Non-Asymptotic Analysis
Complementary research develops non-asymptotic guarantees. Agarwal et al. [1] derive finite-sample upper bounds on prediction error using $\|\beta^{*}\|_{1}^{2}$ and the rank of the design matrix under latent factor models, and they explore the robustness of PCR to noise and missing values in the observed covariates. Bing et al. [9] consider PCR with an adaptively selected number of components under latent factor models and provide alternative finite-sample risk bounds using $\|\beta^{*}\|_{2}^{2}$. Huang et al. [21] derive non-asymptotic risk bounds for PCR in more general settings by analyzing the alignment between population and empirical principal components. Hucker and Wahl [22] derive non-asymptotic error bounds for PCR in kernel regression.